GNU PROLOG

A Native Prolog Compiler with Constraint Solving over Finite Domains

Edition 1.7, for GNU Prolog version 1.2.16 September 25, 2002

by Daniel Diaz

Copyright (C) 1999-2002 Daniel Diaz

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation^{1}, 59 Temple Place - Suite 330, Boston, MA 02111, USA.

1http://www.fsf.org/

Contents

1 Acknowledgements 9

2 Introduction 11

3 Using GNU Prolog 13 3.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.2 The GNU Prolog interactive interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3.2.1 Starting/exiting the interactive interpreter . . . . . . . . . . . . . . . . . . . . . . 13 3.2.2 The interactive interpreter read-execute-write loop . . . . . . . . . . . . . . . . . 14 3.2.3 Consulting a Prolog program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2.4 Interrupting a query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2.5 The line editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3.3 Adjusting the size of Prolog stacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.4 The GNU Prolog compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3.4.1 Different kinds of codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.4.2 Compilation scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.4.3 Using the compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.4 Running an executable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.4.5 Generating a new interactive interpreter . . . . . . . . . . . . . . . . . . . . . . . 25 3.4.6 The hexadecimal predicate name encoding . . . . . . . . . . . . . . . . . . . . . . 26

4 Debugging 29 4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.2 The procedure box model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.3 Debugging predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

4.3.1 Running and stopping the debugger . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.3.2 Leashing ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.3.3 Spy-points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

4.4 Debugging messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.5 Debugger commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.6 The WAM debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

5 Format of definitions 35 5.1 General format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 5.2 Types and modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 5.3 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

5.3.1 General format and error context . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.3.2 Instantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.3.3 Type error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 5.3.4 Domain error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 5.3.5 Existence error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.3.6 Permission error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.3.7 Representation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.3.8 Evaluation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.3.9 Resource error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.3.10 Syntax error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.3.11 System error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

6 Prolog directives and control constructs 41

6.1 Prolog directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.1.2 dynamic/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.1.3 public/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.1.4 multifile/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.1.5 discontiguous/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.1.6 ensure linked/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

6.1.7 built in/0, built in/1, built in fd/0, built in fd/1 . . . . . . . . . . . . . . 43

6.1.8 include/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.1.9 ensure loaded/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.1.10 op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.1.11 char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.1.12 set prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.1.13 initialization/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 6.1.14 foreign/2, foreign/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

6.2 Prolog control constructs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 6.2.1 true/0, fail/0, !/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 6.2.2 (',')/2 - conjunction, (;)/2 - disjunction, (->)/2 - if-then . . . . . . . . . . . . 46 6.2.3 call/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 6.2.4 catch/3, throw/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7 Prolog built-in predicates 49

7.1 Type testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

7.1.1 var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1,
       compound/1, callable/1, list/1, partial list/1, list or partial list/1 . . 49

7.2 Term unification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.2.1 (=)/2 - Prolog unification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.2.2 unify with occurs check/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.2.3 (\=)/2 - not Prolog unifiable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

7.3 Term comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.3.1 Standard total ordering of terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.3.2 (==)/2 - term identical, (\==)/2 - term not identical,

(@<)/2 - term less than, (@=<)/2 - term less than or equal to,

(@>)/2 - term greater than, (@>=)/2 - term greater than or equal to . . . . . . . 51 7.3.3 compare/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

7.4 Term processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.4.1 functor/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.4.2 arg/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.4.3 (=..)/2 - univ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.4.4 copy term/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.4.5 setarg/4, setarg/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

7.5 Variable naming/numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 7.5.1 name singleton vars/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 7.5.2 name query vars/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

7.5.3 bind variables/2, numbervars/3, numbervars/1 . . . . . . . . . . . . . . . . . . 56

7.5.4 term ref/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.6 Arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7.6.1 Evaluation of an arithmetic expression . . . . . . . . . . . . . . . . . . . . . . . . 57 7.6.2 (is)/2 - evaluate expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.6.3 (=:=)/2 - arithmetic equal, (=\=)/2 - arithmetic not equal,

(<)/2 - arithmetic less than, (=<)/2 - arithmetic less than or equal to,

(>)/2 - arithmetic greater than, (>=)/2 - arithmetic greater than or equal to . . 60

7.7 Dynamic clause management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.7.2 asserta/1, assertz/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.7.3 retract/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 7.7.4 retractall/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 7.7.5 clause/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 7.7.6 abolish/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

7.8 Predicate information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.8.1 current predicate/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.8.2 predicate property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

7.9 All solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.9.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.9.2 findall/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.9.3 bagof/3, setof/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7.10 Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.10.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.10.2 current input/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.10.3 current output/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.10.4 set input/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.10.5 set output/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.10.6 open/4, open/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.10.7 close/2, close/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.10.8 flush output/1, flush output/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.10.9 current stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.10.10 stream property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.10.11 at end of stream/1, at end of stream/0 . . . . . . . . . . . . . . . . . . . . . . 74 7.10.12 stream position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 7.10.13 set stream position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.10.14 seek/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.10.15 character count/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 7.10.16 line count/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.10.17 line position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.10.18 stream line column/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.10.19 set stream line column/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 7.10.20 add stream alias/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.10.21 current alias/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.10.22 add stream mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.10.23 remove stream mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 7.10.24 current mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 7.10.25 set stream type/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.10.26 set stream eof action/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.10.27 set stream buffering/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

7.11 Constant term streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 7.11.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

7.11.2 open input atom stream/2, open input chars stream/2,

open input codes stream/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

7.11.3 close input atom stream/1, close input chars stream/1,

close input codes stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

7.11.4 open output atom stream/1, open output chars stream/1,

open output codes stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

7.11.5 close output atom stream/2, close output chars stream/2,

close output codes stream/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

7.12 Character input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

7.12.1 get char/2, get char/1, get code/1, get code/2 . . . . . . . . . . . . . . . . . 85
7.12.2 get key/2, get key/1 get key no echo/2, get key no echo/1 . . . . . . . . . . 86
7.12.3 peek char/2, peek char/1, peek code/1, peek code/2 . . . . . . . . . . . . . . . 87
7.12.4 unget char/2, unget char/1, unget code/2, unget code/1 . . . . . . . . . . . . 88
7.12.5 put char/2, put char/1, put code/1, put code/2, nl/1, nl/0 . . . . . . . . . . 88

7.13 Byte input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.13.1 get byte/2, get byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.13.2 peek byte/2, peek byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.13.3 unget byte/2, unget byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.13.4 put byte/2, put byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

7.14 Term input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

7.14.1 read term/3, read term/2, read/2, read/1 . . . . . . . . . . . . . . . . . . . . . 92
7.14.2 read atom/2, read atom/1, read integer/2, read integer/1,

read number/2, read number/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.14.3 read token/2, read token/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.14.4 syntax error info/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 7.14.5 last read start line column/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

7.14.6 write term/3, write term/2, write/2, write/1, writeq/2, writeq/1,
       write canonical/2, write canonical/1, display/2, display/1, print/2,

print/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.14.7 format/3, format/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 7.14.8 portray clause/2, portray clause/1 . . . . . . . . . . . . . . . . . . . . . . . . 99 7.14.9 get print stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.14.10 op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.14.11 current op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 7.14.12 char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 7.14.13 current char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

7.15 Input/output from/to constant terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

7.15.1 read term from atom/3, read from atom/2, read token from atom/2 . . . . . . 104
7.15.2 read term from chars/3, read from chars/2, read token from chars/2 . . . . 104
7.15.3 read term from codes/3, read from codes/2, read token from codes/2 . . . . 105
7.15.4 write term to atom/3, write to atom/2, writeq to atom/2,
       write canonical to atom/2, display to atom/2, print to atom/2,

format to atom/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

7.15.5 write term to chars/3, write to chars/2, writeq to chars/2,
       write canonical to chars/2, display to chars/2, print to chars/2,

format to chars/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

7.15.6 write term to codes/3, write to codes/2, writeq to codes/2,
       write canonical to codes/2, display to codes/2, print to codes/2,

format to codes/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

7.16 DEC-10 compatibility input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.16.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.16.2 see/1, tell/1, append/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.16.3 seeing/1, telling/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.16.4 seen/0, told/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.16.5 get0/1, get/1, skip/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.16.6 put/1, tab/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

7.17 Term expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.17.1 Definite clause grammars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.17.2 expand term/2, term expansion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.17.3 phrase/3, phrase/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

7.18 Logic, control and exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

7.18.1 abort/0, stop/0, top level/0, break/0, halt/1, halt/0 . . . . . . . . . . . . . 112
7.18.2 once/1, (\+)/1 - not provable, call with args/1-11, call/2 . . . . . . . . . . . 113

7.18.3 repeat/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 7.18.4 for/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

7.19 Atomic term processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.19.1 atom length/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.19.2 atom concat/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 7.19.3 sub atom/5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 7.19.4 char code/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.19.5 lower upper/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.19.6 atom chars/2, atom codes/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

7.19.7 number atom/2, number chars/2, number codes/2 . . . . . . . . . . . . . . . . . 117

7.19.8 name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.19.9 atom hash/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 7.19.10 new atom/3, new atom/2, new atom/1 . . . . . . . . . . . . . . . . . . . . . . . . . 119 7.19.11 current atom/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 7.19.12 atom property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

7.20 List processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.20.1 append/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.20.2 member/2, memberchk/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.20.3 reverse/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 7.20.4 delete/3, select/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 7.20.5 permutation/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 7.20.6 prefix/2, suffix/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 7.20.7 sublist/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 7.20.8 last/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 7.20.9 length/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 7.20.10 nth/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 7.20.11 max list/2, min list/2, sum list/2 . . . . . . . . . . . . . . . . . . . . . . . . . 125

7.20.12 sort/2, sort0/2, keysort/2 sort/1, sort0/1, keysort/1 . . . . . . . . . . . . . 125

7.21 Global variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 7.21.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 7.21.2 g assign/2, g assignb/2, g link/2 . . . . . . . . . . . . . . . . . . . . . . . . . 127 7.21.3 g read/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 7.21.4 g array size/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

7.21.5 g inc/3, g inc/2, g inco/2, g inc/1, g dec/3, g dec/2, g deco/2, g dec/1 . . . 129
7.21.6 g set bit/2, g reset bit/2, g test set bit/2, g test reset bit/2 . . . . . . 130

7.21.7 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 7.22 Prolog state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

7.22.1 set prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.22.2 current prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.22.3 set bip name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.22.4 current bip name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

7.22.5 write pl state file/1, read pl state file/1 . . . . . . . . . . . . . . . . . . . 136

7.23 Program state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.23.1 consult/1, '.'/2 - program consult . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.23.2 load/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.23.3 listing/1, listing/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

7.24 System statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 7.24.1 statistics/0, statistics/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

7.24.2 user time/1, system time/1, cpu time/1, real time/1 . . . . . . . . . . . . . . 139

7.25 Random number generator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 7.25.1 set seed/1, randomize/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 7.25.2 get seed/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.25.3 random/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.25.4 random/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

7.26 File name processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.26.1 absolute file name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.26.2 decompose file name/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.26.3 prolog file name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

7.27 Operating system interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 7.27.1 argument counter/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 7.27.2 argument value/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.27.3 argument list/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.27.4 environ/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

7.27.5 make directory/1, delete directory/1, change directory/1 . . . . . . . . . . 144

7.27.6 working directory/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 7.27.7 directory files/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 7.27.8 rename file/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 7.27.9 delete file/1, unlink/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 7.27.10 file permission/2, file exists/1 . . . . . . . . . . . . . . . . . . . . . . . . . 146 7.27.11 file property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 7.27.12 temporary name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 7.27.13 temporary file/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 7.27.14 date time/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 7.27.15 host name/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.27.16 os version/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.27.17 architecture/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.27.18 shell/2, shell/1, shell/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.27.19 system/2, system/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 7.27.20 spawn/3, spawn/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 7.27.21 popen/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.27.22 exec/5, exec/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.27.23 fork prolog/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.27.24 create pipe/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.27.25 wait/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.27.26 prolog pid/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.27.27 send signal/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.27.28 sleep/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.27.29 select/5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

7.28 Sockets input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 7.28.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 7.28.2 socket/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 7.28.3 socket close/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 7.28.4 socket bind/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 7.28.5 socket connect/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 7.28.6 socket listen/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 7.28.7 socket accept/4, socket accept/3 . . . . . . . . . . . . . . . . . . . . . . . . . 160 7.28.8 hostname address/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

7.29 Linedit management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.29.1 get linedit prompt/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.29.2 set linedit prompt/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.29.3 add linedit completion/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.29.4 find linedit completion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

7.30 Source reader facility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 7.30.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 7.30.2 sr open/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.3 sr change options/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.4 sr close/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.5 sr read term/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.6 sr current descriptor/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.7 sr get stream/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.8 sr get module/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.9 sr get file name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.10 sr get position/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.11 sr get include list/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.12 sr get include stream list/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.13 sr get size counters/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.14 sr get error counters/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.15 sr set error counters/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.30.16 sr error from exception/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

7.30.17 sr write message/8, sr write message/6, sr write message/4 . . . . . . . . . 164
7.30.18 sr write error/6, sr write error/4, sr write error/2 . . . . . . . . . . . . . 164

8 Finite domain solver and built-in predicates 165

8.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

8.1.1 Finite Domain variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 8.2 FD variable parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

8.2.1 fd max integer/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 8.2.2 fd vector max/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 8.2.3 fd set vector max/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

8.3 Initial value constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 8.3.1 fd domain/3, fd domain bool/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 8.3.2 fd domain/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168

8.4 Type testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168

8.4.1 fd var/1, non fd var/1, generic var/1, non generic var/1 . . . . . . . . . . . 168

8.5 FD variable information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 8.5.1 fd min/2, fd max/2, fd size/2, fd dom/2 . . . . . . . . . . . . . . . . . . . . . . 169

8.5.2 fd has extra cstr/1, fd has vector/1, fd use vector/1 . . . . . . . . . . . . . 170

8.6 Arithmetic constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 8.6.1 FD arithmetic expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 8.6.2 Partial AC: (#=)/2 - constraint equal, (#\=)/2 - constraint not equal,

(#<)/2 - constraint less than, (#=<)/2 - constraint less than or equal,

(#>)/2 - constraint greater than, (#>=)/2 - constraint greater than or equal … 171 8.6.3 Full AC: (#=#)/2 - constraint equal, (#\=#)/2 - constraint not equal,

(#<#)/2 - constraint less than, (#=<#)/2 - constraint less than or equal,

(#>#)/2 - constraint greater than, (#>=#)/2 - constraint greater than or equal . . 172 8.6.4 fd prime/1, fd not prime/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

8.7 Boolean and reified constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 8.7.1 Boolean FD expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 8.7.2 (#\)/1 - constraint NOT, (#<=>)/2 - constraint equivalent,

(#\<=>)/2 - constraint different, (##)/2 - constraint XOR, (#==>)/2 - constraint imply, (#\==>)/2 - constraint not imply, (#/\)/2 - constraint AND, (#\/\)/2 - constraint NAND, (#\/)/2 - constraint OR, (#\\/)/2 - constraint NOR . . . . . . . . . . . . . . . . 174

8.7.3 fd cardinality/2, fd cardinality/3, fd at least one/1, fd at most one/1,

fd only one/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

8.8 Symbolic constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 8.8.1 fd all different/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 8.8.2 fd element/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 8.8.3 fd element var/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

8.8.4 fd atmost/3, fd atleast/3, fd exactly/3 . . . . . . . . . . . . . . . . . . . . . 177

8.8.5 fd relation/2, fd relationc/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 8.9 Labeling constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

8.9.1 fd labeling/2, fd labeling/1, fd labelingff/1 . . . . . . . . . . . . . . . . . 179

8.10 Optimization constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 8.10.1 fd minimize/2, fd maximize/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180

9 Interfacing Prolog and C 183

9.1 Calling C from Prolog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 9.1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 9.1.2 foreign/2 directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 9.1.3 The C function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 9.1.4 Input arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 9.1.5 Output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 9.1.6 Input/output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 9.1.7 Writing non-deterministic C code . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 9.1.8 Example: input and output arguments . . . . . . . . . . . . . . . . . . . . . . . . 186 9.1.9 Example: non-deterministic code . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 9.1.10 Example: input/output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . 189

9.2 Manipulating Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 9.2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 9.2.2 Managing Prolog atoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 9.2.3 Reading Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 9.2.4 Unifying Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 9.2.5 Creating Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 9.2.6 Testing the type of Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 9.2.7 Comparing Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 9.2.8 Copying Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 9.2.9 Comparing and evaluating arithmetic expressions . . . . . . . . . . . . . . . . . . 196

9.3 Raising Prolog errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 9.3.1 Managing the error context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 9.3.2 Instantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 9.3.3 Type error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 9.3.4 Domain error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 9.3.5 Existence error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 9.3.6 Permission error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 9.3.7 Representation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 9.3.8 Evaluation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 9.3.9 Resource error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 9.3.10 Syntax error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 9.3.11 System error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

9.4 Calling Prolog from C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 9.4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 9.4.2 Example: my call/1 - a call/1 clone . . . . . . . . . . . . . . . . . . . . . . . . . 200 9.4.3 Example: recovering the list of all operators . . . . . . . . . . . . . . . . . . . . . 202

9.5 Defining a new C main() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 9.5.1 Example: asking for ancestors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

References 207

Index 209

\underline{9}

1 Acknowledgements

I would like to thank the department of computing science^{2} at the university of Paris 1 for allowing me the time and freedom necessary to achieve this project.

I am grateful to the members of the Loco project^{3} at INRIA Rocquencourt^{4} for their encouragement. Their involvement in this work led to useful feedback and exchange.

I would particularly like to thank Jonathan Hodgson^{5} for the time and effort he put into the proofreading of this manual. His suggestions, both regarding ISO technical aspects as well as the language in which it was expressed, proved invaluable.

The on-line HTML version of this document was created using H^{E}V^{E}A^{6} developed by Luc Maranget who kindly devoted so much of his time extending the capabilities of H^{E}V^{E}A in order to handle such a sizeable manual.

Jean-Christophe Aude kindly improved the visual aspect of both the illustrations and the GNU Prolog web pages.

Thanks to Richard A. O’Keefe for his advice regarding the implementation of some Prolog built-in predicates and for suggesting me the in-place installation feature.

Many thanks to the following contributors:

• Alexander Diemand^{7} for his initial port to alpha/linux and more generally for his personal involvement in the development of GNU Prolog.

• Clive Cox^{8} and Edmund Grimley Evans^{9} for their port to ix86/SCO.

• Nicolas Ollinger^{10} to for his port to ix86/FreeBSD.

• Brook Milligan^{11} for his port to ix86/NetBSD and for general configuration improvements.

• Andreas Stolcke^{12} for his port to ix86/Solaris.

• Lindsey Spratt^{13} for his port to powerpc/Darwin (MacOS X).

Many thanks to all those people at GNU^{14} who helped me to finalize the GNU Prolog project.

Finally, I would like to thank everybody who tested preliminary releases and helped me to put the finishing touches to this system.

2http://panoramix.univ-paris1.fr/CRINFO/

3http://loco.inria.fr/

4http://www.inria.fr/Unites/ROCQUENCOURT-eng.html

5http://www.sju.edu/~jhodgson

6http://pauillac.inria.fr/~maranget/hevea/

7ax@apax.net

8clive@laluna.demon.co.uk

9http://www.rano.org/

10nollinge@ens-lyon.fr

11brook@nmsu.edu

12http://www.speech.sri.com/people/stolcke/

13spratt@alum.mit.edu

14http://www.gnu.org

10  1 ACKNOWLEDGEMENTS

2 Introduction

GNU Prolog is a free Prolog compiler with constraint solving over finite domains developed by Daniel Diaz^{15}. For recent information about GNU Prolog please consult the GNU Prolog page^{16}.

GNU Prolog is a Prolog compiler based on the Warren Abstract Machine (WAM) [8, 1]. It first compiles a Prolog program to a WAM file which is then translated to a low-level machine independent language called mini-assembly specifically designed for GNU Prolog. The resulting file is then translated to the assembly language of the target machine (from which an object is obtained). This allows GNU Prolog to produce a native stand alone executable from a Prolog source (similarly to what does a C compiler from a C program). The main advantage of this compilation scheme is to produce native code and to be fast. Another interesting feature is that executables are small. Indeed, the code of most unused built-in predicates is not included in the executables at link-time.

A lot of work has been devoted to the ISO compatibility. Indeed, GNU Prolog is very close to the ISO standard for Prolog^{17} [5].

GNU Prolog also offers various extensions very useful in practice (global variables, OS interface, sockets,...). In particular, GNU Prolog contains an efficient constraint solver over Finite Domains (FD). This opens contraint logic pogramming to the user combining the power of constraint programming to the declarativity of logic programming. The key feature of the GNU Prolog solver is the use of a single (lowlevel) primitive to define all (high-level) FD constraints. There are many advantages of this approach: constraints can be compiled, the user can define his own constraints (in terms of the primitive), the solver is open and extensible (as opposed to black-box solvers like CHIP),… Moreover, the GNU Prolog solver is rather efficient, often more than commercial solvers.

GNU Prolog is inspired from two systems developed by the same author:

• wamcc: a Prolog to C compiler [3]. the key point of wamcc was its ability to produce stand alone executables using an original compilation scheme: the translation of Prolog to C via the WAM. Its drawback was the time needed by gcc to compile the produced sources. GNU Prolog can also produce stand alone executables but using a faster compilation scheme.

• clp(FD): a constraint programming language over FD [4]. Its key feature was the use of a single primitive to define FD constraints. GNU Prolog is based on the same idea but offers an extended constraint definition language. In comparison to clp(FD), GNU Prolog offers new predefined constraints, new predefined heuristics, reified constraints,…

Here are some features of GNU Prolog:

• Prolog system:

– conforms to the ISO standard for Prolog (floating point numbers, streams, dynamic code,…).

– a lot of extensions: global variables, definite clause grammars (DCG), sockets interface, operating system interface,…

– more than 300 Prolog built-in predicates.

– Prolog debugger and a low-level WAM debugger.

– line editing facility under the interactive interpreter with completion on atoms.

– powerful bidirectional interface between Prolog and C.

• Compiler:

– native-code compiler producing stand alone executables.

– simple command-line compiler accepting a wide variety of files: Prolog files, C files, WAM files,… – direct generation of assembly code 15 times faster than wamcc + gcc.

15http://pauillac.inria.fr/~diaz

16http://www.gnu.org/software/prolog

17http://www.logic-programming.org/prolog std.html

– most of unused built-in predicates are not linked (to reduce the size of the executables).

– compiled predicates (native-code) as fast as wamcmcc on average.

– consulted predicates (byte-code) 5 times faster than wamcc.

• Constraint solver:

– FD variables well integrated into the Prolog environment (full compatibility with Prolog variables and integers). No need for explicit FD declarations.

– very efficient FD solver (comparable to commercial solvers).

– high-level constraints can be described in terms of simple primitives.

– a lot of predefined constraints: arithmetic constraints, boolean constraints, symbolic constraints, reified constraints,…

– several predefined enumeration heuristics.

– the user can define his own new constraints.

– more than 50 FD built-in constraints/predicates.

3 Using GNU Prolog

3.1 Introduction

GNU Prolog offers two ways to execute a Prolog program:

• interpreting it using the GNU Prolog interactive interpreter.

• compiling it to a (machine-dependent) executable using the GNU Prolog native-code compiler.

Running a program under the interactive interpreter allows the user to list it and to make full use of the debugger on it (section 4, page 29). Compiling a program to native code makes it possible to obtain a stand alone executable, with a reduced size and optimized for speed. Running a Prolog program compiled to native-code is around 3-5 times faster than running it under the interpreter. However, it is not possible to make full use of the debugger on a program compiled to native-code. Nor is it possible to list the program. In general, it is preferable to run a program under the interpreter for debugging and then use the native-code compiler to produce an autonomous executable. It is also possible to combine these two modes by producing an executable that contains some parts of the program (e.g. already debugged predicates whose execution-time speed is crucial) and interpreting the other parts under this executable. In that case, the executable has the same facilities as the GNU Prolog interpreter but also integrates the native-code predicates. This way to define a new enriched interpreter is detailed later (section 3.4.5, page 25).

3.2 The GNU Prolog interactive interpreter

3.2.1 Starting/exiting the interactive interpreter

GNU Prolog offers a classical Prolog interactive interpreter also called top-level. It allows the user to execute queries, to consult Prolog programs, to list them, to execute them and to debug them. The top-level can be invoked using the following command:

% gprolog [OPTION ]…  (the % symbol is the operating system shell prompt)

Options:

--init-goal GOAL  execute GOAL before top level/0
--entry-goal GOAL  execute GOAL inside top level/0
--query-goal GOAL  execute GOAL as a query for top level/0
--help  print a help and exit
--version  print version number and exit
--  do not parse the rest of the command-line

The main role of the gprolog command is to execute the top-level itself, i.e. to execute the built-in predicate top level/0 (section 7.18.1, page 112) which will produce something like:

     GNU Prolog 1.2.9
     By Daniel Diaz
     Copyright (C) 1999-2001 Daniel Diaz
     | ?-

The top-level is ready to execute your queries as explained in the next section.

To quit the top-level type the end-of-file key sequence (Ctl-D) or its term representation: end of file. It is also possible to use the built-in predicate halt/0 (section 7.18.1, page 112).

However, before entering the top-level itself, the command-line is processed to treat all known options (those listed above). All unrecognized arguments are collected together to form the argument list which will be available using argument value/2 (section 7.27.2, page 143) or argument list/1 (section 7.27.3, page 143). The -- option stops the parsing of the command-line, all remainding options are collected into the argument list.

Several options are provided to execute a goal before entering the interaction with the user:

• The --init-goal option executes the GOAL as soon as it is encountered (while the commnad-line is processed). GOAL is thus executed before entering top level/0.

• The --entry-goal option executes the GOAL at the entry of top level/0 just after the banner is displayed.

• The --query-goal option executes the GOAL as if the user has typed in.

The above order is thus the order in which each kind of goal (init, entry, query) is executed. If there are several goals of a same kind they are executed in the oder of appearance. Thus, all init goals are executed (in the order of appearance) before all entry goals and all entry goals are executed before all query goals.

Each GOAL is passed as a shell argument (i.e. one shell string) and should not contain a terminal dot. Example: --init-goal 'write(hello), nl' under a sh-like. To be executed, a GOAL is transformed

into a term using read term from atom(Goal, Term, [end of term(eof)]). Respecting both the syn-

tax of shell strings and of Prolog can be heavy. For instance, passing a backslash character \ can be difficult since it introduces an escape sequence both in sh and inside Prolog quoted atoms. The use of back quotes can then be useful since, by default, no escape sequence is processed inside back quotes (this behavior can be controlled using the back quotes Prolog flag (section 7.22.1, page 133)).

Since the Prolog argument list is created when the whole command-line is parsed, if a --init-goal option uses argument value/2 or argument list/1 it will obtained the original command-line arguments (i.e. including all recognized arguments).

Here is an example of using execution goal options:

     % gprolog --init-goal 'write(before), nl' --entry-goal 'write(inside), nl'
     --query-goal 'append([a,b],[c,d],X)'

will produce the following:

     before
     GNU Prolog 1.2.9
     By Daniel Diaz
     Copyright (C) 1999-2001 Daniel Diaz
     inside
     | ?- append([a,b],[c,d],X).
     X = [a,b,c,d]
     yes
     | ?-

3.2.2 The interactive interpreter read-execute-write loop

The GNU Prolog top-level is built on a classical read-execute-write loop that also allows for re-executions (when the query is not deterministic) as follows:

• display the prompt, i.e. ’| ?-’.

• read a query (i.e. a goal).

• execute the query.

• in case of success display the values of the variables of the query.

• if there are remaining alternatives (i.e. the query is not deterministic), display a ? and ask the user who can use one of the following commands: RETURN to stop the execution, ; to compute the next solution or a to compute all remaining solution.

Here is an example of execution of a query (“find the lists X and Y such that the concatenation of X and

Y is [a,b]”):
| ?- append(X,Y,[a,b,c]).
X = []

Y = [a,b,c] ?  ;  (here the user presses ; to compute another solution)

X = [a]

Y = [b,c] ?  a  (here the user presses a to compute all remaining solutions)

X = [a,b]

Y = [c]  (here the user is not asked and the next solution is computed)

X = [a,b,c]

Y = []  (here the user is not asked and the next solution is computed)
no  (no more solution)

In some cases the top-level can detect that the current solution is the last one (no more alternatives remaining). In such a case it does not display the ? symbol (and does not ask the user). Example:

| ?- (X=1 ; X=2).

X = 1 ?  ;  (here the user presses ; to compute another solution)
X = 2  (here the user is not prompted since there are no more alternatives)

yes

The user can stop the execution even if there are more alternatives by typing RETURN.

| ?- (X=1 ; X=2).

X = 1 ?  (here the user presses RETURN to stop the execution)

yes

The top-level tries to display the values of the variables of the query in a readable manner. For instance, when a variable is bound to a query variable, the name of this variable appears. When a variable is a singleton an underscore symbol is displayed ( is a generic name for a singleton variable, it is also called an anonymous variable). Other variables are bound to new brand variable names. When a query variable name X appears as the value of another query variable Y it is because X is itself not instantiated otherwise the value of X is displayed. In such a case, nothing is output for X itself (since it is a variable). Example:

| ?- X=f(A,B, ,A), A=k.

A = k  (the value of A is displayed also in f/3 for X)
X = f(k,B, ,k)  (since B is a variable which is also a part of X, B is not displayed)

| ?- functor(T,f,3), arg(1,T,X), arg(3,T,X).

T = f(X, ,X)  (the 1^{st} and 3^{rd} args are equal to X, the 2^{nd} is an anonymous variable)

| ?- read from atom('k(X,Y,X).',T).

T = k(A, ,A)  (the 1^{st} and 3^{rd} args are unified, a new variable name A is introduced)

The top-level uses variable binding predicates (section 7.5, page 55). To display the value of a variable, the top-level calls write term/3 with the following option list: [quoted(true),numbervars(false), namevars(true)] (section 7.14.6, page 96). A term of the form '$VARNAME'(Name) where Name is an atom is displayed as a variable name while a term of the form '$VAR'(N) where N is an integer is displayed as a normal compound term (such a term could be output as a variable name by write term/3). Example:

     | ?- X='$VARNAME'('Y'), Y='$VAR'(1).

X = Y  (the term '$VARNAME'('Y') is displayed as Y)
Y = '$VAR'(1)  (the term '$VAR'(1) is displayed as is)

     | ?- X=Y, Y='$VAR'(1).
     X = '$VAR'(1)
     Y = '$VAR'(1)

In the first example, X is explicitly bound to '$VARNAME'('Y') by the query so the top-level displays Y as the value of X. Y is unified with '$VAR'(1) so the top-level displays it as a normal compound term. It should be clear that X is not bound to Y (whereas it is in the second query). This behavior should be kept in mind when doing variable binding operations.

Finally, the top-level computes the user-time (section 7.24.2, page 139) taken by a query and displays it when it is significant. Example:

     | ?- retractall(p( )), assertz(p(0)),
          repeat,
             retract(p(X)),
             Y is X + 1,
             assertz(p(Y)),
             X = 1000, !.
     X = 1000
     Y = 1001

(180 ms) yes  (the query took 180ms of user time)

3.2.3 Consulting a Prolog program

The top-level allows the user to consult Prolog source files. Consulted predicates can be listed, executed and debugged (while predicates compiled to native-code cannot). For more information about the difference between a native-code predicate and a consulted predicate refer to the introduction of this section (section 3.1, page 13) and to the part devoted to the compiler (section 3.4.1, page 20).

To consult a program use the built-in predicate consult/1 (section 7.23.1, page 136). The argument of this predicate is a Prolog file name or user to specify the terminal. This allows the user to directly input the predicates from the terminal. In that case the input shall be terminated by the end-of-file key sequence (Ctl-D) or its term representation: end of file. A shorthand for consult(FILE) is [FILE]. Example:

     | ?- [user].
     {compiling user for byte code...}
     even(0).
     even(s(s(X))):-
             even(X).

(here the user presses Ctl-D to end the input)

     {user compiled, 3 lines read - 350 bytes written, 1180 ms}
     | ?- even(X).

X = 0 ?  ;  (here the user presses ; to compute another solution)
X = s(s(0)) ?  ;  (here the user presses ; to compute another solution)
X = s(s(s(s(0)))) ? (here the user presses RETURN to stop the execution)

     yes
     | ?- listing.
     even(0).
     even(s(s(A))) :-
             even(A).

When consult/1 (section 7.23.1, page 136) is invoked on a Prolog file it first runs the GNU Prolog compiler (section 3.4, page 20) as a child process to generate a temporary WAM file for byte-code. If the compilation fails a message is displayed and nothing is loaded. If the compilation succeeds, the produced file is loaded into memory using load/1 (section 7.23.2, page 137). Namely, the byte-code of each predicate is loaded. When a predicate P is loaded if there is a previous definition for P it is removed (i.e. all clauses defining P are erased). We say that P is redefined. Note that only consulted predicates can be redefined. If P is a native-code predicate, trying to redefine it will produce an error at load-time: the predicate redefinition will be ignored and the following message displayed:

     native code procedure P cannot be redefined

Finally, an existing predicate will not be removed if it is not re-loaded. This means that if a predicate P is loaded when consulting the file F , and if later the definition of P is removed from the file F , consulting F again will not remove the previously loaded definition of P from the memory.

Consulted predicates can be debugged using the Prolog debugger. Use the debugger predicate trace/0 or debug/0 (section 4.3.1, page 29) to activate the debugger.

3.2.4 Interrupting a query

Under the top-level it is possible to interrupt the execution of a query by typing the interruption key (Ctl-C). This can be used to abort a query, to stop an infinite loop, to activate the debugger,… When an interruption occurs the top-level displays the following message: Prolog interruption (h for help) ? The user can then type one of the following commands:

Command Name Description
a  abort abort the current execution. Same as abort/0 (section 7.18.1, page 112)
e exit quit the current Prolog process. Same as halt/0 (section 7.18.1, page 112)
b  break invoke a recursive top-level. Same as break/0 (section 7.18.1, page 112)
c  continue resume the execution
t  trace start the debugger using trace/0 (section 4.3.1, page 29)
d  debug start the debugger using debug/0 (section 4.3.1, page 29)
h or ?  help  display a summary of available commands

3.2.5 The line editor

The line editor (linedit) allows the user to build/update the current input line using a variety of commands. This facility is available if the linedit part of GNU Prolog has been installed. linedit is implicitly called by any built-in predicate reading from a terminal (e.g. get char/1, read/1,…). This is the case when the top-level reads a query.

Bindings: each command of linedit is activated using a key. For some commands another key is also available to invoke the command (on some terminals this other key may not work properly while the primary key always works). Here is the list of available commands:

Key  Alternate key Description
Ctl-B  ←  go to the previous character
Ctl-F  →  go to the next character
Esc-B       Ctl-←  go to the previous word
Esc-F       Ctl-→  go to the next word
Ctl-A        Home  go to the beginning of the line
Ctl-E         End  go to the end of the line
Ctl-H      Backspace delete the previous character
Ctl-D       Delete  delete the current character
Ctl-U      Ctl-Home  delete from beginning of the line to the current character
Ctl-K       Ctl-End  delete from the current character to the end of the line
Esc-L  lower case the next word
Esc-U  upper case the next word
Esc-C  capitalize the next word
Ctl-T  exchange last two characters
Ctl-V       Insert  switch on/off the insert/replace mode
Ctl-I         Tab  complete word (twice displays all possible completions)
Esc-Ctl-I     Esc-Tab  insert spaces to emulate a tabulation
Ctl-space  mark beginning of the selection
Esc-W  copy (from the begin selection mark to the current character)
Ctl-W  cut (from the begin selection mark to the current character)

Ctl-Y  paste

Ctl-P  ↑  recall previous history line
Ctl-N  ↓  recall next history line
Esc-P  recall previous history line beginning with the current prefix
Esc-N  recall next history line beginning with the current prefix
Esc-<       Page Up  recall first history line
Esc->      Page Down recall last history line
Ctl-C  generate an interrupt signal (section 3.2.4, page 17)
Ctl-D  generate an end-of-file character (at the begin of the line)
RETURN  validate a line
Esc-?  display a summary of available commands

History: when a line is entered (i.e. terminated by RETURN), linedit records it in an internal list called history. It is later possible to recall history lines using appropriate commands (e.g. Ctl-P recall the last entered line) and to modify them as needed. It is also possible to recall a history line beginning with a given prefix. For instance to recall the previous line beginning with write simply type write followed by Esc-P. Another Esc-P will recall an earlier line beginning with write,…

Completion: another important feature of linedit is its completion facility. Indeed, linedit maintains a list of known words and uses it to complete the prefix of a word. Initially this list contains all predefined atoms and the atoms corresponding to available predicates. This list is dynamically updated when a new atom appears in the system (whether read at the top-level, created with a built-in predicate, associated to a new consulted predicate,…). When the completion key (Tab) is pressed linedit acts as follows:

• use the current word as a prefix.

• collect all words of the list that begin with this prefix.

• complete the current word with the longest common part of all matching words.

• if more than one word matches emit a beep (a second Tab will display all possibilities).

Example:

| ?- argu  (here the user presses Tab to complete the word)

     | ?- argument  (linedit completes argu with argument and emits a beep)

(the user presses again Tab to see all possible completions)

argument counter  (linedit shows 3 possible completions)

     argument list
     argument value

| ?- argument  (linedit redisplays the input line)

     | ?- argument c  (to select argument counter the user presses c and Tab)
     | ?- argument counter (linedit completes with argument counter)

Finally, linedit allows the user to check that (square/curly) brackets are well balanced. For this, when a close bracket symbol, i.e. ), ] or }, is typed, linedit determines the associated open bracket, i.e. (, [ or {, and temporarily repositions the cursor on it to show the match.

3.3 Adjusting the size of Prolog stacks

GNU Prolog uses several stacks to execute a Prolog program. Each stack has a static size and cannot be dynamically increased during the execution. For each stack there is a default size but the user can define a new size by setting an environment variable. When a GNU Prolog program is run it first consults these variables and if they are not defined uses the default sizes. The following table presents each stack of GNU Prolog with its default size and the name of its associated environment variable:

Stack  Default Environment Description
name size (Kb)  variable
local  4096  LOCALSZ  control stack (environments and choice-points)
global  8192  GLOBALSZ heap (compound terms)
trail  3072  TRAILSZ  conditional bindings (bindings to undo at backtracking)
cstr  3072  CSTRSZ  finite domain constraint stack (FD variables and constraints)

If the size of a stack is too small an overflow will occur during the execution. In that case GNU Prolog emits the following error message before stopping:

     S stack overflow (size: N Kb, environment variable used: E)

where S is the name of the stack, N is the current stack size in Kb and E the name of the associated environment variable. When such a message occurs it is possible to (re)define the variable E with the new size. For instance to allocate 8192 Kb to the local stack under a Unix shell use:

     LOCALSZ=8192; export LOCALS  (under sh or bash)
     setenv LOCALSZ 8192  (under csh or tcsh)

This method allows the user to adjust the size of Prolog stacks. However, in some cases it is preferable not to allow the user to modify these sizes. For instance, when providing a stand alone executable whose behavior should be independent of the environment in which it is run. In that case the program should not consult environment variables and the programmer should be able to define new default stack sizes. The GNU Prolog compiler offers this facilities via several command-line options such as --local-size or --fixed-sizes (section 3.4.3, page 22).

Finally note that GNU Prolog stacks are virtually allocated (i.e. use virtual memory). This means that a physical memory page is allocated only when needed (i.e. when an attempt to read/write it occurs). Thus it is possible to define very large stacks. At the execution, only the needed amount of space will be physically allocated.

3.4 The GNU Prolog compiler

3.4.1 Different kinds of codes

One of the main advantages of GNU Prolog is its ability to produce stand alone executables. A Prolog program can be compiled to native code to give rise to a machine-dependent executable using the GNU Prolog compiler. However native-code predicates cannot be listed nor fully debugged. So there is an alternative to native-code compilation: byte-code compilation. By default the GNU Prolog compiler produces native-code but via a command-line option it can produce a file ready for byte-code loading. This is exactly what consult/1 does as was explained above (section 3.2.3, page 16). GNU Prolog also manages interpreted code using a Prolog interpreter written in Prolog. Obviously interpreted code is slower than byte-code but does not require the invocation of the GNU Prolog compiler. This interpreter is used each time a meta-call is needed as by call/1 (section 6.2.3, page 47). This also the case of dynamically asserted clauses. The following table summarizes these three kinds of codes:

Type  Speed  Debug ? For what
interpreted-code slow  yes  meta-call and dynamically asserted clauses
byte-code  medium  yes  consulted predicates
native-code  fast  no  compiled predicates

3.4.2 Compilation scheme

Native-code compilation: a Prolog source is compiled in several stages to produce an object file that is linked to the GNU Prolog libraries to produce an executable. The Prolog source is first compiled to obtain a WAM [8] file. For a detailed study of the WAM the interested reader can refer to “Warren’s Abstract Machine: A Tutorial Reconstruction”^{18} [1]. The WAM file is translated to a machine-independent language specifically designed for GNU Prolog. This language is close to a (universal) assembly language and is based on a very reduced instruction set. For this reason this language is called mini-assembly (MA). The mini-assembly file is then mapped to the assembly language of the target machine. This assembly file is assembled to give rise to an object file which is then linked with the GNU Prolog libraries to provide an executable. The compiler also takes into account Finite Domain constraint definition files. It translates them to C and invoke the C compiler to obtain object files. The following figure presents this compilation scheme:

18http://www.isg.sfu.ca/~hak/documents/wam.html

Obviously all intermediate stages are hidden to the user who simply invokes the compiler on his Prolog file(s) (plus other files: C,…) and obtains an executable. However, it is also possible to stop the compiler at any given stage. This can be useful, for instance, to see the WAM code produced (perhaps when learning the WAM). Finally it is possible to give any kind of file to the compiler which will insert it in the compilation chain at the stage corresponding to its type. The type of a file is determined using the suffix of its file name. The following table presents all recognized types/suffixes:

Suffix of the file  Type of the file  Handled by:
.pl, .pro  Prolog source file  pl2wam
.wam  WAM source file  wam2ma
.ma  Mini-assembly source file  ma2asm
.s  Assembly source file  the assembler
.c, .C, .CC, .cc, .cxx, .c++, .cpp C or C++ source file  the C compiler
.fd  Finite Domain constraint source file fd2c
any other suffix (.o, .a,…)  any other type (object, library,…) the linker (C linker)

Byte-code compilation: the same compiler can be used to compile a source Prolog file for byte-code. In that case the Prolog to WAM compiler is invoked using a specific option and produces a WAM for byte-code source file (suffixed .wbc) that can be later loaded using load/1 (section 7.23.2, page 137). Note that this is exactly what consult/1 (section 7.23.1, page 136) does as explained above (section 3.2.3, page 16).

3.4.3 Using the compiler

The GNU Prolog compiler is a command-line compiler similar in spirit to a Unix C compiler like gcc. To invoke the compiler use the gplc command as follows:

% gplc [OPTION ]… FILE …  (the % symbol is the operating system shell prompt)

The arguments of gplc are file names that are dispatched in the compilation scheme depending on the type determined from their suffix as was explained previously (section 3.4.2, page 20). All object files are then linked to produce an executable. Note however that GNU Prolog has no module facility (since there is not yet an ISO reference for Prolog modules) thus a predicate defined in a Prolog file is visible from any other predicate defined in any other file. GNU Prolog allows the user to split a big Prolog source into several files but does not offer any way to hide a predicate from others.

The simplest way to obtain an executable from a Prolog source file prog.pl is to use:

    % gplc prog.pl

This will produce an native executable called prog which can be executed as follows:

    % prog

However, there are several options that can be used to control the compilation:

General options:

-o FILE , --output FILE  use FILE as the name of the output file
-W, --wam-for-native  stop after producing WAM files(s)
-w, --wam-for-byte-code  stop after producing WAM for byte-code file(s) (force --no-call-c)
-M, --mini-assembly  stop after producing mini-assembly files(s)
-S, --assembly  stop after producing assembly files (s)
-F, --fd-to-c  stop after producing C files(s) from FD constraint definition file(s)
-c, --object  stop after producing object files(s)
--temp-dir PATH  use PATH as directory for temporary files
--no-del-temp  do not delete temporary files
--no-decode-hexa  do not decode hexadecimal predicate names
-v, --verbose  print executed commands
-h, --help  print a help and exit
--version  print version number and exit

Prolog to WAM compiler options:

--pl-state FILE  read FILE to set the initial Prolog state
--no-susp-warn  do not show warnings for suspicious predicates
--no-singl-warn  do not show warnings for named singleton variables
--no-redef-error  no not show errors for built-in predicate redefinitions

--foreign-only  only compile foreign/1-2 directives
--no-call-c  do not allow the use of fd tell, '$call c’,. . .

--no-inline  do not inline predicates
--no-reorder  do not reorder predicate arguments
--no-reg-opt  do not optimize registers
--min-reg-opt  minimally optimize registers
--no-opt-last-subterm  do not optimize last subterm compilation
--fast-math  use fast mathematical mode (assume integer arithmetics)
--keep-void-inst  keep void WAM instructions in the output file
--compile-msg  print a compile message
--statistics  print statistics information

WAM to mini-assembly translator options:

--comment  include comments in the output file

Mini-assembly to assembly translator options:

--comment  include comments in the output file

C compiler options:

--c-compiler FILE  use FILE as C compiler

-C OPTION  pass OPTION to the C compiler

Assembler options:

-A OPTION  pass OPTION to the assembler

Linker options:

--local-size N  set default local stack size to N Kb
--global-size N  set default global stack size to N Kb
--trail-size N  set default trail stack size to N Kb
--cstr-size N  set default constraint stack size to N Kb
--fixed-sizes  do not consult environment variables at run-time (use default sizes)
--no-top-level  do not link the top-level (force --no-debugger)
--no-debugger  do not link the Prolog/WAM debugger
--min-pl-bips  link only used Prolog built-in predicates
--min-fd-bips  link only used FD solver built-in predicates

--min-bips  shorthand for: --no-top-level --min-pl-bips --min-fd-bips
--min-size  shorthand for: --min-bips --strip

--no-fd-lib  do not look for the FD library (maintenance only)
-s, --strip  strip the executable
-L OPTION  Pass OPTION to the linker

It is possible to only give the prefix of an option if there is no ambiguity.

The name of the output file is controlled via the -o FILE option. If present the output file produced will be named FILE . If not specified, the output file name depends on the last stage reached by the compiler. If the link is not done the output file name(s) is the input file name(s) with the suffix associated to the last stage. If the link is done, the name of the executable is the name (without suffix) of the first file name encountered in the command-line. Note that if the link is not done -o has no sense in the presence of multiple input file names. For this reason, several meta characters are available for substitution in

FILE :

• %f is substitued by the whole input file name.

• %F is similar to %f but the directory part is omitted.

• %p is substitued by the whole prefix file name (omitting the suffix).

• %P is similar to %p but the directory part is omitted.

• %s is substitued by the file suffix (including the dot).

• %d is substitued by the directory part (empty if no directory is specified).

• %c is substitued by the value of an internal counter starting from 1 and auto-incremented.

By default the compiler runs in the native-code compilation scheme. To generate a WAM file for byte-code use the --wam-for-byte-code option. The resulting file can then be loaded using load/1 (section 7.23.2, page 137).

To execute the Prolog to WAM compiler in a given read environment (operator definitions, character conversion table,…) use --pl-stateFILE . The state file should be produced by write pl state file/1 (section 7.22.5, page 136).

By default the Prolog to WAM compiler inlines calls to some deterministic built-in predicates (e.g. arg/3 and functor/3). Namely a call to such a predicate will not yield a classical predicate call but a simple C function call (which is obviously faster). It is possible to avoid this using --no-inline.

Another optimization performed by the Prolog to WAM compiler is unification reordering. The arguments of a predicate are reordered to optimize unification. This can be deactivated using --no-reorder. The compiler also optimizes the unification/loading of nested compound terms. More precisely, the compiler emits optimized instructions when the last subterm of a compound term is itself a compound term (e.g. lists). This can be deactivated using --no-opt-last-subterm.

By default the Prolog to WAM compiler fully optimizes the allocation of registers to decrease both the number of instruction produced and the number of used registers. A good allocation will generate many void instructions that are removed from the produced file except if --keep-void-inst is specified. To prevent any optimization use --no-reg-opt while --min-reg-opt forces the compiler to only perform simple register optimizations.

The Prolog to WAM compiler emits an error when a control construct or a built-in predicate is redefined. This can be avoided using --no-redef-error. The compiler also emits warnings for suspicious predicate definitions like -/2 since this often corresponds to an earlier syntax error (e.g. - instead of . This can be deactivated by specifying --no-susp-warn. Finally, the compiler warns when a singleton variable has a name (i.e. not the generic anonymous name ). This can be deactivated specifying --no-singl-warn.

Predicate names are encoded with an hexadecimal representation. This is explained in more detail later (section 3.4.6, page 26). By default the error messages from the linker (e.g. multiple definitions for a given predicate, reference to an undefined predicate,…) are filtered to replace any hexadecimal representation by the real predicate name. Specifying the --no-decode-hexa prevents gplc from filtering linker output messages and hexadecimal representations are then shown.

When producing an executable it is possible to specify default stack sizes (using --STACK NAME-size) and to prevent it from consulting environment variables (using --fixed-sizes) as was explained above (section 3.3, page 19). By default the produced executable will include the top-level, the Prolog/WAM debugger and all Prolog and FD built-in predicates. It is possible to avoid linking the top-level (section 3.2, page 13) by specifying --no-top-level. In this case, at least one initialization/1 directive (section 6.1.13, page 45) should be defined. The option --no-debugger does not link the debugger. To include only used built-in predicates that are actually used the options --no-pl-bips and/or --no-fd-bips can be specified. For the smallest executable all these options should be specified. This can be abbreviated by using the shorthand option --min-bips. By default, executables are not stripped, i.e. their symbol table is not removed. This table is only useful for the C debugger (e.g. when interfacing Prolog and C). To remove the symbol table (and then to reduce the size of the final executable) use --strip. Finally --min-size is a shortcut for --min-bips and --strip, i.e. the produced executable is as small as possible.

Example: compile and link two Prolog sources prog1.pl and prog2.pl. The resulting executable will be named prog1 (since -o is not specified):

     % gplc prog1.pl prog2.pl

Example: compile the Prolog file prog.pl to study basic WAM code. The resulting file will be named

prog.wam:
     % gplc -W --no-inline --no-reorder --keep-void-inst prog.pl

Example: compile the Prolog file prog.pl and its C interface file utils.c to provide an autonomous executable called mycommand. The executable is not stripped to allow the use of the C debugger:

     % gplc -o mycommand prog.pl utils.c

Example: detail all steps to compile the Prolog file prog.pl (the resulting executable is stripped). All intermediate files are produced (prog.wam, prog.ma, prog.s, prog.o and the executable prog):

     % gplc -W prog.pl
     % gplc -M --comment prog.wam
     % gplc -S --comment prog.ma
     % gplc -c prog.s
     % gplc -o prog -s prog.o

3.4.4 Running an executable

In this section we explain what happens when running an executable produced by the GNU Prolog nativecode compiler. The default main function first starts the Prolog engine. This function collects all linked objects (issued from the compilation of Prolog files) and initializes them. The initialization of a Prolog object file consists in adding to appropriate tables new atoms, new predicates and executing its system directives. A system directive is generated by the Prolog to WAM compiler to reflect a (user) directive executed at compile-time such as op/3 (section 6.1.10, page 44). Indeed, when the compiler encounters such a directive it immediately executes it and also generates a system directive to execute it at the start of the executable. When all system directives have been executed the Prolog engine executes all initialization directives defined with initialization/1 (section 6.1.13, page 45). If several initialization directives appear in the same file they are executed in the order of appearance. If several initialization directives appear in different files the order in which they are executed is machine-dependant. However, on most machines the order will be the reverse order in which the associated files have been linked (this is not true under native win32). When all initialization directives have been executed the default main function looks for the GNU Prolog top-level. If present (i.e. it has been linked) it is called otherwise the program simply ends. Note that if the top-level is not linked and if there is no initialization directive the program is useless since it simply ends without doing any work. The default main function detects such a behavior and emits a warning message.

Example: compile an empty file prog.pl without linking the top-level and execute it:

     % gplc --no-top-level prog.pl
     % prog
     Warning: no initial goal executed
        use a directive :- initialization(Goal)
        or remove the link option --no-top-level (or --min-bips or --min-size)

3.4.5 Generating a new interactive interpreter

In this section we show how to define a new top-level extending the GNU Prolog interactive interpreter with new predicate definitions. The obtained top-level can then be considered as an enriched version of the basic GNU Prolog top-level (section 3.2, page 13). Indeed, each added predicate can be viewed as a predefined predicate just like any other built-in predicate. This can be achieved by compiling these predicates and including the top-level at link-time.

The real question is: why would we include some predicates in a new top-level instead of simply consulting them under the GNU Prolog top-level ? There are two reasons for this:

• the predicate cannot be consulted. This is the case of a predicate calling foreign code, like a predicate interfacing with C (section 9, page 183) or a predicate defining a new FD constraint.

• the performance of the predicate is crucial. Since it is compiled to native-code such a predicate will be executed very quickly. Consulting will load it as byte-code. The gain is much more noticeable if the program is run under the debugger. The included version will not be affected by the debugger while the consulted version will be several times slower. Obviously, a predicate should be included in a new top-level only when it is itself debugged since it is difficult to debug native-code.

To define a new top-level simply compile the set of desired predicates and linking them with the GNU Prolog top-level (this is the default) using gplc (section 3.4.3, page 22).

Example: let us define a new top-level called my top level including all predicates defined in prog.pl:

     % gplc -o my top level prog.pl

By the way, note that if prog.pl is an empty Prolog file the previous command will simply create a new interactive interpreter similar to the GNU Prolog top-level.

Example: as before where some predicates of prog.pl call C functions defined in utils.c:

     % gplc -o my top level prog.pl utils.c

In conclusion, defining a particular top-level is nothing else but a particular case of the native-code compilation. It is simple to do and very useful in practice.

3.4.6 The hexadecimal predicate name encoding

When the GNU Prolog compiler compiles a Prolog source to an object file it has to associate a symbol to each predicate name. However, the syntax of symbols is restricted to identifiers: string containing only letters, digits or underscore characters. On the other hand, predicate names (i.e. atoms) can contain any character with quotes if necessary (e.g. 'x+y=z' is a valid predicate name). The compiler has then to encode predicate names respecting the syntax of identifiers. To achieve this, GNU Prolog uses an hexadecimal representation where each predicate name is translated to a symbol beginning with an X followed by the hexadecimal notation of the code of each character of the name.

Example: 'x+y=z' will be encoded as X782B793D7A since 78 is the hexadecimal representation of the code of x, 2B of the code of +, etc.

Since Prolog allows the user to define several predicates with the same name but with a different arity GNU Prolog encodes predicate indicators (predicate name followed by the arity). The symbol associated to the predicate name is then followed by an underscore and by the decimal notation of the arity.

Example: 'x+y=z'/3 will be encoded as X782B793D7A 3.

So, from the mini-assembly stage, each predicate indicator is replaced by its hexadecimal encoding. The knowledge of this encoding is normally not of interest for the user, i.e. the Prolog programmer. For this reason the GNU Prolog compiler hides this encoding. When an error occurs on a predicate (undefined predicate, predicate with multiple definitions,…) the compiler has to decode the symbol associated to the predicate indicator. For this gplc filters each message emitted by the linker to locate and decode eventual predicate indicators. This filtering can be deactivated specifying --no-decode-hexa when invoking gplc (section 3.4.3, page 22).

This filter is provided as an utility that can be invoked using the hexgplc command as follows:

% hexgplc [OPTION ]… FILE … (the % symbol is the operating system shell prompt)

Options:

--encode  encoding mode (default mode is decoding)
--relax  decode also predicate names (not only predicate indicators)
--printf FORMAT  pass encoded/decoded string to C printf(3) with FORMAT
--aux-father  decode an auxiliary predicate as its father
--aux-father2  decode an auxiliary predicate as its father + auxiliary number
--cmd-line  encode/decode each argument of the command-line

 -H  same as: --cmd-line --encode
 -P  same as: --cmd-line --relax

--help  print a help and exit
--version  print version number and exit

It is possible to give a prefix of an option if there is no ambiguity.

Without arguments hexgplc runs in decoding mode reading its standard input and decoding each symbol corresponding to a predicate indicator. To use hexgplc in the encoding mode the --encode option must be specified. By default hexgplc only decodes predicate indicators, this can be relaxed using --relax to also take into account simple predicate names (the arity can be omitted). It is possible to format the output of an encoded/decoded string using --printf FORMAT in that case each string S is passed to the

C printf(3) function as printf(FORMAT,S).

Auxiliary predicates are generated by the Prolog to WAM compiler when simplifying some control constructs like ';'/2 present in the body of a clause. They are of the form '$NAME/ARITY $auxN' where NAME/ARITY is the predicate indicator of the simplified (i.e. father) predicate and N is a sequential number (a predicate can give rise to several auxiliary predicates). It is possible to force hexgplc to decode an auxiliary predicate as its father predicate indicator using --aux-father or as its father predicate indicator followed by the sequential number using --aux-father2.

If no file is specified, hexgplc processes its standard input otherwise each file is treated sequentially. Specifying the --cmd-line option informs hexgplc that each argument is not a file name but a string that must be encoded (or decoded). This is useful to encode/decode a particular string. For this reason the option -H (encode to hexadecimal) and -P (decode to Prolog) are provided as shorthand. Then, to obtain the hexadecimal representation of a predicate P use:

     % hexgplc -H P

Example:

     % hexgplc -H 'x+y=z'
     X782B793D7A

28  3 USING GNU PROLOG

4 Debugging

4.1 Introduction

The GNU Prolog debugger provides information concerning the control flow of the program. The debugger can be fully used on consulted predicates (i.e. byte-code). For native compiled code only the calls/exits are traced, no internal behavior is shown. Under the debugger it is possible to exhaustively trace the execution or to set spy-points to only debug a specific part of the program. Spy-points allow the user to indicate on which predicates the debugger has to stop to allow the user to interact with it. The debugger uses the “procedure box control flow model”, also called the Byrd Box model since it is due to Lawrence Byrd.

4.2 The procedure box model

The procedure box model of Prolog execution provides a simple way to show the control flow. This model is very popular and has been adopted in many Prolog systems (e.g. SICStus Prolog, Quintus Prolog,…). A good introduction is the chapter 8 of “Programming in Prolog” of Clocksin & Mellish [2]. The debugger executes a program step by step tracing an invocation to a predicate (call) and the return from this predicate due to either a success (exit) or a failure (fail). When a failure occurs the execution backtracks to the last predicate with an alternative clause. The predicate is then reinvoked (redo). Another source of change of the control flow is due to exceptions. When an exception is raised from a predicate (exception) by throw/1 (section 6.2.4, page 47) the control is given back to the most recent predicate that has defined a handler to recover this exception using catch/3 (section 6.2.4, page 47). The procedure box model shows these different changes in the control flow, as illustrated here:

Each arrow corresponds to a port. An arrow to the box indicates that the control is given to this predicate while an arrow from the box indicates that the control is given back from the procedure. This model visualizes the control flow through these five ports and the connections between the boxes associated to subgoals. Finally, it should be clear that a box is associated to one invocation of a given predicate. In particular, a recursive predicate will give raise to a box for each invocation of the predicate with different entries/exits in the control flow. Since this might get confusing for the user, the debugger associates to each box a unique identifier (i.e. the invocation number).

4.3 Debugging predicates

4.3.1 Running and stopping the debugger

trace/0 activates the debugger. The next invocation of a predicate will be traced.

debug/0 activates the debugger. The next invocation of a predicate on which a spy-point has been set will be traced.

It is important to understand that the information associated to the control flow is only available when the debugger is on. For efficiency reasons, when the debugger is off the information concerning the control flow (i.e. the boxes) is not retained. So, if the debugger is activated in the middle of a computation (by a call to debug/0 or trace/0 in the program or after the interrupt key sequence (Ctl-C) by choosing trace or debug), information prior to this point is not available.

debugging/0: prints onto the terminal information about the current debugging state (whether the debugger is switched on, what are the leashed ports, spy-points defined,…).

notrace/0 or nodebug/0 switches the debugger off.

wam debug/0 invokes the sub-debugger devoted to the WAM data structures (section 4.6, page 32). It can be also invoked using the W debugger command (section 4.5, page 31).

4.3.2 Leashing ports

leash(Ports) requests the debugger to prompt the user, as he creeps through the program, for every port defined in the Ports list. Each element of Ports is an atom in call, exit, redo, fail, exception. Ports can also be an atom defining a shorthand:

   • full: equivalent to [call, exit, redo, fail, exception]
   • half: equivalent to [call, redo]

• loose: equivalent to [call]

• none: equivalent to []

   • tight: equivalent to [call, redo, fail, exception]

When an unleashed port is encountered the debugger continues to show the associated goal but does not stop the execution to prompt the user.

4.3.3 Spy-points

When dealing with big sources it is not very practical to creep through the entire program. It is preferable to define a set of spy-points on interesting predicates to be prompted when the debugger reaches one of these predicates. Spy-points can be added either using spy/1 (or spypoint condition/3) or dynamically when prompted by the debugger using the + (or *) debugger command (section 4.5, page 31). The current mode of leashing does not affect spy-points in the sense that user interaction is requested on every port.

spy(PredSpec) sets a spy-point on all the predicates given by PredSpec. PredSpec defines one or several predicates and has one of the following forms:

• [PredSpec1, PredSpec2,...]: set a spy-point for each element of the list.

• Name: set a spy-point for any predicate whose name is Name (whatever the arity).

• Name/Arity: set a spy-point for the predicate whose name is Name and arity is Arity.

• Name/A1-A2: set a spy-point for the each predicate whose name is Name and arity is between A1 and A2.

It is not possible to set a spy-point on an undefined predicate.

The following predicate is used to remove one or several spy-points:

nospy(PredSpec) removes the spy-points from the specified predicates.

nospyall/0 removes all spy-points:

It is also possible to define conditional spy-points.

spypoint condition(Goal, Port, Test) sets a conditional spy-point on the predicate for Goal. When the debugger reaches a conditional spy-point it only shows the associated goal if the following conditions are verified:

• the actual goal unifies with Goal.

• the actual port unifies with Port.

• the Prolog goal Test succeeds.

4.4 Debugging messages

We here described which information is displayed by the debugger when it shows a goal. The basic format is as follows:

     S N M Port: Goal ?

S is a spy-point indicator: if there is a spy-point on the current goal the + symbol is displayed else a space is displayed. N is the invocation number. This unique number can be used to correlate the trace messages for the various ports, since it is unique for every invocation. M is an index number which represents the number of direct ancestors of the goal (i.e. the current depth of the goal). Port specifies the particular port (call, exit, fail, redo, exception). Goal is the current goal (it is then possible to inspect its current instantiation) which is displayed using write term/3 with quoted(true) and max depth(D) options (section 7.14.6, page 96). Initially D (the print depth) is set to 10 but can be redefined using the < debugger command (section 4.5, page 31). The ? symbol is displayed when the debugger is waiting a command from the user. (i.e. Port is a leashed port). If the port is unleashed, this symbol is not displayed and the debugger continues the execution displaying the next goal.

4.5 Debugger commands

When the debugger reaches a leashed port it shows the current goal followed by the ? symbol. At this point there are many commands available. Typing RETURN will creep into the program. Continuing to creep will show all the control flow. The debugger shows every port for every predicate encountered during the execution. It is possible to select the ports at which the debugger will prompt the user using the built-in predicate leash/1 (section 4.3.2, page 30). Each command is only one character long:

Command  Name  Description
RET or c  creep  single-step to the next port
l  leap  continue the execution only stopping when a goal with a spy-point

is reached

s skip skip over the entire execution of the current goal. No message will

be shown until control returns

G go to ask for an invocation number and continue the execution until a port

is reached for that invocation number

r retry try to restart the invocation of the current goal by failing until reaching the invocation of the goal. The state of execution is the same as when the goal was initially invoked (except when using side-effect predicates)

f  fail  force the current goal to fail immediately
w  write  show the current goal using write/2 (section 7.14.6, page 96)
d  display  show the current goal using display/2 (section 7.14.6, page 96)
p  print  show the current goal using print/2 (section 7.14.6, page 96)
e  exception  show the pending exception. Only applicable to an exception port
g  ancestors  show the list of ancestors of the current goal
A  alternatives  show the list of ancestors of the current goal combined with choice-

points

u unify ask for a term and unify the current goal with this term. This is convenient for getting a specific solution. Only available at a call port

. father file show the Prolog file name and the line number where the current

predicate is defined

n  no debug  switch the debugger off. Same as nodebug/0 (section 4.3.1, page 29)
=  debugging  show debugger information. Same as debugging/0 (section 4.3.1,

page 29)

+ spy this set a spy-point on the current goal. Uses spy/1 (section 4.3.3,

page 30)

- nospy this remove a spy-point on the current goal. Uses nospy/1 (section 4.3.3, page 30)

* spy conditionally ask for a term Goal, Port, Test (terminated by a dot) and set a conditional spy-point on the current predicate. Goal and the current goal must have the same predicate indicator. Uses spypoint condition/3 (section 4.3.3, page 30)

L listing list the clauses associated to the current predicate. Uses listing/1

(section 7.23.3, page 138)

a abort abort the current execution. Same as abort/0 (section 7.18.1,

page 112)

b break invoke a recursive top-level. Same as break/0 (section 7.18.1,

page 112)

@ execute goal ask for a goal and execute it < set print depth ask for an integer and set the print depth to this value (-1 for no

depth limit)

h or ? help display a summary of available commands W WAM debugger invoke the low-level WAM debugger (section 4.6, page 32)

4.6 The WAM debugger

In some cases it is interesting to have access to the WAM data structures. This sub-debugger allows the user to inspect/modify the contents of any stack or register of the WAM. The WAM debugger is invoked using the built-in predicate wam debug/0 (section 4.3.1, page 29) or the W debugger command (section 4.5, page 31). The following table presents the specific commands of the WAM debugger:

Command  Description
write A [N ]  write N terms starting at the address A using write/1 (section 7.14.6, page 96)
data A [N ]  display N words starting at the address A
modify A [N ]  display and modify N words starting at the address A
where A  display the real address corresponding to A
what RA  display what corresponds to the real address RA
deref A  display the dereferenced word starting at the address A
envir [SA ]  display the contents of the environment located at SA (or the current one)
backtrack [SA ] display the contents of the choice-point located at SA (or the current one)
backtrack all display all choice-points
quit  quit the WAM debugger
help  display a summary of available commands

In the above table the following conventions apply:

• elements between [ and ] are optional.

• N is an optional integer (defaults to 1).

• A is a WAM address, its syntax is: BANK NAME [ [ N ] ], i.e. a bank name possibly followed by an

index (defaults to 0). BANK NAME is either: – reg: WAM general register (stack pointers, continuation, ...).

– x: WAM X register (temporary variables, i.e. arguments).

– y: WAM Y register (permanent variables).

– ab: WAM X register saved in the current choice-point.

      – STACK NAME : WAM stack (STACK NAME in local, global, trail, cstr).

• SA is a WAM stack address, i.e. STACK NAME [ [ N ] ] (special case of WAM addresses).

• RA is a real address, its syntax is the syntax of C integers (in particular the notation 0x... is

recognized). It is possible to only use the first letters of a commands and bank names when there is no ambiguity. Also the square brackets [ ] enclosing the index of a bank name can be omitted. For instance the following command (showing the contents of 25 consecutive words of the global stack from the index 3): data global[3] 25 can be abbreviated as: d g 3 25.

34  4 DEBUGGING

\underline{35}

5 Format of definitions

5.1 General format

The definition of control constructs, directives and built-in predicates is presented as follows: Templates Specifies the types of the arguments and which of them shall be instantiated (mode). Types and modes are described later (section 5.2, page 35). Description Describes the behavior (in the absence of any error conditions). It is explicitly mentioned when a builtin predicate is re-executable on backtracking. Predefined operators involved in the definition are also mentioned. Errors Details the error conditions. Possible errors are detailed later (section 5.3, page 37). For directives, this part is omitted. Portability Specifies whether the definition conforms to the ISO standard or is a GNU Prolog extension.

5.2 Types and modes

The templates part defines, for each argument of the concerned built-in predicate, its mode and type. The mode specifies whether or not the argument must be instantiated when the built-in predicate is called. The mode is encoded with a symbol just before the type. Possible modes are:

• +: the argument must be instantiated.

• -: the argument must be a variable (will be instantiated if the built-in predicate succeeds).

• ?: the argument can be instantiated or a variable. The type of an argument is defined by the following table:

Type  Description

TYPE list a list whose the type of each element is TYPE TYPE1 or TYPE2 a term whose type is either TYPE1 or TYPE2

atom                 an atom

atom property an atom property (section 7.19.12, page 120)

boolean  the atom true or false

byte an integer ≥ 0 and ≤ 255 callable term an atom or a compound term character a single character atom character code an integer ≥ 1 and ≤ 255 clause a clause (fact or rule) close option a close option (section 7.10.7, page 72) compound term a compound term evaluable an arithmetic expression (section 7.6.1, page 57) fd bool evaluable a boolean FD expression (section 8.7.1, page 173) fd labeling option an FD labeling option (section 8.9.1, page 179) fd evaluable an arithmetic FD expression (section 8.6.1, page 170) fd variable an FD variable flag a Prolog flag (section 7.22.1, page 133) float a floating point number head a head of a clause (atom or compound term)

integer              an integer

in byte an integer ≥ 0 and ≤ 255 or -1 (for the end-of-file) in character a single character atom or the atom end of file (for the end-of-file) in character code an integer ≥ 1 and ≤ 255 or -1 (for the end-of-file)

io mode  an atom in: read, write or append

list the empty list [] or a non-empty list [ | ] nonvar any term that is not a variable number an integer or a floating point number operator specifier an operator specifier (section 7.14.10, page 100) os file property an operating system file property (section 7.27.11, page 147)

predicate indicator a term Name/Arity where Name is an atom and Arity an integer ≥ 0. A

callable term can be given if the strict iso Prolog flag is switched off (section 7.22.1, page 133)

predicate property a predicate property (section 7.8.2, page 65) read option a read option (section 7.14.1, page 92)

socket address  a term of the form 'AF UNIX'(A) or 'AF INET'(A,N) where A is an atom

and N an integer

socket domain  an atom in: 'AF UNIX' or 'AF INET'

source sink an atom identifying a source or a sink stream a stream-term: a term of the form '$stream'(N) where N is an integer ≥ 0 stream option a stream option (section 7.10.6, page 70) stream or alias a stream-term or an alias (atom)

stream position  a stream position: a term '$stream position'(I1, I2, I3, I4) where

I1, I2, I3 and I4 are integers

stream property a stream property (section 7.10.10, page 73)

stream seek method an atom in: bof, current or eof

term                 any term

var binding option a variable binding option (section 7.5.3, page 56) write option a write option (section 7.14.6, page 96)

\underline{5.3 Errors 37}

5.3 Errors

5.3.1 General format and error context

When an error occurs an exception of the form: error(ErrorTerm, Caller) is raised. ErrorTerm is a term specifying the error (detailed in next sections) and Caller is a term specifying the context of the error. The context is either the predicate indicator of the last invoked built-in predicate or an atom giving general context information.

Using exceptions allows the user both to recover an error using catch/3 (section 6.2.4, page 47) and to raise an error using throw/1 (section 6.2.4, page 47).

To illustrate how to write error cases, let us write a predicate my pred(X) where X must be an integer:

     my_pred(X) :-
             (   nonvar(X) ->
                 true
             ;   throw(error(instantiation_error), my_pred/1)),
             ),
             (   integer(X) ->
                 true
             ;   throw(error(type_error(integer, X), my_pred/1))
             ),
             ...

To help the user to write these error cases, a set of system predicates is provided to raise errors. These predicates are of the form '$pl err ...' and they all refer to the implicit error context. The predicates set bip name/2 (section 7.22.3, page 135) and current bip name/2 (section 7.22.4, page 135) are provided to set and recover the name and the arity associated to this context (an arity < 0 means that only the atom corresponding to the functor is significant). Using these system predicates the user could define the above predicate as follow:

     my_pred(X) :-
             set_bip_name(my_pred,1),
             (   nonvar(X) ->
                 true
             ;   '$pl_err_instantiation'
             ),
             (   integer(X) ->
                 true
             ;   '$pl_err_type'(integer, X)
             ),
             ...

The following sections detail each kind of errors (and associated system predicates).

5.3.2 Instantiation error

An instantiation error occurs when an argument or one of its components is variable while an instantiated argument was expected. ErrorTerm has the following form: instantiation error.

The system predicate '$pl err instantiation' raises this error in the current error context (section 5.3.1, page 37).

5.3.3 Type error

A type error occurs when the type of an argument or one of its components is not the expected type (but not a variable). ErrorTerm has the following form: type error(Type, Culprit) where Type is the expected type and Culprit the argument which caused the error. Type is one of:

• atom
• atomic
• boolean
• byte
• callable
• character
• compound

                              • evaluable
                              • fd bool evaluable
                              • fd evaluable
                              • fd variable
                              • float
                              • in byte
                              • in character

                                                           • integer
                                                           • list
                                                           • number
                                                           • predicate indicator
                                                           • variable

The system predicate '$pl err type'(Type, Culprit) raises this error in the current error context (section 5.3.1, page 37).

5.3.4 Domain error

A domain error occurs when the type of an argument is correct but its value is outside the expected

domain. ErrorTerm has the following form: domain error(Domain, Culprit) where Domain is the

expected domain and Culprit the argument which caused the error. Domain is one of:

• atom property
• buffering mode
• character code list
• close option
• date time
• eof action
• fd labeling option
• flag value
• format control sequence
• g array index
• io mode
• non empty list
• not less than zero

                              • operator priority
                              • operator specifier
                              • os file permission
                              • os file property
                              • os path
                              • predicate property
                              • prolog flag
                              • read option
                              • selectable item
                              • socket address
                              • socket domain
                              • source sink
                              • statistics key

                                                           • statistics value
                                                           • stream
                                                           • stream option
                                                           • stream or alias
                                                           • stream position
                                                           • stream property
                                                           • stream seek method
                                                           • stream type
                                                           • term stream or alias
                                                           • var binding option
                                                           • write option

The system predicate '$pl err domain'(Domain, Culprit) raises this error in the current error context (section 5.3.1, page 37).

5.3.5 Existence error

an existence error occurs when an object on which an operation is to be performed does not exist.

ErrorTerm has the following form: existence error(Object, Culprit) where Object is the type of

the object and Culprit the argument which caused the error. Object is one of:

   • procedure  • source sink  • stream

The system predicate '$pl err existence'(Object, Culprit) raises this error in the current error context (section 5.3.1, page 37).

5.3.6 Permission error

A permission error occurs when an attempt to perform a prohibited operation is made. ErrorTerm

has the following form: permission error(Operation, Permission, Culprit) where Operation is

the operation which caused the error, Permission the type of the tried permission and Culprit the argument which caused the error. Operation is one of:

   • access • add alias • close

                                • create • input • modify

                                                              • open
                                                              • output
                                                              • reposition

and Permission is one of:

   • binary stream
   • flag
   • operator

                                • past end of stream
                                • private procedure
                                • source sink

                                                              • static procedure
                                                              • stream
                                                              • text stream
The system predicate '$pl err permission'(Operation, Permission, Culprit) raises this error in

the current error context (section 5.3.1, page 37).

5.3.7 Representation error

A representation error occurs when an implementation limit has been breached. ErrorTerm has the following form: representation error(Limit) where Limit is the name of the reached limit. Limit is one of:

   • character
   • character code
   • in character code

                                • max arity
                                • max integer
                                • min integer

                                                              • too many variables

The errors max integer and min integer are not currently implemented.

The system predicate '$pl err representation'(Limit) raises this error in the current error context (section 5.3.1, page 37).

5.3.8 Evaluation error

An evaluation error occurs when an arithmetic expression gives rise to an exceptional value. ErrorTerm has the following form: evaluation error(Error) where Error is the name of the error. Error is one of:

   • float overflow
   • int overflow

                                • undefined • underflow

                                                              • zero divisor

The errors float overflow, int overflow, undefined and underflow are not currently implemented.

The system predicate '$pl err evaluation'(Error) raises this error in the current error context (section 5.3.1, page 37).

5.3.9 Resource error

A resource error occurs when GNU Prolog does not have enough resources. ErrorTerm has the following form: resource error(Resource) where Resource is the name of the resource. Resource is one of:

   • print object not linked  • too big fd constraint  • too many open streams

The system predicate '$pl err resource'(Resource) raises this error in the current error context (section 5.3.1, page 37).

5.3.10 Syntax error

A syntax error occurs when a sequence of character does not conform to the syntax of terms. ErrorTerm has the following form: syntax error(Error) where Error is an atom explaining the error.

The system predicate '$pl err syntax'(Error) raises this error in the current error context (section 5.3.1, page 37).

5.3.11 System error

A system error can occur at any stage. A system error is generally associated to an external component (e.g. operating system). ErrorTerm has the following form: system error(Error) where Error is an atom explaining the error. This is an extension to ISO which only defines system error without arguments.

The system predicate '$pl err system'(Error) raises this error in the current error context (section 5.3.1, page 37).

6 Prolog directives and control constructs

6.1 Prolog directives

6.1.1 Introduction

Prolog directives are annotations inserted in Prolog source files for the compiler. A Prolog directive is used to specify:

• the properties of some procedures defined in the source file.

• the format and the syntax for read-terms in the source file (using changeable Prolog flags).

• included source files.

• a goal to be executed at run-time.

6.1.2 dynamic/1

Templates

     dynamic(+predicate indicator)
     dynamic(+predicate indicator list)
     dynamic(+predicate indicator sequence)

Description

dynamic(Pred) specifies that the procedure whose predicate indicator is Pred is a dynamic procedure. This directive makes it possible to alter the definition of Pred by adding or removing clauses. For more information refer to the section about dynamic clause management (section 7.7.1, page 60).

This directive shall precede the definition of Pred in the source file.

If there is no clause for Pred in the source file, Pred exists however as an empty predicate (this means

that current predicate(Pred) succeeds).

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

ISO directive.

6.1.3 public/1

Templates

     public(+predicate indicator)
     public(+predicate indicator list)
     public(+predicate indicator sequence)

Description

public(Pred) specifies that the procedure whose predicate indicator is Pred is a public procedure. This directive makes it possible to inspect the clauses of Pred. For more information refer to the section about dynamic clause management (section 7.7.1, page 60).

This directive shall precede the definition of Pred in the source file. Since a dynamic procedure is also public. It is useless (but correct) to define a public directive for a predicate already declared as dynamic.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

GNU Prolog directive. The ISO reference does not define any directive to declare a predicate public but it does distinguish public predicates. It is worth noting that in most Prolog systems the public/1 directive is as a visibility declaration. Indeed, declaring a predicate as public makes it visible from any predicate defined in any other file (otherwise the predicate is only visible from predicates defined in the same source file as itself). When a module system is incorporated in GNU Prolog a more general visibility declaration shall be provided conforming to the ISO reference.

6.1.4 multifile/1

Templates

     multifile(+predicate indicator)
     multifile(+predicate indicator list)
     multifile(+predicate indicator sequence)

Description

multifile(Pred) is not supported by GNU Prolog. When such a directive is encountered it is simply ignored. All clauses for a given predicate must reside in a single file.

Portability

ISO directive. Not supported.

6.1.5 discontiguous/1

Templates

     discontiguous(+predicate indicator)
     discontiguous(+predicate indicator list)
     discontiguous(+predicate indicator sequence)

Description

discontiguous(Pred) specifies that the procedure whose predicate indicator is Pred is a discontiguous procedure. Namely, the clauses defining Pred are not restricted to be consecutive but can appear anywhere in the source file.

This directive shall precede the definition of Pred in the source file.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

ISO directive. The ISO reference document states that if there is no clause for Pred in the source file, Pred exists however as an empty predicate (i.e. current predicate(Pred) will succeed). This is not the case for GNU Prolog.

6.1.6 ensure linked/1

Templates

     ensure linked(+predicate indicator)
     ensure linked(+predicate indicator list)
     ensure linked(+predicate indicator sequence)

Description

ensure linked(Pred) specifies that the procedure whose predicate indicator is Pred must be included by the linker. This directive is useful when compiling to native code to force the linker to include the code of a given predicate. Indeed, if the gplc is invoked with an option to reduce the size of the executable (section 3.4.3, page 22), the linker only includes the code of predicates that are statically referenced. However, the linker cannot detect dynamically referenced predicates (used as data passed to a meta-call predicate). The use of this directive prevents it to exclude the code of such predicates.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

GNU Prolog directive.

6.1.7 built in/0, built in/1, built in fd/0, built in fd/1

Templates

     built in
     built in(+predicate indicator)
     built in(+predicate indicator list)
     built in(+predicate indicator sequence)
     built in fd
     built in fd(+predicate indicator)
     built in fd(+predicate indicator list)
     built in fd(+predicate indicator sequence)

Description

built in specifies that the procedures defined from now have the built in property (section 7.8.2, page 65).

built in(Pred) is similar to built in/0 but only affects the procedure whose predicate indicator is

Pred.

This directive shall precede the definition of Pred in the source file.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

built in fd (resp. built in fd(Pred)) is similar to built in (resp. built in(Pred)) but sets the

built in fd predicate property (section 7.8.2, page 65).

Portability

GNU Prolog directives.

6.1.8 include/1

Templates

     include(+atom)

Description include(File) specifies that the content of the Prolog source File shall be inserted. The resulting Prolog text is identical to the Prolog text obtained by replacing the directive by the content of the Prolog source File. See absolute file name/2 for information about the syntax of File (section 7.26.1, page 141). Portability ISO directive.

6.1.9 ensure loaded/1

Templates

     ensure loaded(+atom)

Description ensure loaded(File) is not supported by GNU Prolog. When such a directive is encountered it is simply ignored. Portability ISO directive. Not supported.

6.1.10 op/3 Templates

     op(+integer, +operator specifier, +atom or atom list)

Description op(Priority, OpSpecifier, Operator) alters the operator table. This directive is executed as soon as it is encountered by calling the built-in predicate op/3 (section 7.14.10, page 100). A system directive is also generated to reflect the effect of this directive at run-time (section 3.4.4, page 25). Portability ISO directive.

6.1.11 char conversion/2

Templates

     char conversion(+character, +character)

Description char conversion(InChar, OutChar) alters the character-conversion mapping. This directive is executed as soon as it is encountered by a call to the built-in predicate char conversion/2 (section 7.14.12, page 102). A system directive is also generated to reflect the effect of this directive at run-time (section 3.4.4, page 25). Portability ISO directive.

6.1.12 set prolog flag/2

Templates

     set prolog flag(+flag, +term)

Description set prolog flag(Flag, Value) sets the value of the Prolog flag Flag to Value. This directive is executed as soon as it is encountered by a call to the built-in predicate set prolog flag/2 (section 7.22.1, page 133). A system directive is also generated to reflect the effect of this directive at run-time (section 3.4.4, page 25). Portability ISO directive.

6.1.13 initialization/1

Templates

     initialization(+callable term)

Description initialization(Goal) adds Goal to the set of goal which shall be executed at run-time. A user directive is generated to execute Goal at run-time. If several initialization directives appear in the same file they are executed in the order of appearance (section 3.4.4, page 25). Portability ISO directive.

6.1.14 foreign/2, foreign/1

Templates

     foreign(+callable term, +foreign option list)
     foreign(+callable term)

Description foreign(Template, Options) defines an interface predicate whose prototype is Template according to the options given by Options. Refer to the foreign code interface for more information (section 9.1, page 183).

foreign(Template) is equivalent to foreign(Template, []).

Portability GNU Prolog directive.

6.2 Prolog control constructs

6.2.1 true/0, fail/0, !/0

Templates

     true
     fail
     !

Description true always succeeds. fail always fails (enforces backtracking). ! always succeeds and the for side-effect of removing all choice-points created since the invocation of the predicate activating it. Errors None. Portability ISO control constructs.

6.2.2 (',')/2 - conjunction, (;)/2 - disjunction, (->)/2 - if-then Templates

     ','(+callable term, +callable term)
     ;(+callable term, +callable term)
     ->(+callable term, +callable term)

Description Goal1 , Goal2 executes Goal1 and, in case of success, executes Goal2. Goal1 ; Goal2 first creates a choice-point and executes Goal1. On backtracking Goal2 is executed. Goal1 -> Goal2 first executes Goal1 and, in case of success, removes all choice-points created by Goal1 and executes Goal2. This control construct acts like an if-then (Goal1 is the test part and Goal2 the then part). Note that if Goal1 fails ->/2 fails also. ->/2 is often combined with ;/2 to define an if-then-else as follows: Goal1 -> Goal2 ; Goal3. Note that Goal1 -> Goal2 is the first argument of the (;)/2 and Goal3 (the else part) is the second argument. Such an if-then-else control construct first creates a choice-point for the else-part (intuitively associated to ;/2) and then executes Goal1. In case of success, all choice-points created by Goal1 together with the choice-point for the else-part are removed and Goal2 is executed. If Goal1 fails then Goal3 is executed.

',', ; and -> are predefined infix operators (section 7.14.10, page 100).

Errors

 Goal1 or Goal2 is a variable  instantiation error

Goal1 is neither a variable nor a callable term type error(callable, Goal1) Goal2 is neither a variable nor a callable term type error(callable, Goal2) The predicate indicator Pred of Goal1 or Goal2 does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

Portability

ISO control constructs.

6.2.3 call/1

Templates

     call(+callable term)

Description

call(Goal) executes Goal. call/1 succeeds if Goal represents a goal which is true. When Goal contains a cut symbol ! (section 6.2.1, page 46) as a subgoal, the effect of ! does not extend outside Goal.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

Portability

ISO control construct.

6.2.4 catch/3, throw/1

Templates

     catch(?callable term, ?term, ?term)
     throw(+nonvar)

catch(Goal, Catcher, Recovery) is similar to call(Goal) (section 6.2.3, page 47). If this succeeds or fails, so does the call to catch/3. If however, during the execution of Goal, there is a call to throw(Ball), the current flow of control is interrupted, and control returns to a call of catch/3 that is being executed. This can happen in one of two ways:

• implicitly, when an error condition for a built-in predicate is satisfied.

• explicitly, when the program executes a call of throw/1 because the program wishes to abandon the current processing, and instead to take an alternative action.

throw(Ball) causes the normal flow of control to be transferred back to an existing call of catch/3. When a call to throw(Ball) happens, Ball is copied and the stack is unwound back to the call to catch/3, whereupon the copy of Ball is unified with Catcher. If this unification succeeds, then catch/3 executes the goal Recovery using call/1 (section 6.2.3, page 47) in order to determine the success or failure of catch/3. Otherwise, in case the unification fails, the stack keeps unwinding, looking for an earlier invocation of catch/3. Ball may be any non-variable term.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

 Ball is a variable  instantiation error

If Ball does not unify with the Catcher argument of any call of catch/3, a system error message is displayed and throw/1 fails.

When catch/3 calls Recovery it uses call/1 (section 6.2.3, page 47), an instantiation error, a

type error or an existence error can then occur depending on Recovery.

Portability

ISO control constructs.

7 Prolog built-in predicates

7.1 Type testing

7.1.1 var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1,
       compound/1, callable/1, list/1, partial list/1, list or partial list/1

Templates

     var(?term)
     nonvar(?term)
     atom(?term)
     integer(?term)
     float(?term)
     number(?term)

                                                 atomic(?term)
                                                 compound(?term)
                                                 callable(?term)
                                                 list(?term)
                                                 partial list(?term)
                                                 list or partial list(?term)

Description

var(Term) succeeds if Term is currently uninstantiated (which therefore has not been bound to anything, except possibly another uninstantiated variable).

nonvar(Term) succeeds if Term is currently instantiated (opposite of var/1).

atom(Term) succeeds if Term is currently instantiated to an atom.

integer(Term) succeeds if Term is currently instantiated to an integer.

float(Term) succeeds if Term is currently instantiated to a floating point number.

number(Term) succeeds if Term is currently instantiated to an integer or a floating point number.

atomic(Term) succeeds if Term is currently instantiated to an atom, an integer or a floating point number.

compound(Term) succeeds if Term is currently instantiated to a compound term, i.e. a term of arity > 0 (a list or a structure).

callable(Term) succeeds if Term is currently instantiated to a callable term, i.e. an atom or a compound term.

list(Term) succeeds if Term is currently instantiated to a list, i.e. the atom [] (empty list) or a term with principal functor '.'/2 and with second argument (the tail) a list.

partial list(Term) succeeds if Term is currently instantiated to a partial list, i.e. a variable or a term whose the main functor is '.'/2 and the second argument (the tail) is a partial list.

list or partial list(Term) succeeds if Term is currently instantiated to a list or a partial list.

Errors

None.

Portability

var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1, compound/1 and callable/1 are

ISO predicates.

list/1, partial list/1 and list or partial list/1 are GNU Prolog predicates.

7.2 Term unification

7.2.1 (=)/2 - Prolog unification Templates

     =(?term, ?term)

Description Term1 = Term2 unifies Term1 and Term2. No occurs check is done, i.e. this predicate does not check if a variable is unified with a compound term containing this variable (this can lead to an infinite loop). = is a predefined infix operator (section 7.14.10, page 100). Errors None. Portability ISO predicate.

7.2.2 unify with occurs check/2

Templates

     unify with occurs check(?term, ?term)

Description

unify with occurs check(Term1, Term2) unifies Term1 and Term2. The occurs check test is done (i.e.

the unification fails if a variable is unified with a compound term containing this variable). Errors None. Portability ISO predicate.

7.2.3 (\=)/2 - not Prolog unifiable Templates

     \=(?term, ?term)

Description Term1 \= Term2 succeeds if Term1 and Term2 are not unifiable (no occurs check is done).

\= is a predefined infix operator (section 7.14.10, page 100).

Errors

None.

Portability

ISO predicate.

7.3 Term comparison

7.3.1 Standard total ordering of terms

The built-in predicates described in this section allows the user to compare Prolog terms. Prolog terms are totally ordered according to the standard total ordering of terms which is as follows (from the smallest term to the greatest):

• variables, oldest first.

• finite domain variables (section 8.1.1, page 165), oldest first.

• floating point numbers, in numeric order.

• integers, in numeric order.

• atoms, in alphabetical (i.e. character code) order.

• compound terms, ordered first by arity, then by the name of the principal functor and by the arguments in left-to-right order.

A list is treated as a compound term (whose principal functor is '.'/2).

The portability of the order of variables is not guaranteed (in the ISO reference the oder of variables is system dependent).

7.3.2 (==)/2 - term identical, (\==)/2 - term not identical, (@<)/2 - term less than, (@=<)/2 - term less than or equal to, (@>)/2 - term greater than, (@>=)/2 - term greater than or equal to

Templates

     ==(?term, ?term)
     \==(?term, ?term)
     @<(?term, ?term)

                                                 @=<(?term, ?term)
                                                 @>(?term, ?term)
                                                 @>=(?term, ?term)

Description

These predicates compare two terms according to the standard total ordering of terms (section 7.3.1, page 51).

Term1 == Term2 succeeds if Term1 and Term2 are equal.

Term1 \== Term2 succeeds if Term1 and Term2 are different.

Term1 @< Term2 succeeds if Term1 is less than Term2.

Term1 @=< Term2 succeeds if Term1 is less than or equal to Term2. Term1 @> Term2 succeeds if Term1 is greater than Term2. Term1 @>= Term2 succeeds if Term1 is greater than or equal to Term2. ==, \==, @<, @=<, @> and @>= are predefined infix operators (section 7.14.10, page 100). Errors None. Portability ISO predicates.

7.3.3 compare/3

Templates

     compare(?atom, +term, +term)

Description compare(Result, Term1, Term2) compares Term1 and Term2 according to the standard (section 7.3.1, page 51) and unifies Result with:

• the atom < if Term1 is less than Term2.

• the atom = if Term1 and Term2 are equal.

• the atom > if Term1 is greater than Term2. Errors

 Result is neither a variable nor an atom  type error(atom, Result)

Portability GNU Prolog predicate.

7.4 Term processing

7.4.1 functor/3

Templates

     functor(+nonvar, ?atomic, ?integer)
     functor(-nonvar, +atomic, +integer)

Description functor(Term, Name, Arity) succeeds if the principal functor of Term is Name and its arity is Arity. This predicate can be used in two ways:

• Term is not a variable: extract the name (an atom or a number if Term is a number) and the arity of Term (if Term is atomic Arity = 0).

• Term is a variable: unify Term with a general term whose principal functor is given by Name and

arity is given by Arity. Errors

 Term and Name are both variables  instantiation error
 Term and Arity are both variables  instantiation error

Term is a variable and Name is neither a variable nor an atomic term

                                             type error(atomic, Name)

Term is a variable and Arity is neither a variable nor an integer

                                             type error(integer, Arity)

Term is a variable, Name is a constant but not an atom and Arity is an integer > 0

                                             type error(atom, Name)

Term is a variable and Arity is an integer > max arity flag (section 7.22.1, page 133)

                                             representation error(max arity)
 Term is a variable and Arity is an integer < 0  domain error(not less than zero, Arity)

Portability ISO predicate.

7.4.2 arg/3

Templates

     arg(+integer, +compound term, ?term)

Description arg(N, Term, Arg) succeeds if the Nth argument of Term is Arg. Errors

 N is a variable  instantiation error
 Term is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N) Term is neither a variable nor a compound term type error(compound, Term)

 N is an integer < 0  domain error(not less than zero, N)

Portability ISO predicate.

7.4.3 (=..)/2 - univ Templates

     =..(+nonvar, ?list)
     =..(-nonvar, +list)

Description Term =.. List succeeds if List is a list whose head is the atom corresponding to the principal functor of Term and whose tail is a list of the arguments of Term. =.. is a predefined infix operator (section 7.14.10, page 100).

Errors

Term is a variable and List is a partial list instantiation error List is neither a partial list nor a list type error(list, List) Term is a variable and List is a list whose head is a variable

                                             instantiation error

List is a list whose head H is neither an atom nor a variable and whose tail is not the empty list

                                             type error(atom, H)

List is a list whose head H is a compound term and whose tail is the empty list

                                             type error(atomic, H)
 Term is a variable and List is the empty list  domain error(non empty list, [])

Term is a variable and the tail of List has a length > max arity flag (section 7.22.1, page 133)

                                             representation error(max arity)

Portability

ISO predicate.

7.4.4 copy term/2

Templates

     copy term(?term, ?term)

Description

copy term(Term1, Term2) succeeds if Term2 unifies with a term T which is a renamed copy of Term1.

Errors

None.

Portability

ISO predicate.

7.4.5 setarg/4, setarg/3

Templates

     setarg(+integer, +compound term, +term, +boolean)
     setarg(+integer, +compound term, +term)

Description

setarg(N, Term, NewValue, Undo) replaces destructively the Nth argument of Term with NewValue.

This assignment is undone on backtracking if Undo = true. This should only used if there is no further use of the old value of the replaced argument. If Undo = false then NewValue must be either an atom or an integer.

setarg(N, Term, NewValue) is equivalent to setarg(N, Term, NewValue, true).

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)
 Term is a variable  instantiation error

Term is neither a variable nor a compound term type error(compound, Term) NewValue is neither an atom nor an integer and

 Undo = false

                                             type error(atomic, NewValue)
 Undo is a variable  instantiation error

Undo is neither a variable nor a boolean type error(boolean, Undo) Portability GNU Prolog predicate.

7.5 Variable naming/numbering

7.5.1 name singleton vars/1

Templates

     name singleton vars(?term)

Description name singleton vars(Term) binds each singleton variable appearing in Term with a term of the form '$VARNAME'(' '). Such a term can be output by write term/3 as a variable name (section 7.14.6, page 96). Errors None. Portability GNU Prolog predicates.

7.5.2 name query vars/2

Templates

     name query vars(+list, ?list)

Description

name query vars(List, Rest) for each element of List of the form Name = Var where Name is an

atom and Var a variable, binds Var with the term '$VARNAME'(Name). Such a term can be output by write term/3 as a variable name (section 7.14.6, page 96). Rest is unified with the list of elements of List that have not given rise to a binding. This predicate is provided as a way to name the vari-

able lists obtained returned by read term/3 with variable names(List) or singletons(List) options

(section 7.14.1, page 92). Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) Rest is neither a partial list nor a list type error(list, Rest)

Portability

GNU Prolog predicate.

7.5.3 bind variables/2, numbervars/3, numbervars/1

Templates

     bind variables(?term, +var binding option list)
     numbervars(?term, +integer, ?integer)
     numbervars(?term)

Description

bind variables(Term, Options) binds each variable appearing in Term according to the options given

by Options.

Variable binding options: Options is a list of variable binding options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• numbervars: specifies that each variable appearing in Term should be bound to a term of the form '$VAR'(N) where N is an integer. Such a term can be output by write term/3 as a variable name (section 7.14.6, page 96). This is the default.

• namevars: specifies that each variables appearing in Term shall be bound to a term of the form '$VARNAME'(Name) where Name is the atom that would be output by write term/3 seeing a term of the '$VAR'(N) where N is an integer. Such a term can be output by write term/3 as a variable name (section 7.14.6, page 96). This is the alternative to numbervars.

• from(From): the first integer N to use for number/name variables of Term is From. The default value is 0.

• next(Next): when bind variables/2 succeeds, Next is unified with the (last integer N)+1 used to bind the variables of Term.

• exclude(List): collects all variable names appearing in List to avoid a clash when binding a

variable of Term. Precisely a number N ≥ From will not be used to bind a variable of Term if:

– there is a sub-term of List of the form '$VAR'(N) or '$VARNAME'(Name) where Name is the constant that would be output by write term/3 seeing a term of the '$VAR'(N).

– an element of List is of the form Name = Var where Name is an atom that would be output by write term/3 on seeing a term of the from '$VAR'(N). This case allows for lists returned

         by read term/3 (with variable names(List) or singletons(List) options) (section 7.14.1,

page 92) and by name query vars/2 (section 7.5.2, page 55).

numbervars(Term, From, Next) is equivalent to bind variables(Term, [from(From), next(Next)],

i.e. each variable of Term is bound to '$VAR'(N) where From ≤ N < Next.

numbervars(Term) is equivalent to numbervars(Term, 0,  ).

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a variable binding option

                                             domain error(var binding option, E)
 From is a variable  instantiation error

From is neither a variable nor an integer type error(integer, From) Next is neither a variable nor an integer type error(integer, Next)

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List)

Portability

GNU Prolog predicates.

7.5.4 term ref/2

Templates

     term ref(+term, ?integer)
     term ref(?term, +integer)

Description

term ref(Term, Ref) succeeds if the internal reference of Term is Ref. This predicate can be used either to obtain the internal reference of a term or to obtain the term associated to a given reference. Note that two identical terms can have different internal references. A good way to use this predicate is to first record the internal reference of a given term and to later re-obtain the term via this reference.

Errors

 Term and Ref are both variables  instantiation error

Ref is neither a variable nor an integer type error(integer, Ref)

 Ref is an integer < 0  domain error(not less than zero, Ref)

Portability

GNU Prolog predicate.

7.6 Arithmetic

7.6.1 Evaluation of an arithmetic expression

An arithmetic expression is a Prolog term built from numbers, variables, and functors (or operators) that represent arithmetic functions. When an expression is evaluated each variable must be bound to a non-variable expression. An expression evaluates to a number, which may be an integer or a floating point number. The following table details the components of an arithmetic expression, how they are evaluated, the types expected/returned and if they are ISO or an extension:

Expression  Result = eval(Expression)  Signature ISO

Variable must be bound to a non-variable expression E.

The result is eval(E)

IF → IF Y
integer number  this number  I → I  Y
floating point number  this number  F → F  Y
+ E                       eval(E)  IF → IF  N
- E                       - eval(E)  IF → IF  Y
inc(E)                    eval(E) + 1  IF → IF  N
dec(E)                    eval(E) - 1  IF → IF  N
E1 + E2  eval(E1) + eval(E2)  IF, IF → IF Y
E1 - E2  eval(E1) - eval(E2)  IF, IF → IF Y
E1 * E2  eval(E1) * eval(E2)  IF, IF → IF Y
E1 / E2  eval(E1) / eval(E2)  IF, IF → F Y
E1 // E2  rnd(eval(E1) / eval(E2))  I, I → I  Y
E1 rem E2  eval(E1) - (rnd(eval(E1) / eval(E2))*eval(E2))  I, I → I  Y
E1 mod E2  eval(E1) - ( ⌊eval(E1) / eval(E2)⌋ *eval(E2))  I, I → I  Y
E1 /\ E2  eval(E1) bitwise and eval(E2)  I, I → I  Y
E1 \/ E2  eval(E1) bitwise or eval(E2)  I, I → I  Y
E1 ^E2  eval(E1) bitwise xor eval(E2)  I, I → I  N
\ E  bitwise not eval(E)  I → I  Y
E1 << E2  eval(E1) integer shift left eval(E2)  I, I → I  Y
E1 >> E2  eval(E1) integer shift right eval(E2)  I, I → I  Y
abs(E)  absolute value of eval(E)  IF → IF  Y
sign(E)  sign of eval(E) (-1 if < 0, 0 if = 0, +1 if > 0)  IF → IF  Y
min(E1,E2)  minimal value between eval(E1) and eval(E2)  IF, IF → ? N
max(E1,E2)  maximal value between eval(E1) and eval(E2)  IF, IF → ? N
E1 ** E2  eval(E1) raised to the power of eval(E2)  IF, IF → F Y
sqrt(E)  square root of eval(E)  IF → F  Y
atan(E)  arc tangent of eval(E)  IF → F  Y
cos(E)  cosine of eval(E)  IF → F  Y
acos(E)  arc cosine of eval(E)  IF → F  N
sin(E)  sine of eval(E)  IF → F  Y
asin(E)  arc sine of eval(E)  IF → F  N
exp(E)  e raised to the power of eval(E)  IF → F  Y
log(E)  natural logarithms of eval(E)  IF → F  Y
float(E)  the floating point number equal to eval(E)  IF → F  Y
ceiling(E)  rounds eval(E) upward to the nearest integer  F → I  Y
floor(E)  rounds eval(E) downward to the nearest integer  F → I  Y
round(E)  rounds eval(E) to the nearest integer  F → I  Y
truncate(E)  the integer value of eval(E)  F → I  Y
float fractional part(E) the float equal to the fractional part of eval(E)  F → F  Y
float integer part(E)  the float equal to the integer part of eval(E)  F → F  Y

The meaning of the signature field is as follows:

• I → I: unary function, the operand must be an integer and the result is an integer.

• F → F: unary function, the operand must be a floating point number and the result is a floating point number.

• F → I: unary function, the operand must be a floating point number and the result is an integer.

• IF → F: unary function, the operand can be an integer or a floating point number and the result is a floating point number.

• IF → IF: unary function, the operand can be an integer or a floating point number and the result has the same type as the operand.

• I, I → I: binary function: each operand must be an integer and the result is an integer.

• IF, IF → IF: binary function: each operand can be an integer or a floating point number and the result is a floating point number if at least one operand is a floating point number, an integer otherwise.

• IF, IF → ?: binary function: each operand can be an integer or a floating point number and the result has the same type as the selected operand. This is used for min and max. Note that in case of equality between an integer and a floating point number the result is an integer.

is, +, -, *, //, /, rem, and mod are predefined infix operators. + and - are predefined prefix operators (section 7.14.10, page 100).

Integer division rounding function: the integer division rounding function rnd(X) rounds the floating point number X to an integer. There are two possible definitions (depending on the target machine) for this function which differ on negative numbers:

• rnd(X) = integer part of X, e.g. rnd(-1.5) = -1 (round toward 0)

• rnd(X) = ⌊X⌋, e.g. rnd(-1.5) = -2 (round toward −∞)

The definition of this function determines the precise definition of the integer division (//)/2 and of the integer remainder (rem)/2. Rounding toward zero is the most common case. In any case it is possible to test the value (toward zero or down) of the integer rounding function Prolog flag to determine which function being used (section 7.22.1, page 133).

Fast mathematical mode: in order to speed-up integer computations, the GNU Prolog compiler can generate faster code when invoked with the --fast-math option (section 3.4.3, page 22). In this mode only integer operations are allowed and a variable in an expression must be bound at evaluation time to an integer. No type checking is done.

Errors

a sub-expression E is a variable instantiation error a sub-expression E is neither a number nor an evaluable functor

                                             type error(evaluable, E)

a sub-expression E is a floating point number while an integer is expected

                                             type error(integer, E)

a sub-expression E is an integer while a floating point number is expected

                                             type error(float, E)
 a division by zero occurs  evaluation error(zero divisor)

Portability

Refer to the above table to determine which evaluable functors are ISO and which are GNU Prolog extensions. For efficiency reasons, GNU Prolog does not detect the following ISO arithmetic errors:

float overflow, int overflow, int underflow, and undefined.

7.6.2 (is)/2 - evaluate expression

Templates

     is(?nonvar, +evaluable)

Description

Result is Expression succeeds if Result can be unified with eval(Expression). Refer to the evaluation of an arithmetic expression for the definition of the eval function (section 7.6.1, page 57). is is a predefined infix operator (section 7.14.10, page 100). Errors Refer to the evaluation of an arithmetic expression for possible errors (section 7.6.1, page 57). Portability ISO predicate.

7.6.3 (=:=)/2 - arithmetic equal, (=\=)/2 - arithmetic not equal, (<)/2 - arithmetic less than, (=<)/2 - arithmetic less than or equal to, (>)/2 - arithmetic greater than, (>=)/2 - arithmetic greater than or equal to

Templates

     =:=(+evaluable, +evaluable)
     =\=(+evaluable, +evaluable)
     <(+evaluable, +evaluable)

                                                 =<(+evaluable, +evaluable)
                                                 >(+evaluable, +evaluable)
                                                 >=(+evaluable, +evaluable)

Description Expr1 =:= Expr2 succeeds if eval(Expr1) = eval(Expr2). Expr1 =\= Expr2 succeeds if eval(Expr1) ≠ eval(Expr2). Expr1 < Expr2 succeeds if eval(Expr1) < eval(Expr2). Expr1 =< Expr2 succeeds if eval(Expr1) ≤ eval(Expr2). Expr1 > Expr2 succeeds if eval(Expr1) > eval(Expr2). Expr1 >= Expr2 succeeds if eval(Expr1) ≥ eval(Expr2). Refer to the evaluation of an arithmetic expression for the definition of the eval function (section 7.6.1, page 57). =:=, =\=, <, =<, > and >= are predefined infix operators (section 7.14.10, page 100). Errors Refer to the evaluation of an arithmetic expression for possible errors (section 7.6.1, page 57). Portability ISO predicates.

7.7 Dynamic clause management

7.7.1 Introduction Static and dynamic procedures: a procedure is either dynamic or static. All built-in predicates are static. A user-defined procedure is static by default unless a dynamic/1 directive precedes its definition (section 6.1.2, page 41). Adding a clause to a non-existent procedure creates a dynamic procedure. The clauses of a dynamic procedure can be altered (e.g. using asserta/1), the clauses of a static procedure cannot be altered.

Private and public procedures: each procedure is either public or private. A dynamic procedure is always public. Each built-in predicate is private, and a static user-defined procedure is private by default unless a public/1 directive precedes its definition (section 6.1.3, page 41). If a dynamic declaration exists it is unnecessary to add a public declaration since a dynamic procedure is also public. A clause of a public procedure can be inspected (e.g. using clause/2), a clause of a private procedure cannot be inspected.

A logical database update view: any change in the database that occurs as the result of executing a goal (e.g. when a sub-goal is a call of assertz/1 or retract/1) only affects subsequent activations. The change does not affect any activation that is currently being executed. Thus the database is frozen during the execution of a goal, and the list of clauses defining a predication is fixed at the moment of its execution.

7.7.2 asserta/1, assertz/1

Templates

     asserta(+clause)
     assertz(+clause)

Description

asserta(Clause) first converts the term Clause to a clause and then adds it to the current internal database. The predicate concerned must be dynamic (section 7.7.1, page 60) or undefined and the clause is inserted before the first clause of the predicate. If the predicated is undefined it is created as a dynamic procedure.

assertz(Clause) acts like asserta/1 except that the clause is added at the end of all existing clauses of the concerned predicate.

Converting a term Clause to a clause Clause1:

• extract the head and the body of Clause: either Clause = (Head :- Body) or Clause = Head and

     Body = true.

• Head must be a callable term (or else the conversion fails).

• convert Body to a body clause (i.e. a goal) Body1.

• the converted clause Clause1 = (Head :- Body1).

Converting a term T to a goal:

• if T is a variable it is replaced by the term call(T).

• if T is a control construct (',')/2, (;)/2 or (->)/2 each argument of the control construct is recursively converted to a goal.

• if T is a callable term it remains unchanged.

• otherwise the conversion fails (T is neither a variable nor a callable term).

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head)

 Body cannot be converted to a goal  type error(callable, Body)

The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability

ISO predicates.

7.7.3 retract/1

Templates

     retract(+clause)

Description

retract(Clause) erases the first clause of the database that unifies with Clause. The concerned predicate must be a dynamic procedure (section 7.7.1, page 60). Removing all clauses of a procedure does not erase the procedure definition. To achieve this use abolish/1 (section 7.7.6, page 63). retract/1 is re-executable on backtracking.

Errors

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head) The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability

ISO predicate. In the ISO reference, the operation associated to the permission error is access while it is modify in GNU Prolog. This seems to be an error of the ISO reference since for asserta/1 (which is similar in spirit to retract/1) the operation is also modify.

7.7.4 retractall/1

Templates

     retractall(+head)

Description

retractall(Head) erases all clauses whose head unifies with Head. The concerned predicate must be a dynamic procedure (section 7.7.1, page 60). The procedure definition is not removed so that it is found by current predicate/1 (section 7.8.1, page 64). abolish/1 should be used to remove the procedure (section 7.7.6, page 63).

Errors

 Head is a variable  instantiation error
 Head is not a callable term  type error(callable, Head)

The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability GNU Prolog predicate.

7.7.5 clause/2

Templates

     clause(+head, ?callable term)

Description clause(Head, Body) succeeds if there exists a clause in the database that unifies with Head :- Body. The predicate in question must be a public procedure (section 7.7.1, page 60). Clauses are delivered from the first to the last. This predicate is re-executable on backtracking. Errors

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head) The predicate indicator Pred of Head is that of a private procedure

                                             permission error(access,
                                             private procedure, Pred)

Body is neither a variable nor a callable term type error(callable, Body) Portability ISO predicate.

7.7.6 abolish/1

Templates

     abolish(+predicate indicator)

Description abolish(Pred) removes from the database the procedure whose predicate indicator is Pred. The concerned predicate must be a dynamic procedure (section 7.7.1, page 60). Errors

Pred is a variable  instantiation error
Pred is a term Name/Arity and either Name or

Arity is a variable

                                           instantiation error

Pred is neither a variable nor a predicate indicator

                                           type error(predicate indicator, Pred)
Pred is a term Name/Arity and Arity is neither

a variable nor an integer

                                           type error(integer, Arity)

Pred is a term Name/Arity and Name is neither a variable nor an atom

                                           type error(atom, Name)
Pred is a term Name/Arity and Arity is an

integer < 0

                                           domain error(not less than zero, Arity)
Pred is a term Name/Arity and Arity is an

integer > max arity flag (section 7.22.1, page 133)

                                           representation error(max arity)

The predicate indicator Pred is that of a static procedure

                                           permission error(modify,
                                           static procedure, Pred)

Portability

ISO predicate.

7.8 Predicate information

7.8.1 current predicate/1

Templates

    current predicate(?predicate indicator)

Description

current predicate(Pred) succeeds if there exists a predicate indicator of a defined procedure that unifies with Pred. All user defined procedures are found, whether static or dynamic. Internal system procedures whose name begins with '$' are not found. A user-defined procedure is found even when it has no clauses. A user-defined procedure is not found if it has been abolished. To conform to the ISO reference, built-in predicates are not found except if the strict iso Prolog flag is switched off (section 7.22.1, page 133). This predicate is re-executable on backtracking.

Errors

Pred is neither a variable nor a predicate indicator

                                           type error(predicate indicator, Pred)
Pred is a term Name/Arity and Arity is neither

a variable nor an integer

                                           type error(integer, Arity)

Pred is a term Name/Arity and Name is neither a variable nor an atom

                                           type error(atom, Name)
Pred is a term Name/Arity and Arity is an

integer < 0

                                           domain error(not less than zero, Arity)
Pred is a term Name/Arity and Arity is an

integer > max arity flag (section 7.22.1, page 133)

                                           representation error(max arity)

Portability

ISO predicate.

7.8.2 predicate property/2

Templates

    predicate property(?predicate indicator, ?predicate property)

Description predicate property(Pred, Property) succeeds if there exists a predicate indicator of a defined procedure that unifies with Pred and if Property unifies with one of the properties of the procedure. All user defined procedures and built-in predicates are found. Internal system procedures whose name begins with '$' are not found. This predicate is re-executable on backtracking. Predicate properties:

• static: if the procedure is static.

• dynamic: if the procedure is dynamic.

• private: if the procedure is private.

• public: if the procedure is public.

• user: if the procedure is a user-defined procedure.

• built in: if the procedure is a Prolog built-in predicate.

• built in fd: if the procedure is an FD built-in predicate.

• native code: if the procedure is compiled in native code.

• prolog file(File): source file from which the predicate has been read.

• prolog line(Line): line number of the source file. Errors Pred is neither a variable nor a predicate indicator

                                           type error(predicate indicator, Pred)
Pred is a term Name/Arity and Arity is neither

a variable nor an integer

                                           type error(integer, Arity)

Pred is a term Name/Arity and Name is neither a variable nor an atom

                                           type error(atom, Name)
Pred is a term Name/Arity and Arity is an

integer < 0

                                           domain error(not less than zero, Arity)
Pred is a term Name/Arity and Arity is an

integer > max arity flag (section 7.22.1, page 133)

                                           representation error(max arity)

Property is neither a variable nor a predicate property term

                                           domain error(predicate property,
                                           Property)
Property = prolog file(File) and File is

neither a variable nor an atom

                                           type error(atom, File)
Property = prolog line(Line) and Line is

neither a variable nor an integer

                                           type error(integer, Line)

Portability GNU Prolog predicate.

7.9 All solutions

7.9.1 Introduction

It is sometimes useful to collect all solutions for a goal. This can be done by repeatedly backtracking and gradually building the list of solutions. The following built-in predicates are provided to automate this process.

The built-in predicates described in this section invoke call/1 (section 6.2.3, page 47) on the argument Goal. When efficiency is crucial and Goal is complex it is better to define an auxiliary predicate which can then be compiled, and have Goal call this predicate.

7.9.2 findall/3

Templates

     findall(?term, +callable term, ?list)

Description

findall(Template, Goal, Instances) succeeds if Instances unifies with the list of values to which a variable X not occurring in Template or Goal would be instantiated by successive re-executions of call(Goal), X = Template after systematic replacement of all variables in X by new variables. Thus, the order of the list Instances corresponds to the order in which the proofs are found.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

 Instances is neither a partial list nor a list  type error(list, Instances)

Portability

ISO predicate.

7.9.3 bagof/3, setof/3

Templates

     bagof(?term, +callable term, ?list)
     setof(?term, +callable term, ?list)

Description

bagof(Template, Goal, Instances) assembles as a list the set of solutions of Goal for each different instantiation of the free variables in Goal. The elements of each list are in order of solution, but the order in which each list is found is undefined. This predicate is re-executable on backtracking.

\underline{7.10 Streams 67}

Free variable set: bagof/3 groups the solutions of Goal according to the free variables in Goal. This set corresponds to all variables occurring in Goal but not in Template. It is sometimes useful to exclude some additional variables of Goal. For that, bagof/3 recognizes a goal of the form T^Goal and exclude all variables occuring in T from the free variable set. (^)/2 can be viewed as an existential quantifier (the logical reading of X^Goal being “there exists an X such that Goal is true”). The use of this existential qualifier is superfluous outside bagof/3 (and setof/3) and then is not recognized.

(^)/2 is a predefined infix operator (section 7.14.10, page 100).

setof(Template, Goal, Instances) is equivalent to bagof(Template,Goal,I), sort(I,Instances).

Each list is then a sorted list (duplicate elements are removed).

From the implementation point of view setof/3 is as fast as bagof/3. Both predicates use an in-place (i.e. destructive) sort (section 7.20.12, page 125) and require the same amount of memory.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

 Instances is neither a partial list nor a list  type error(list, Instances)

Portability

ISO predicates.

7.10 Streams

7.10.1 Introduction

A stream provides a logical view of a source/sink.

Sources and sinks: a program can output results to a sink or input data from a source. A source/sink may be a file (regular file, terminal, device,…), a constant term, a pipe, a socket,…

Associating a stream to a source/sink: to manipulate a source/sink it must be associated to a stream. This provides a logical and uniform view of the source/sink whatever its type. Once this association has been established, i.e. a stream has been created, all subsequent references to the source/sink are made by referring the stream. A stream is unidirectional: it is either an input stream or an output stream. For a classical file, the association is done by opening the file (whose name is specified as an atom) with the open/4 (section 7.10.6, page 70). GNU Prolog makes it possible to treat a Prolog constant term as a source/sink and provides built-in predicates to associate a stream to such a term (section 7.11, page 82). GNU Prolog provides operating system interface predicates defining pipes between GNU Prolog and child processes with streams associated to these pipes, e.g. popen/3 (section 7.27.21, page 153). Similarly, socket interface predicates associate streams to a socket to allow the communication, e.g. socket connect/4 (section 7.28.5, page 159).

Stream-term: a stream-term identifies a stream during a call of an input/output built-in predicate. It is created as a result of associating a stream to a source/sink (section above). A stream-term is a compound term of the form '$stream'(I) where I is an integer.

Stream aliases: any stream may be associated with a stream alias which is an atom which may be used to refer to that stream. The association can be done at open time or using add stream alias/2 (section 7.10.20, page 79). Such an association automatically ends when the stream is closed. A particular alias only refers to at most one stream at any one time. However, more than one alias can be associated to a stream. Most built-in predicates which have a stream-term as an input argument also accept a stream alias as that argument. However, built-in predicates which return a stream-term do not accept a stream alias.

Standard streams: two streams are predefined and open during the execution of every goal: the standard input stream which has the alias user input and the standard output stream which has the alias user output. A goal which attempts to close either standard stream succeeds, but does not close the stream.

Current streams: during execution there is a current input stream and a current output stream. By default, the current input and output streams are the standard input and output streams, but the builtin predicates set input/1 (section 7.10.4, page 69) and set output/1 (section 7.10.5, page 70) can be used to change them. When the current input stream is closed, the standard input stream becomes the current input stream. When the current output stream is closed, the standard output stream becomes the current output stream.

Text streams and binary streams: a text stream is a sequence of characters. A text stream is also regarded as a sequence of lines where each line is a possibly empty sequence of characters followed by a new line character. GNU Prolog may add or remove space characters at the ends of lines in order to conform to the conventions for representing text streams in the operating system. A binary stream is a sequence of bytes. Only a few built-in predicates can deal with binary streams, e.g. get byte/2 (section 7.13, page 89).

Stream positions: the stream position of a stream identifies an absolute position of the source/sink to which the stream is connected and defines where in the source/sink the next input or output will take place. A stream position is a ground term of the form '$stream position'(I1, I2, I3, I4) where I1, I2, I3 and I4 are integers. Stream positions are used to reposition a stream (when possible) using for instance set stream position/2 (section 7.10.13, page 75).

The position end of stream: when all data of a stream S has been input S has a stream position end-of-stream. At this stream position a goal to input more data will return a specific value to indicate that end of stream has been reached (e.g. -1 for get code/2 or end of file for get char/2,…). When this terminating value has been input, the stream has a stream position past-end-of-stream.

Buffering mode: input/output on a stream can be buffered (line-buffered or block-buffered) or not buffered at all. The buffering mode can be specified at open time or using set stream buffering/2 (section 7.10.27, page 82). Line buffering is used on output streams, output data are only written to the sink when a new-line character is output (or at the close time). Block buffering is used on input or output. On input streams, when an input is requested on the source, if the buffer is empty, all available characters are read (within the limits of the size of the buffer), subsequent reads will first use the characters in the buffer. On output streams, output data are stored in the buffer and only when the buffer is full is it physically written on the sink. Thus, an output to a buffered stream may not be sent immediately to the sink connected to that stream. When it is necessary to be certain that output has been delivered, the built-in predicate flush output/1 (section 7.10.8, page 72) should be used. Finally, it is also possible to use non-buffered streams, in that case input/output are directly done on the connected source/sink. This can be useful for communication purposes (e.g. sockets) or when a precise control is needed, e.g. select/5 (section 7.27.25, page 155).

Stream mirrors: any stream may be associated with mirror streams specified at open time or using add stream mirror/2 (section 7.10.22, page 79). Then, all characters/bytes read from/written to the stream are also written on each mirror stream. The association automatically ends when either the stream or the mirror stream is closed. It is also possible to explicitely remove a mirror stream using remove stream mirror/2 (section 7.10.23, page 80).

\underline{7.10 Streams 69}

7.10.2 current input/1

Templates

     current input(?stream)

Description current input(Stream) unifies Stream with the stream-term identifying the current input stream. Errors

 Stream is neither a variable nor a stream  domain error(stream, Stream)

Portability ISO predicate.

7.10.3 current output/1

Templates

     current output(?stream)

Description current output(Stream) unifies Stream with the stream-term identifying the current output stream. Errors

 Stream is neither a variable nor a stream  domain error(stream, Stream)

Portability ISO predicate.

7.10.4 set input/1

Templates

     set input(+stream or alias)

Description set input(SorA) sets the current input stream to be the stream associated with the stream-term or alias

SorA.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

Portability ISO predicate.

7.10.5 set output/1

Templates

     set output(+stream or alias)

Description

set output(SorA) sets the current output stream to be the stream associated with the stream-term or alias SorA.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

Portability

ISO predicate.

7.10.6 open/4, open/3

Templates

     open(+source sink, +io mode, -stream, +stream option list)
     open(+source sink, +io mode, -stream)

Description

open(SourceSink, Mode, Stream, Options) opens the source/sink SourceSink for input or output as

indicated by Mode and the list of stream-options Options and unifies Stream with the stream-term which is associated to this stream. See absolute file name/2 for information about the syntax of SourceSink (section 7.26.1, page 141).

Input/output modes: Mode is an atom which defines the input/output operations that may be performed the stream. Possible modes are:

• read: the source/sink is a source and must already exist. Input starts at the beginning of the source.

• write: the source/sink is a sink. If the sink already exists then it is emptied else an empty sink is created. Output starts at the beginning of that sink.

• append: the source/sink is a sink. If the sink does not exist it is created. Output starts at the end of that sink.

Stream options: Options is a list of stream options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• type(text/binary): specifies whether the stream is a text stream or a binary stream. The default value is text.

• reposition(true/false): specifies whether it is possible to reposition the stream. The default value is true except if the stream cannot be repositioned (e.g. a terminal).

• eof action(error/eof code/reset): specifies the effect of attempting to input from a stream

whose stream position is past-end-of-stream:

– error: a permission error is raised signifying that no more input exists in this stream.

– eof code: the result of input is as if the stream position is end-of-stream.

– reset: the stream position is reset so that it is not past-end-of-stream, and another attempt

is made to input from it (this is useful when inputting from a terminal). The default value is eof code.

• alias(Alias): specifies that the atom Alias is to be an alias for the stream. By default no alias is attached to the stream. Several aliases can be defined for a same stream.

• mirror(Mirror): specifies the stream associated with the stream-term or alias Mirror is a mirror for the stream. By default no mirro is attached to the stream. Several mirrors can be defined for a same stream.

• buffering(none/line/block): specifies which type of buffering is used by input/output operations

on this stream:

– none: no buffering.

– line: output operations buffer data emitted until a new-line occurs

– block: input/output operations buffer data until a given number (implementation dependant)

of characters/bytes have been treated. The default value is line for a terminal (TTY), block otherwise.

open(SourceSink, Mode, Stream, Options) is equivalent to open(SourceSink, Mode, Stream, []).

Errors

SourceSink is a variable  instantiation error
Mode is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                           instantiation error

Mode is neither a variable nor an atom type error(atom, Mode)

Options is neither a partial list nor a list  type error(list, Options)
Stream is not a variable  type error(variable, Stream)

SourceSink is neither a variable nor a source/sink

                                           domain error(source sink, SourceSink)

Mode is an atom but not an input/output mode domain error(io mode, Mode) an element E of the Options list is neither a variable nor a stream-option

                                           domain error(stream option, E)

the source/sink specified by SourceSink does not exist

                                           existence error(source sink, SourceSink)

the source/sink specified by SourceSink cannot be opened

                                           permission error(open, source sink,
                                           SourceSink)

an element E of the Options list is alias(A) and A is already associated with an open stream

                                           permission error(open, source sink,
                                           alias(A))

an element E of the Options list is mirror(M) and M is not associated with an open stream

                                           existence error(stream, M)

an element E of the Options list is mirror(M) and M iis an input stream

                                           permission error(output, stream, M)

an element E of the Options list is reposition(true) and it is not possible to reposition this stream

                                           permission error(open, source sink,
                                           reposition(true))

Portability

ISO predicates. The mirror/1 and buffering/1 stream options are GNU Prolog extensions.

7.10.7 close/2, close/1

Templates

     close(+stream or alias, +close option list)
     close(+stream or alias)

Description

close(SorA, Options) closes the stream associated with the stream-term or alias SorA. If SorA is the standard input stream or the standard output stream close/2 simply succeeds else the associated source/sink is physically closed. If SorA is the current input stream the current input stream becomes the standard input stream user input. If SorA is the current output stream the current output stream becomes the standard output stream user output.

Close options: Options is a list of close options. For the moment only one option is available:

• force(true/false): with false, if an error occurs when trying to close the source/sink, the stream is not closed and an error (system error or resource error) is raised (but close/2 succeeds). With true, if an error occurs it is ignored and the stream is closed. The purpose of force/1 option is to allow an error handling routine to do its best to reclaim resources. The default value is false.

close(SorA) is equivalent to close(SorA, []).

Errors

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

an element E of the Options list is neither a variable nor a close-option

                                             domain error(close option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)

SorA needs a special close (section 7.11, page 82) system error(needs special close)

Portability

ISO predicates. The system error(needs special close) is a GNU Prolog extension.

7.10.8 flush output/1, flush output/0

Templates

     flush output(+stream or alias)
     flush output

Description

flush output(SorA) sends any buffered output characters/bytes to the stream.

flush output/0 applies to the current output stream.

\underline{7.10 Streams 73}

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

Portability ISO predicates.

7.10.9 current stream/1

Templates

     current stream(?stream)

Description current stream(Stream) succeeds if there exists a stream-term that unifies with Stream. This predicate is re-executable on backtracking. Errors Stream is neither a variable nor a stream-term domain error(stream, Stream) Portability GNU Prolog predicate.

7.10.10 stream property/2

Templates

     stream property(?stream, ?stream property)

Description

stream property(Stream, Property) succeeds if current stream(Stream) succeeds (section 7.10.9,

page 73) and if Property unifies with one of the properties of the stream. This predicate is re-executable on backtracking. Stream properties:

• file name(F): the name of the connected source/sink.

   • mode(M): M is the open mode (read, write, append).

• input: if it is an input stream.

• output: if it is an output stream.

• alias(A): A is an alias of the stream.

• mirror(M): M is a mirror stream of the stream.

• type(T): T is the type of the stream (text, binary).

• reposition(R): R is the reposition boolean (true, false).

   • eof action(A): A is the end-of-file action (error, eof code, reset).
   • buffering(B): B is the buffering mode (none, line, block).

• end of stream(E): E is the current end-of-stream status (not, at, past). If the stream position is end-of-stream then E is unified with at else if the stream position is past-end-of-stream then E is unified with past else E is unified with not.

• position(P): P is the stream-position term associated to the current position.

Errors

 Stream is a variable  instantiation error

Stream is neither a variable nor a stream-term domain error(stream, Stream) Property is neither a variable nor a stream property

                                             domain error(stream property, Property)
 Property = file name(E), mode(E), alias(E),
 end of stream(E), eof action(E),
 reposition(E), type(E) or buffering(E) and

E is neither a variable nor an atom

                                             type error(atom, E)

Portability

ISO predicate. The buffering/1 property is a GNU Prolog extension.

7.10.11 at end of stream/1, at end of stream/0

Templates

     at end of stream(+stream or alias)
     at end of stream

Description

at end of stream(SorA) succeeds if the stream associated with stream-term or alias SorA has a stream position end-of-stream or past-end-of-stream. This predicate can be defined using stream property/2 (section 7.10.10, page 73).

at end of stream/0 applies to the current input stream.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

Portability

ISO predicates. The permission error(input, stream, SorA) is a GNU Prolog extension.

7.10.12 stream position/2

     stream position(+stream or alias, ?stream position)

Description

stream position(SorA, Position) succeeds unifying Position with the stream-position term associated to the current position of the stream-term or alias SorA. This predicate can be defined using stream property/2 (section 7.10.10, page 73).

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Position is neither a variable nor a stream-position term

                                             domain error(stream position, Position)
 SorA is not associated with an open stream  existence error(stream, SorA)

Portability

GNU Prolog predicate.

7.10.13 set stream position/2

Templates

     set stream position(+stream or alias, +stream position)

Description

set stream position(SorA, Position) sets the position of the stream associated with the streamterm or alias SorA to Position. Position should have previously been returned by stream property/2 (section 7.10.10, page 73) or by stream position/2 (section 7.10.12, page 74).

Errors

 SorA is a variable  instantiation error
 Position is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Position is neither a variable nor a stream-position term

                                             domain error(stream position, Position)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA has stream property reposition(false)    permission error(reposition, stream,
                                             SorA)

Portability

ISO predicate.

7.10.14 seek/4

Templates

     seek(+stream or alias, +stream seek method, +integer, ?integer)

Description

seek(SorA, Whence, Offset, NewOffset) sets the position of the stream associated with the streamterm or alias SorA to Offset according to Whence and unifies NewOffset with the new offset from the beginning of the file. seek/4 can only be used on binary streams. Whence is an atom from:

• bof: the position is set relatively to the begin of the file (Offset should be ≥ 0).

• current: the position is set relatively to the current position (Offset can be ≥ 0 or ≤ 0).

• eof: the position is set relatively to the end of the file (Offset should be ≤ 0).

This predicate is an interface to the C Unix function lseek(2).

Errors

SorA is a variable  instantiation error
Whence is a variable  instantiation error
Offset is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
Whence is neither a variable nor an atom  type error(atom, Whence)

Whence is an atom but not a valid stream seek method

                                           domain error(stream seek method, Whence)
Offset is neither a variable nor an integer  type error(integer, Offset)
NewOffset is neither a variable nor an integer  type error(integer, NewOffset)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA has stream property reposition(false)    permission error(reposition, stream,
                                           SorA)
SorA is associated with a text stream  permission error(reposition, text stream,
                                           SorA)

Portability

GNU Prolog predicate.

7.10.15 character count/2

Templates

    character count(+stream or alias, ?integer)

Description

character count(SorA, Count) unifies Count with the number of characters/bytes read/written on the stream associated with stream-term or alias SorA.

Errors

SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)

Portability \underline{7.10 Streams 77}

7.10.16 line count/2

Templates

     line count(+stream or alias, ?integer)

Description line count(SorA, Count) unifies Count with the number of lines read/written on the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Errors

 SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is associated with a binary stream  permission error(access, binary stream,
                                             SorA)

Portability GNU Prolog predicate.

7.10.17 line position/2

Templates

     line position(+stream or alias, ?integer)

Description line position(SorA, Count) unifies Count with the number of characters read/written on the current line of the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Errors

 SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is associated with a binary stream  permission error(access, binary stream,
                                             SorA)

Portability GNU Prolog predicate.

7.10.18 stream line column/3

Templates

     stream line column(+stream or alias, ?integer, ?integer)

Description

stream line column(SorA, Line, Column) unifies Line (resp. Column) with the current line number

(resp. column number) of the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Note that Line corresponds to the value returned by line count/2 + 1 (section 7.10.16, page 77) and Column to the value returned by line position/2 + 1 (section 7.10.17, page 77).

Errors

 SorA is a variable  instantiation error

Line is neither a variable nor an integer type error(integer, Line)

 Column is neither a variable nor an integer  type error(integer, Column)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is associated with a binary stream  permission error(access, binary stream,
                                             SorA)

Portability

GNU Prolog predicate.

7.10.19 set stream line column/3

Templates

     set stream line column(+stream or alias, +integer, +integer)

Description

set stream line column(SorA, Line, Column) sets the stream position of the stream associated with the stream-term or alias SorA according to the line number Line and the column number Column. This predicate can only be used on text streams. It first repositions the stream to the beginning of the file and then reads character by character until the required position is reached.

Errors

 SorA is a variable  instantiation error
 Line is a variable  instantiation error
 Column is a variable  instantiation error

Line is neither a variable nor an integer type error(integer, Line)

 Column is neither a variable nor an integer  type error(integer, Column)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is associated with a binary stream  permission error(reposition,
                                             binary stream, SorA)
 SorA has stream property reposition(false)    permission error(reposition, stream,
                                             SorA)

Portability \underline{7.10 Streams 79}

7.10.20 add stream alias/2

Templates

     add stream alias(+stream or alias, +atom)

Description add stream alias(SorA, Alias) adds Alias as a new alias to the stream associated with the streamterm or alias SorA. Errors

 SorA is a variable  instantiation error
 Alias is a variable  instantiation error

Alias is neither a variable nor an atom type error(atom, Alias) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 Alias is already associated with an open stream permission error(add alias, source sink,
                                             alias(Alias))

Portability GNU Prolog predicate.

7.10.21 current alias/2

Templates

     current alias(?stream, ?atom)

Description

current alias(Stream, Alias) succeeds if current stream(Stream) succeeds (section 7.10.9, page 73)

and if Alias unifies with one of the aliases of the stream. It can be defined using stream property/2 (section 7.10.10, page 73). This predicate is re-executable on backtracking. Errors Stream is neither a variable nor a stream-term domain error(stream, Stream) Alias is neither a variable nor an atom type error(atom, Alias) Portability GNU Prolog predicate.

7.10.22 add stream mirror/2

Templates

     add stream mirror(+stream or alias, +stream or alias)

Description add stream mirror(SorA, Mirror) adds the stream associated with the stream-term or alias Mirror as a new mirror to the stream associated with the stream-term or alias SorA. After this, all characters (or bytes) read from (or written to) SorA are also written to Mirror. This mirroring occurs until Mirror is explicitely removed using remove stream mirror/2 (section 7.10.23, page 80) or implicitely when Mirror is closed. Several mirror streams can be associated to a same stream. If Mirror represents the same stream as SorA or if Mirror is already a mirror for SorA, no mirror is added.

Errors

 SorA is a variable  instantiation error
 Mirror is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Mirror is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, Mirror)
 SorA is not associated with an open stream  existence error(stream, SorA)
 Mirror is not associated with an open stream  existence error(stream, Mirror)
 Mirror is an input stream  permission error(output, stream, Mirror)

Portability

GNU Prolog predicate.

7.10.23 remove stream mirror/2

Templates

     remove stream mirror(+stream or alias, +stream or alias)

Description

remove stream mirror(SorA, Mirror) removes the stream associated with the stream-term or alias Mirror from the list of mirrors of the stream associated with the stream-term or alias SorA. This predicate fails if Mirror is not a mirror stream for SorA.

Errors

 SorA is a variable  instantiation error
 Mirror is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Mirror is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, Mirror)
 SorA is not associated with an open stream  existence error(stream, SorA)
 Mirror is not associated with an open stream  existence error(stream, Mirror)

Portability

GNU Prolog predicate.

7.10.24 current mirror/2

Templates

     current mirror(?stream, ?stream)

\underline{7.10 Streams 81}

Description

current mirror(Stream, M) succeeds if current stream(Stream) succeeds (section 7.10.9, page 73)

and if M unifies with one of the mirrors of the stream. It can be defined using stream property/2 (section 7.10.10, page 73). This predicate is re-executable on backtracking. Errors Stream is neither a variable nor a stream-term domain error(stream, Stream) M is neither a variable nor a stream-term domain error(stream, M) Portability GNU Prolog predicate.

7.10.25 set stream type/2

Templates

     set stream type(+stream or alias, +atom)

Description set stream type(SorA, Type) updates the type associated with stream-term or alias SorA. The value of Type is an atom in text or binary as for open/4 (section 7.10.6, page 70). The type of a stream can only be changed before any input/output operation is executed. Errors

 SorA is a variable  instantiation error
 Type is a variable  instantiation error
 Type is neither a variable nor a valid type  domain error(stream type, Type)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

An I/O operation has already been executed on

 SorA

                                             permission error(modify, stream, SorA)

Portability GNU Prolog predicate.

7.10.26 set stream eof action/2

Templates

     set stream eof action(+stream or alias, +atom)

Description

set stream eof action(SorA, Action) updates the eof action option associated with the stream-

term or alias SorA. The value of Action is one of the atoms error, eof code, reset as for open/4 (section 7.10.6, page 70). Errors

 SorA is a variable  instantiation error
 Action is a variable  instantiation error

Action is neither a variable nor a valid eof action domain error(eof action, Action) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(modify, stream, SorA)

Portability

GNU Prolog predicate.

7.10.27 set stream buffering/2

Templates

     set stream buffering(+stream or alias, +atom)

Description

set stream buffering(SorA, Buffering) updates the buffering mode associated with the stream-term or alias SorA. The value of Buffering is one of the atoms none, line or block as for open/4 (section 7.10.6, page 70). This predicate may only be used after opening a stream and before any other operations have been performed on it.

Errors

 SorA is a variable  instantiation error
 Buffering is a variable  instantiation error

Buffering is neither a variable nor a valid buffering mode

                                             domain error(buffering mode, Buffering)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

Portability

GNU Prolog predicate.

7.11 Constant term streams

7.11.1 Introduction

Constant term streams allow the user to consider a constant term (atom, character list or character code list) as a source/sink by associating to them a stream. Reading from a constant term stream will deliver the characters of the constant term as if they had been read from a standard file. Characters written on a constant term stream are stored to form the final constant term when the stream is closed. The built-in predicates described in this section allow the user to open and close a constant term stream for input or output. However, very often, a constant term stream is created to be only read or written once and then closed. To avoid the creation and the destruction of such a stream, GNU Prolog offers several built-in predicates to perform single input/output from/to constant terms (section 7.15, page 104).

7.11.2 open input atom stream/2, open input chars stream/2,
        open input codes stream/2

Templates

     open input atom stream(+atom, -stream)
     open input chars stream(+character list, -stream)
     open input codes stream(+character code list, -stream)

Description

open input atom stream(Atom, Stream) unifies Stream with the stream-term which is associated to a new input text-stream whose data are the characters of Atom.

open input chars stream(Chars, Stream) is similar to open input atom stream/2 except that data

are the content of the character list Chars.

open input codes stream(Codes, Stream) is similar to open input atom stream/2 except that data

are the content of the character code list Codes.

Errors

 Stream is not a variable  type error(variable, Stream)
 Atom is a variable  instantiation error

Chars is a partial list or a list with an element E which is a variable

                                             instantiation error

Codes is a partial list or a list with an element E which is a variable

                                             instantiation error

Atom is neither a variable nor a an atom type error(atom, Atom) Chars is neither a partial list nor a list type error(list, Chars) Codes is neither a partial list nor a list type error(list, Codes) an element E of the Chars list is neither a variable nor a character

                                             type error(character, E)

an element E of the Codes list is neither a variable nor an integer

                                             type error(integer, E)

an element E of the Codes list is an integer but not a character code

                                             representation error(character code)

Portability

GNU Prolog predicates.

7.11.3 close input atom stream/1, close input chars stream/1,
        close input codes stream/1

Templates

     close input atom stream(+stream or alias)
     close input chars stream(+stream or alias)
     close input codes stream(+stream or alias)

Description

close input atom stream(SorA) closes the constant term stream associated with the stream-term or alias SorA. SorA must a stream open with open input atom stream/2 (section 7.11.1, page 82).

close input chars stream(SorA) acts similarly for a character list stream. close input codes stream(SorA) acts similarly for a character code list stream.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(close, stream, SorA)

SorA is a stream-term or alias but does not refer to a constant term stream.

                                             domain error(term stream or alias, SorA)

Portability

GNU Prolog predicates.

7.11.4 open output atom stream/1, open output chars stream/1,
        open output codes stream/1

Templates

     open output atom stream(-stream)
     open output chars stream(-stream)
     open output codes stream(-stream)

Description

open output atom stream(Stream) unifies Stream with the stream-term which is associated to a new output text-stream. All characters written to this stream are collected and will be returned as an atom when the stream is closed by close ouput atom stream/2 (section 7.11.5, page 84).

open output chars stream(Stream) is similar to open output atom stream/1 except that the result

will be a character list.

open output codes stream(Stream) is similar to open output atom stream/1 except that the result

will be a character code list.

Errors

 Stream is not a variable  type error(variable, Stream)

Portability

GNU Prolog predicates.

7.11.5 close output atom stream/2, close output chars stream/2,
        close output codes stream/2

Templates

     close output atom stream(+stream or alias, ?atom)
     close output chars stream(+stream or alias, ?character list)
     close output codes stream(+stream or alias, ?character code list)

close output atom stream(SorA, Atom) closes the constant term stream associated with the streamterm or alias SorA. SorA must be associated to a stream open with open output atom stream/1 (section 7.11.4, page 84). Atom is unified with an atom formed with all characters written on the stream.

close output chars stream(SorA, Chars) acts similarly for a character list stream.

close output codes stream(SorA, Codes) acts similarly for a character code list stream.

Errors

 SorA is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) Chars is neither a partial list nor a list type error(list, Chars) Codes is neither a partial list nor a list type error(list, Codes) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(close, stream, SorA)

SorA is a stream-term or alias but does not refer to a constant term stream

                                             domain error(term stream or alias, SorA)

Portability

GNU Prolog predicates.

7.12 Character input/output

These built-in predicates enable a single character or character code to be input from and output to a text stream. The atom end of file is returned as character to indicate the end-of-file. -1 is returned as character code to indicate the end-of-file.

7.12.1 get char/2, get char/1, get code/1, get code/2

Templates

     get char(+stream or alias, ?in character)
     get char(?in character)
     get code(+stream or alias, ?in character code)
     get code(?in character code)

Description

get char(SorA, Char) succeeds if Char unifies with the next character read from the stream associated with the stream-term or alias SorA.

get code/2 is similar to get char/2 but deals with character codes.

get char/1 and get code/1 apply to the current input stream.

 SorA is a variable  instantiation error

Char is neither a variable nor an in-character type error(in character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a binary stream  permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

The entity input from the stream is not a character

                                             representation error(character)
 Code is an integer but not an in-character code  representation error(in character code)

Portability

ISO predicates.

7.12.2 get key/2, get key/1 get key no echo/2, get key no echo/1

Templates

     get key(+stream or alias, ?integer)
     get key(?integer)
     get key no echo(+stream or alias, ?integer)
     get key no echo(?integer)

Description

get key(Code, SorA) succeeds if Code unifies with the character code of the next key read from the stream associated with the stream-term or alias SorA. It is intended to read a single key from the keyboard (thus SorA should refer to current input stream). No buffering is performed (a character is read as soon as available) and function keys can also be read (in that case, Code is an integer > 255). The read character is echoed if it is printable.

This facility is only possible if the linedit facility has been installed (section 3.2.5, page 18) otherwise get key/2 behaves similarly to get code/2 (section 7.12.1, page 85) (the code of the first character is returned) but also pumps remaining characters until a character < space (0x20) is read (in particular RETURN). The same behavior occurs if SorA does not refer to the current input stream or if this stream is not attached to a terminal.

get key no echo/2 behaves similarly to get key/2 except that the read character is not echoed.

get key/1 and get key no echo/1 apply to the current input stream.

SorA is a variable  instantiation error

Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)
SorA is associated with a binary stream  permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

Portability

GNU Prolog predicates.

7.12.3 peek char/2, peek char/1, peek code/1, peek code/2

Templates

    peek char(+stream or alias, ?in character)
    peek char(?in character)
    peek code(+stream or alias, ?in character code)
    peek code(?in character code)

Description

peek char(SorA, Char) succeeds if Char unifies with the next character that will be read from the stream associated with the stream-term or alias SorA. The character is not read.

peek code/2 is similar to peek char/2 but deals with character codes.

peek char/1 and peek code/1 apply to the current input stream.

Errors

SorA is a variable  instantiation error

Char is neither a variable nor an in-character type error(in character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)
SorA is associated with a binary stream  permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

The entity input from the stream is not a character

                                           representation error(character)
Code is an integer but not an in-character code  representation error(in character code)

Portability

ISO predicates.

7.12.4 unget char/2, unget char/1, unget code/2, unget code/1

Templates

     unget char(+stream or alias, +character)
     unget char(+character)
     unget code(+stream or alias, +character code)
     unget code(+character code)

Description

unget char(SorA, Char) pushes back Char onto the stream associated with the stream-term or alias SorA. Char will be the next character read by get char/2. The maximum number of characters that can be cumulatively pushed back is given by the max unget Prolog flag (section 7.22.1, page 133).

unget code/2 is similar to unget char/2 but deals with character codes.

unget char/1 and unget code/1 apply to the current input stream.

Errors

 SorA is a variable  instantiation error
 Char is a variable  instantiation error
 Code is a variable  instantiation error

Char is neither a variable nor a character type error(character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a binary stream  permission error(input, binary stream,
                                             SorA)
 Code is an integer but not a character code  representation error(character code)

Portability

GNU Prolog predicates.

7.12.5 put char/2, put char/1, put code/1, put code/2, nl/1, nl/0

Templates

     put char(+stream or alias, +character)
     put char(+character)
     put code(+stream or alias, +character code)
     put code(+character code)
     nl(+stream or alias)
     nl

Description

put char(SorA, Char) writes Char onto the stream associated with the stream-term or alias SorA.

put code/2 is similar to put char/2 but deals with character codes.

nl(SorA) writes a new-line character onto the stream associated with the stream-term or alias SorA. This is equivalent to put char(SorA, '\n'). put char/1, put code/1 and nl/0 apply to the current output stream.

Errors

SorA is a variable  instantiation error
Char is a variable  instantiation error
Code is a variable  instantiation error

Char is neither a variable nor a character type error(character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an input stream  permission error(output, stream, SorA)
SorA is associated with a binary stream  permission error(output, binary stream,
                                           SorA)
Code is an integer but not a character code  representation error(character code)

Portability

ISO predicates.

7.13 Byte input/output

These built-in predicates enable a single byte to be input from and output to a binary stream. -1 is returned to indicate the end-of-file.

7.13.1 get byte/2, get byte/1

Templates

    get byte(+stream or alias, ?in byte)
    get byte(?in byte)

Description

get byte(SorA, Byte) succeeds if Byte unifies with the next byte read from the stream associated with the stream-term or alias SorA.

get byte/1 applies to the current input stream.

Errors

SorA is a variable  instantiation error

Byte is neither a variable nor an in-byte type error(in byte, Byte) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)
SorA is associated with a text stream  permission error(input, text stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

ISO predicates.

7.13.2 peek byte/2, peek byte/1

Templates

     peek byte(+stream or alias, ?in byte)
     peek byte(?in byte)

Description

peek byte(SorA, Byte) succeeds if Byte unifies with the next byte that will be read from the stream associated with the stream-term or alias SorA. The byte is not read.

peek byte/1 applies to the current input stream.

Errors

 SorA is a variable  instantiation error

Byte is neither a variable nor an in-byte type error(in byte, Byte) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a text stream  permission error(input, text stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

Portability

ISO predicates.

7.13.3 unget byte/2, unget byte/1

Templates

     unget byte(+stream or alias, +byte)
     unget byte(+byte)

Description

unget byte(SorA, Byte) pushes back Byte onto the stream associated with the stream-term or alias SorA. Byte will be the next byte read by get byte/2. The maximum number of bytes that can be successively pushed back is given by the max unget Prolog flag (section 7.22.1, page 133).

unget byte/1 applies to the current input stream.

 SorA is a variable  instantiation error
 Byte is a variable  instantiation error

Byte is neither a variable nor a byte type error(byte, Byte) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a text stream  permission error(input, text stream,
                                             SorA)

Portability

GNU Prolog predicates.

7.13.4 put byte/2, put byte/1

Templates

     put byte(+stream or alias, +byte)
     put byte(+byte)

Description

put byte(SorA, Byte) writes Byte onto the stream associated with the stream-term or alias SorA.

put byte/1 applies to the current output stream.

Errors

 SorA is a variable  instantiation error
 Byte is a variable  instantiation error

Byte is neither a variable nor a byte type error(byte, Byte) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(output, stream, SorA)
 SorA is associated with a text stream  permission error(output, text stream,
                                             SorA)

Portability

GNU Prolog predicates.

7.14 Term input/output

These built-in predicates enable a Prolog term to be input from or output to a text stream. The atom end of file is returned as term to indicate the end-of-file. The syntax of such terms can also be altered by changing the operators (section 7.14.10, page 100), and making some characters equivalent to others (section 7.14.12, page 102) if the char conversion Prolog flag is on (section 7.22.1, page 133). Double quoted tokens will be returned as an atom or a character list or a character code list depending on the value of the double quotes Prolog flag (section 7.22.1, page 133). Similarly, back quoted tokens are returned depending on the value of the back quotes Prolog flag.

7.14.1 read term/3, read term/2, read/2, read/1

Templates

read term(+stream or alias, ?term, +read option list)
read term(?term, +read option list)
read(+stream or alias, ?term)
read(?term)

Description

read term(SorA, Term, Options) is true if Term unifies with the next term read from the stream associated with the stream-term or alias SorA according to the options given by Options.

Read options: Options is a list of read options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• variables(VL): VL is unified with the list of all variables of the input term, in left-to-right traversal order. Anonymous variables are included in the list VL.

• variable names(VNL): VNL is unified with the list of pairs Name = Var where Var is a named variable of the term and Name is the atom associated to the name of Var. Anonymous variables are not included in the list VNL.

• singletons(SL): SL is unified with the list of pairs Name = Var where Var is a named variable which occurs only once in the term and Name is the atom associated to the name of Var. Anonymous variables are not included in the list SL.

• syntax error(error/warning/fail): specifies the effect of a syntax error:

  – error: a syntax error is raised.

– warning: a warning message is displayed and the predicate fails.

– fail: the predicate quietly fails. The default value is the value of the syntax error Prolog flag (section 7.22.1, page 133).

• end of term(dot/eof): specifies the end-of-term delimiter: dot is the classical full-stop delimiter (a dot followed with a layout character), eof is the end-of-file delimiter. This option is useful for predicates like read term from atom/3 (section 7.15.1, page 104) to avoid to add a terminal dot at the end of the atom. The default value is dot.

read(SorA, Term) is equivalent to read term(SorA, Term, []).

read term/2 and read/1 apply to the current input stream.

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a valid read option

                                             domain error(read option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a binary stream  permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

a syntax error occurs and the value of the

 syntax error Prolog flag is error

(section 7.22.1, page 133)

                                             syntax error(atom explaining the error)

Portability

ISO predicates. The ISO reference raises a representation error(Flag) where Flag is max arity, max integer, or min integer when the read term breaches an implementation defined limit specified by Flag. GNU Prolog detects neither min integer nor max integer violation and treats a max arity violation as a syntax error. The read options syntax error/1 and end of term/1 are GNU Prolog extensions.

7.14.2 read atom/2, read atom/1, read integer/2, read integer/1,
        read number/2, read number/1

Templates

     read atom(+stream or alias, ?atom)
     read atom(?atom)
     read integer(+stream or alias, ?integer)
     read integer(?integer)
     read number(+stream or alias, ?number)
     read number(?number)

Description

read atom(SorA, Atom) succeeds if Atom unifies with the next atom read from the stream associated with the stream-term or alias SorA.

read integer(SorA, Integer) succeeds if Integer unifies with the next integer read from the stream associated with the stream-term or alias SorA.

read number(SorA, Number) succeeds if Number unifies with the next number (integer or floating point number) read from the stream associated with the stream-term or alias SorA.

read atom/1, read integer/1 and read number/1 apply to the current input stream.

SorA is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

Integer is neither a variable nor an integer  type error(integer, Integer)
Number is neither a variable nor a number  type error(number, Number)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)
SorA is associated with a binary stream  permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

a syntax error occurs and the value of the

syntax error Prolog flag is error

(section 7.22.1, page 133)

                                           syntax error(atom explaining the error)

Portability GNU Prolog predicates.

7.14.3 read token/2, read token/1

Templates

    read token(+stream or alias, ?nonvar)
    read token(?nonvar)

Description read token(SorA, Token) succeeds if Token unifies with the encoding of the next Prolog token read from the stream associated with stream-term or alias SorA. Token encoding:

• var(A): a variable is read whose name is the atom A.

• an atom A: an atom A is read.

• integer N: an integer N is read.

• floating point number N: a floating point number N is read.

• string(A): a string (double quoted item) is read whose characters forms the atom A.

• punct(P): a punctuation character P is read (P is a one-character atom in ()[]{|}, the atom

    full stop or the atom end of file).

• back quotes(A): a back quoted item is read whose characters forms the atom A.

• extended(A): an extended character A (an atom) is read. As for read term/3, the behavior of read token/2 can be affected by some Prolog flags (section 7.14, page 91). read token/1 applies to the current input stream. Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)
 SorA is associated with a binary stream  permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

a syntax error occurs and the value of the

 syntax error Prolog flag is error

(section 7.22.1, page 133)

                                             syntax error(atom explaining the error)

Portability

GNU Prolog predicates.

7.14.4 syntax error info/4

Templates

     syntax error info(?atom, ?integer, ?integer, ?atom)

Description

syntax error info(FileName, Line, Column, Error) returns the information associated to the last

syntax error. Line is the line number of the error, Column is the column number of the error and Error is an atom explaining the error.

Errors

 FileName is neither a variable nor an atom  type error(atom, FileName)

Line is neither a variable nor an integer type error(integer, Line)

 Column is neither a variable nor an integer  type error(integer, Column)

Error is neither a variable nor an atom type error(atom, Error)

Portability

GNU Prolog predicate.

7.14.5 last read start line column/2

Templates

     last read start line column(?integer, ?integer)

Description

last read start line column(Line, Column) unifies Line and Column with the line number and the

column number associated to the start of the last read predicate. This predicate can be used after calling one of the following predicates: read term/3, read term/2, read/2, read/1 (section 7.14.1, page 92),

read atom/2, read atom/1, read integer/2, read integer/1, read number/2, read number/1 (sec-

tion 7.14.2, page 93) or read token/2, read token/1 (section 7.14.3, page 94).

Line is neither a variable nor an integer type error(integer, Line)

Column is neither a variable nor an integer  type error(integer, Column)

Portability

GNU Prolog predicate.

7.14.6 write term/3, write term/2, write/2, write/1, writeq/2, writeq/1,
   write canonical/2, write canonical/1, display/2, display/1, print/2,
   print/1

Templates

write term(+stream or alias, ?term, +write option list)
write term(?term, +write option list)
write(+stream or alias, ?term)
write(?term)
writeq(+stream or alias, ?term)
writeq(?term)
write canonical(+stream or alias, ?term)
write canonical(?term)
display(+stream or alias, ?term)
display(?term)
print(+stream or alias, ?term)
print(?term)

Description

write term(SorA, Term, Options) writes Term to the stream associated with the stream-term or alias SorA according to the options given by Options.

Write options: Options is a list of write options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• quoted(true/false): if true each atom and functor is quoted if this would be necessary for the term to be input by read term/3. If false no extra quotes are written. The default value is false.

• ignore ops(true/false): if true each compound term is output in functional notation (neither operator notation nor list notation is used). If false operator and list notations are used. The default value is false.

• numbervars(true/false): if true a term of the form '$VAR'(N), where N is an integer, is output as a variable name (see below). If false such a term is output normally (according to the other options). The default value is true.

• namevars(true/false): if true a term of the form '$VARNAME'(Name), where Name is an atom, is

output as a variable name (see below). If false such a term is output normally (according to the other options). The default value is true.

• space args(true/false): if true an extra space character is emitted after each comma separating the arguments of a compound term in functional notation or of a list. If false no extra space is emitted. The default value is false.

• portrayed(true/false): if true and if there exists a predicate portray/1, write term/3 acts

as follows: if Term is a variable it is simply written. If Term is non-variable then it is passed to portray/1. If this succeeds then it is assumed that Term has been output. Otherwise write term/3 outputs the principal functor of Term (Term itself if it is atomic) according to other options and recursively calls portray/1 on the components of Term (if it is a compound term). With ignore ops(false) a list is first passed to portray/1 and only if this call fails each element of the list is passed to portray/1 (thus every sub-list is not passed). The default value is false.

• max depth(N): controls the depth of output for compound terms. N is an integer specifying the depth. The output of a term whose depth is greater than N gives rise to the output of ... (3 dots). By default there is no depth limit.

• priority(N): specifies the starting priority to output the term. This option controls if Term should be enclosed in brackets. N is a positive integer ≤ 1200. By default N = 1200.

Variable numbering: when the numbervars(true) option is passed to write term/3 any term of the form '$VAR'(N) where N is an integer is output as a variable name consisting of a capital letter possibly followed by an integer. The capital letter is the (I+1)th letter of the alphabet and the integer is J, where I = N mod 26 and J = N // 26. The integer J is omitted if it is zero. For example:

'$VAR'(0) is written as A
'$VAR'(1) is written as B

     ...

     '$VAR'(25) is written as Z
     '$VAR'(26) is written as A1
     '$VAR'(27) is written as B1

Variable naming: when the namevars(true) option is passed to write term/3 any term of the form '$VARNAME'(Name) where Name is an atom is output as a variable name consisting of the characters Name. For example: '$VARNAME'('A') is written as A (even in the presence of the quoted(true) option).

write(SorA, Term) is equivalent to write term(SorA, Term, []).
writeq(SorA, Term) is equivalent to write term(SorA, Term, [quoted(true)]).
write canonical(SorA, Term) is equivalent to write term(SorA, Term, [quoted(true),
ignore ops(true), numbervars(false)]).
display(SorA, Term) is equivalent to write term(SorA, Term, [ignore ops(true),
numbervars(false)]).
print(SorA, Term) is equivalent to write term(SorA, Term, [numbervars(false),
portrayed(true)]).
write term/2, write/1, writeq/1, write canonical/1, display/1 and print/1 apply to the current

output stream.

Errors

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

an element E of the Options list is neither a variable nor a valid write-option

                                             domain error(write option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)
 SorA is associated with a binary stream  permission error(output, binary stream,
                                             SorA)

ISO predicates except display/1-2 and print/1-2 that are GNU Prolog predicates. namevars/1,

space args/1, portrayed/1, max depth/1 and priority/1 options are GNU Prolog extensions.

7.14.7 format/3, format/2

Templates

format(+stream or alias, +character code list or atom, +list)
format(+character code list or atom, +list)

Description

format(SorA, Format, Arguments) writes the Format string replacing each format control sequence F by the corresponding element of Arguments (formatted according to F) to the stream associated with the stream-term or alias SorA.

Format control sequences: the general format of a control sequence is '~NC'. The character C determines the type of the control sequence. N is an optional numeric argument. An alternative form of N is '*'. '*' implies that the next argument Arg in Arguments should be used as a numeric argument in the control sequence. The use of C printf() formatting sequence (beginning by the character %) is also allowed. The following control sequences are available:

Format sequence type of the argument

Description

~Na

atom print the atom without quoting. N is minimal number of characters to print using spaces on the rigth if needed (default: the length of the atom)

~Nc

character code print the character associated to the code. N is the number of times to print the character (default: 1)

~Nf ~Ne ~NE ~Ng ~NG

float expression pass the argument Arg and N to the C printf() function as:

if N is not specified printf("%f",Arg) else printf("%.Nf",Arg).

Similarly for ~Ne, ~NE, ~Ng and ~NG

~Nd

integer expression print the argument. N is the number of digits after the decimal point.

If N is 0 no decimal point is printed (default: 0)

~ND

integer expression identical to ~Nd except that ',' separates groups of three digits to

the left of the decimal point

~Nr

integer expression print the argument according to the radix N. 2 ≤ N ≤ 36 (default: 8).

The letters a-z denote digits > 9

~NR

integer expression identical to ~Nr except that the letters A-Z denote digits > 9

~Ns

character code list print exactly N characters (default: the length of the list)

~NS

character list print exactly N characters (default: the length of the list)

~i

term

ignore the current argument

~k

term

pass the argument to write canonical/1 (section 7.14.6, page 96)

~p

term

pass the argument to print/1 (section 7.14.6, page 96)

~q

term

pass the argument to writeq/1 (section 7.14.6, page 96)

~w

term

pass the argument to write/1 (section 7.14.6, page 96)

~~

none

print the character '~'

~Nn

none

print N new-line characters (default: 1)

~N

none

print a new-line character if not at the beginning of a line

~?

atom

use the argument as a nested format string

%F

atom, integer or float expression interface to the C function printf(3) for outputting atoms (C string), integers and floating point numbers. * are also allowed.

format/2 applies to the current output stream.

Errors

SorA is a variable  instantiation error

Format is a partial list or a list with an element E which is a variable

                                           instantiation error
Arguments is a partial list  instantiation error

Format is neither a partial list nor a list or an atom

                                           type error(list, Format)
Arguments is neither a partial list nor a list  type error(list, Arguments)

an element E of the Format list is neither a variable nor a character code

                                           representation error(character code, E)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)

an element E of Format is not a valid format control sequence

                                           domain error(format control sequence, E)

the Arguments list does not contain sufficient elements

                                           domain error(non empty list, [])

an element E of the Arguments list is a variable while a non-variable term was expected

                                           instantiation error

an element E of the Arguments list is neither variable nor an atom while an atom was expected

                                           type error(atom, E)

an element E of the Arguments cannot be evaluated as an arithmetic expression while an integer or a floating point number was expected

an arithmetic error (section 7.6.1, page 57)

an element E of the Arguments list is neither variable nor character code while a character code was expected

                                           representation error(character code, E)

SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an input stream  permission error(output, stream, SorA)
SorA is associated with a binary stream  permission error(output, binary stream,
                                           SorA)

Portability

GNU Prolog predicates.

7.14.8 portray clause/2, portray clause/1

Templates

    portray clause(+stream or alias, +clause)
    portray clause(+clause)

Description

portray clause(SorA, Clause) pretty prints Clause to the stream associated with the stream-term or alias SorA. portray clause/2 uses the variable binding predicates name singleton vars/1 (section 7.5.1, page 55) and numbervars/1 (section 7.5.3, page 56). This predicate is used by listing/1 (section 7.23.3, page 138).

portray clause/1 applies to the current output stream.

 Clause is a variable  instantiation error

Clause is neither a variable nor a callable term type error(callable, Clause)

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)
 SorA is associated with a binary stream  permission error(output, binary stream,
                                             SorA)

Portability

GNU Prolog predicates.

7.14.9 get print stream/1

Templates

     get print stream(?stream)

Description

get print stream(Stream) unifies Stream with the stream-term associated to the output stream used by print/2 (section 7.14.6, page 96). The purpose of this predicate is to allow a user-defined portray/1 predicate to identify the output stream in use.

Errors

Stream is neither a variable nor a stream-term domain error(stream, Stream)

Portability

GNU Prolog predicate.

7.14.10 op/3

Templates

     op(+integer, +operator specifier, +atom or atom list)

Description

op(Priority, OpSpecifier, Operator) alters the operator table. Operator is declared as an operator with properties defined by specifier OpSpecifier and Priority. Priority must be an integer ≥ 0 and ≤ 1200. If Priority is 0 then the operator properties of Operator (if any) are canceled. Operator may also be a list of atoms in which case all of them are declared to be operators. In general, operators can be removed from the operator table and their priority or specifier can be changed. However, it is an error to attempt to change the ',' operator from its initial status. An atom can have multiple operator definitions (e.g. prefix and infix like +) however an atom cannot have both an infix and a postfix operator definitions.

Operator specifiers: the following specifiers are available:

Specifier Type Associativity
fx  prefix  no
fy  prefix  yes
xf  postfix  no
yf  postfix  yes
xfx  infix  no
yfx  infix  left
xfy  infix  right

Prolog predefined operators:

Priority Specifier Operators

1200    xfx    :-  -->
1200     fx     :-
1100    xfy    ;
1050    xfy    ->
1000    xfy    ,
900     fy     \+
700    xfx    {=  \=  =..   ==  \==  @<  @=<  @>  @>=  is  =:=  =\=  <  =< \atop >  >=}
600    xfy    :
500    yfx    +  -  /\  \/
400    yfx    *  /  //  rem  mod  <<  >>
200    xfy    **  ^
200     fy     +  -  \

FD predefined operators:

Priority Specifier Operators

750    xfy    #<=>  #\<=>
740    xfy    #==>  #\==>
730    xfy    ##  #\/  #\\/
720    yfx    #/\  #\/\
710     fy     #\
700    xfx    #=  #\=  #<  #=<  #>  #>=  #=#  #\=#  #<#  #=<#  #>#  #>=#
500    yfx    +  -
400    yfx    *  /  //  rem
200    xfy    **
200     fy     +  -

Priority is a variable  instantiation error
OpSpecifier is a variable  instantiation error

Operator is a partial list or a list with an element E which is a variable

                                           instantiation error
Priority is neither a variable nor an integer  type error(integer, Priority)
OpSpecifier is neither a variable nor an atom  type error(atom, OpSpecifier)

Operator is neither a partial list nor a list nor an atom

                                           type error(list, Operator)

an element E of the Operator list is neither a variable nor an atom

                                           type error(atom, E)
Priority is an integer not ≥ 0 and ≤ 1200  domain error(operator priority, Priority)
OpSpecifier is not a valid operator specifier  domain error(operator specifier,
                                           OpSpecifier)
Operator is ',' or an element of the Operator

list is ','

                                           permission error(modify, operator, ',')
OpSpecifier is a specifier such that Operator

would have a postfix and an infix definition.

                                           permission error(create, operator,
                                           Operator)

Portability

ISO predicate.

The ISO reference implies that if a program calls current op/3, then modifies an operator definition by calling op/3 and backtracks into the call to current op/3, then the changes are guaranteed not to affect that current op/3 goal. This is not guaranteed by GNU Prolog.

7.14.11 current op/3

Templates

    current op(?integer, ?operator specifier, ?atom)

Description

current op(Priority, OpSpecifier, Operator) succeeds if Operator is an operator with properties

defined by specifier OpSpecifier and Priority. This predicate is re-executable on backtracking.

Errors

Priority is neither a variable nor an operator priority

                                           domain error(operator priority, Priority)

OpSpecifier is neither a variable nor an operator specifier

                                           domain error(operator specifier,
                                           OpSpecifier)
Operator is neither a variable nor an atom  type error(atom, Operator)

Portability

ISO predicate.

7.14.12 char conversion/2

Templates

    char conversion(+character, +character)

Description

char conversion(InChar, OutChar) alters the character-conversion mapping. This mapping is used by the following read predicates: read term/3 (section 7.14.1, page 92), read atom/2, read integer/2, read number/2 (section 7.14.2, page 93) and read token/2 (section 7.14.3, page 94) to replace any occurrence of a character InChar by OutChar. However the conversion mechanism should have been previously activated by switching on the char conversion Prolog flag (section 7.22.1, page 133). When InChar and OutChar are the same, the effect is to remove any conversion of a character InChar.

Note that the single character read predicates (e.g. get char/2) never do character conversion. If such behavior is required, it must be explicitly done using current char conversion/2 (section 7.14.13, page 103).

Errors

 InChar is a variable  instantiation error
 OutChar is a variable  instantiation error
 InChar is neither a variable nor a character  type error(character, InChar)
 OutChar is neither a variable nor a character  type error(character, OutChar)

Portability

ISO predicate. The type error(character,...) is a GNU Prolog behavior, the ISO reference instead defines a representation error(character) in this case. This seems to be an error of the ISO reference since, for many other built-in predicates accepting a character (e.g. char code/2, put char/2), a

type error is raised.

The ISO reference implies that if a program calls current char conversion/2, then modifies the character mapping by calling char conversion/2, and backtracks into the call to current char conversion/2 then the changes are guaranteed not to affect that current char conversion/2 goal. This is not guaranteed by GNU Prolog.

7.14.13 current char conversion/2

Templates

     current char conversion(?character, ?character)

Description

current char conversion(InChar, OutChar) succeeds if the conversion of InChar is OutChar according

to the character-conversion mapping. In that case, InChar and OutChar are different. This predicate is re-executable on backtracking.

Errors

 InChar is neither a variable nor a character  type error(character, InChar)
 OutChar is neither a variable nor a character  type error(character, OutChar)

Portability

ISO predicate. Same remark as for char conversion/2 (section 7.14.12, page 102).

7.15 Input/output from/to constant terms

These built-in predicates enable a Prolog term to be input from or output to a Prolog constant term (atom, character list or character code list). All these predicates can be defined using constant term streams (section 7.11, page 82). They are however simpler to use.

7.15.1 read term from atom/3, read from atom/2, read token from atom/2

Templates

     read term from atom(+atom ?term, +read option list)
     read from atom(+atom, ?term)
     read token from atom(+atom, ?nonvar)

Description

Like read term/3, read/2 (section 7.14.1, page 92) and read token/2 (section 7.14.3, page 94) except that characters are not read from a text-stream but from Atom; the atom given as first argument.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) see associated predicate errors (section 7.14.1, page 92) and (section 7.14.3,

page 94)

Portability

GNU Prolog predicates.

7.15.2 read term from chars/3, read from chars/2, read token from chars/2

Templates

     read term from chars(+character list ?term, +read option list)
     read from chars(+character list, ?term)
     read token from chars(+character list, ?nonvar)

Description

Like read term/3, read/2 (section 7.14.1, page 92) and read token/2 (section 7.14.3, page 94) except that characters are not read from a text-stream but from Chars; the character list given as first argument.

Errors

Chars is a partial list or a list with an element E which is a variable

                                             instantiation error

Chars is neither a partial list nor a list type error(list, Chars) an element E of the Chars list is neither a variable nor a character

                                             type error(character, E)

see associated predicate errors (section 7.14.1, page 92) and (section 7.14.3, page 94)

Portability

7.15.3 read term from codes/3, read from codes/2, read token from codes/2

Templates

     read term from codes(+character code list ?term, +read option list)
     read from codes(+character code list, ?term)
     read token from codes(+character code list, ?nonvar)

Description

Like read term/3, read/2 (section 7.14.1, page 92) and read token/2 (section 7.14.3, page 94) except that characters are not read from a text-stream but from Codes; the character code list given as first argument.

Errors

Codes is a partial list or a list with an element E which is a variable

                                             instantiation error

Codes is neither a partial list nor a list type error(list, Codes) an element E of the Codes list is neither a variable nor an integer

                                             type error(integer, E)

an element E of the Codes list is an integer but not a character code

                                             representation error(character code, E)

see associated predicate errors (section 7.14.1, page 92) and (section 7.14.3, page 94)

Portability

GNU Prolog predicates.

7.15.4 write term to atom/3, write to atom/2, writeq to atom/2,
        write canonical to atom/2, display to atom/2, print to atom/2,
        format to atom/3

Templates

     write term to atom(?atom, ?term, +write option list)
     write to atom(?atom, ?term)
     writeq to atom(?atom, ?term)
     write canonical to atom(?atom, ?term)
     display to atom(?atom, ?term)
     print to atom(?atom, ?term)
     format to atom(?atom, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 7.14.6,

page 96) and format/3 (section 7.14.7, page 98) except that characters are not written onto a text-stream but are collected as an atom which is then unified with the first argument Atom.

Errors

Atom is neither a variable nor an atom type error(atom, Atom) see associated predicate errors (section 7.14.6, page 96) and (section 7.14.7,

page 98)

GNU Prolog predicates.

7.15.5 write term to chars/3, write to chars/2, writeq to chars/2,
        write canonical to chars/2, display to chars/2, print to chars/2,
        format to chars/3

Templates

     write term to chars(?character list, ?term, +write option list)
     write to chars(?character list, ?term)
     writeq to chars(?character list, ?term)
     write canonical to chars(?character list, ?term)
     display to chars(?character list, ?term)
     print to chars(?character list, ?term)
     format to chars(?character list, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 7.14.6,

page 96) and format/3 (section 7.14.7, page 98) except that characters are not written onto a text-stream but are collected as a character list which is then unified with the first argument Chars.

Errors

Chars is neither a partial list nor a list type error(list, Chars) see associated predicate errors (section 7.14.6, page 96) and (section 7.14.7,

page 98)

Portability

GNU Prolog predicates.

7.15.6 write term to codes/3, write to codes/2, writeq to codes/2,
        write canonical to codes/2, display to codes/2, print to codes/2,
        format to codes/3

Templates

     write term to codes(?character code list, ?term, +write option list)
     write to codes(?character code list, ?term)
     writeq to codes(?character code list, ?term)
     write canonical to codes(?character code list, ?term)
     display to codes(?character code list, ?term)
     print to codes(?character code list, ?term)
     format to codes(?character code list, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 7.14.6,

page 96) and format/3 (section 7.14.7, page 98) except that characters are not written onto a text-stream but are collected as a character code list which is then unified with the first argument Codes.

Codes is neither a partial list nor a list type error(list, Codes) see associated predicate errors (section 7.14.6, page 96) and (section 7.14.7,

page 98)

Portability

GNU Prolog predicates.

7.16 DEC-10 compatibility input/output

7.16.1 Introduction

The DEC-10 Prolog I/O predicates manipulate streams implicitly since they only refer to current input/output streams (section 7.10.1, page 67). The current input and output streams are initially set to user input and user output respectively. The predicate see/1 (resp. tell/1, append/1) can be used for setting the current input (resp. output) stream to newly opened streams for particular files. The predicate seen/0 (resp. told/0) close the current input (resp. output) stream, and resets it to the standard input (resp. output). The predicate seeing/1 (resp. telling/1) is used for retrieving the file name associated with the current input (resp. output) stream. The file name user stands for the standard input or output, depending on context (user input and user output can also be used). The DEC-10 Prolog I/O predicates are only provided for compatibility, they are now obsolete and their use is discouraged. The predicates for explicit stream manipulation should be used instead (section 7.10, page 67).

7.16.2 see/1, tell/1, append/1

Templates

     see(+source sink)
     see(+stream)
     tell(+source sink)
     tell(+stream)
     append(+source sink)
     append(+stream)

Description

see(FileName) sets the current input stream to FileName. If there is a stream opened by see/1 associated with the same FileName already, then it becomes the current input stream. Otherwise, FileName is opened for reading and becomes the current input stream.

tell(FileName) sets the current output stream to FileName. If there is a stream opened by tell/1 associated with the same FileName already, then it becomes the current output stream. Otherwise, FileName is opened for writing and becomes the current output stream.

append(FileName) like tell/1 but FileName is opened for writing + append.

A stream-term (obtained with any other built-in predicate) can also be provided as FileName to these predicates.

Errors

See errors associated to open/4 (section 7.10.6, page 70).

Portability GNU Prolog predicates.

7.16.3 seeing/1, telling/1

Templates

     seeing(?source sink)
     telling(?source sink)

Description seeing(FileName) succeeds if FileName unifies with the name of the current input file, if it was opened by see/1; else with the current input stream-term, if this is not user input, otherwise with user. telling(FileName) succeeds if FileName unifies with the name of the current output file, if it was opened by tell/1 or append/1; else with the current output stream-term, if this is not user output, otherwise with user. Errors None. Portability GNU Prolog predicates.

7.16.4 seen/0, told/0

Templates

     seen
     told

Description seen closes the current input, and resets it to user input. told closes the current output, and resets it to user output. Errors None. Portability GNU Prolog predicates.

7.16.5 get0/1, get/1, skip/1

Templates

     get0(?in character code)
     get(?in character code)
     skip(+character code)

Description get0(Code) succeeds if Code unifies with the next character code read from the current input stream. Thus it is equivalent to get code(Code) (section 7.12.1, page 85). get(Code) succeeds if Code unifies with the next character code read from the current input stream that is not a layout character. skip(Code) skips just past the next character code Code from the current input stream. Errors See errors for get code/2 (section 7.12.1, page 85). Portability GNU Prolog predicates.

7.16.6 put/1, tab/1

Templates

     put(+character code)
     tab(+evaluable)

Description put(Code) writes the character whose code is Code onto the current output stream. It is equivalent to put code(Code) (section 7.12.5, page 88). tab(N) writes N spaces onto the current output stream. N may be an arithmetic expression. Errors See errors for put code/2 (section 7.12.5, page 88) and for arithmetic expressions (section 7.6.1, page 57). Portability GNU Prolog predicates.

7.17 Term expansion

7.17.1 Definite clause grammars Definite clause grammars are a useful notation to express grammar rules. However the ISO reference does not include them, so they should be considered as a system dependent feature. Definite clause grammars are an extension of context-free grammars. A grammar rule is of the form:

head --> body. --> is a predefined infix operator (section 7.14.10, page 100).

Here are some features of definite clause grammars:

• a non-terminal symbol may be any callable term.

• a terminal symbol may be any Prolog term and is written as a list. The empty list represents an empty sequence of terminals.

• a sequence is expressed using the Prolog conjunction operator ((’,’)/2).

• the head of a grammar rule consists of a non-terminal optionally followed by a sequence of terminals (i.e. a Prolog list).

• the body of a grammar rule consists of a sequence of non-terminals, terminals, predicate call, disjunction (using ;/2), if-then (using (->)/2) or cut (using !).

• a predicate call must be enclosed in curly brackets (using {}/1). This makes it possible to express

an extra condition. A grammar rule is nothing but a “syntactic sugar” for a Prolog clause. Each grammar rule accepts as input a list of terminals (tokens), parses a prefix of this list and gives as output the rest of this list (possibly enlarged). This rest is generally parsed later. So, each a grammar rule is translated into a Prolog clause that explicitly the manages the list. Two arguments are then added: the input list (Start) and the output list (Stop). For instance:

     p --> q.

is translated into:

     p(Start, End) :- q(Start, End).

Extra arguments can be provided and the body of the rule can contain several non-terminals. Example:

     p(X, Y) -->
             q(X),
             r(X, Y),
             s(Y).

is translated into:

     p(X, Y, Start, End) :-
             q(X, Start, A),
             r(X, Y, A, B),
             s(Y, B, End).

Terminals are translated using unification:

     assign(X,Y) --> left(X), [:=], right(Y), [;].

is translated into:

     assign(X,Y,Start,End) :-
             left(X, Start, A),
             A=[:=|B],
             right(Y, B, C),
             C=[;|End].

Terminals appearing on the left-hand side of a rule are connected to the output argument of the head. It is possible to include a call to a prolog predicate enclosing it in curly brackets (to distinguish them from non-terminals):

     assign(X,Y) --> left(X), [:=], right(Y0), {Y is Y0 }, [;].

is translated into:

     assign(X,Y,Start,End) :-
             left(X, Start, A),
             A=[:=|B],
             right(Y0, B, C),
             Y is Y0,
             C=[;|End].

Cut, disjunction and if-then(-else) are translated literally (and do not need to be enclosed in curly brackets).

7.17.2 expand term/2, term expansion/2

Templates

     expand term(?term, ?term)
     term expansion(?term, ?term)

Description

expand term(Term1, Term2) succeeds if Term2 is a transformation of Term1. The transformation steps are as follows:

• if Term1 is a variable, it is unified with Term2

• if term expansion(Term1, Term2) succeeds Term2 is assumed to be the transformation of Term1.

• if Term1 is a DCG then Term2 is its translation (section 7.17.1, page 109).

• otherwise Term2 is unified with Term1.

term expansion(Term1, Term2) is a hook predicate allowing the user to define a specific transformation.

The GNU Prolog compiler (section 3.4, page 20) automatically calls expand term/2 on each Term1 read in. However, in the current release, only DCG transformation are done by the compiler (i.e.

term expansion/2 cannot be used). To use term expansion/2, it is necessary to call expand term/2

explicitly.

Errors

None.

Portability

GNU Prolog predicate.

7.17.3 phrase/3, phrase/2

Templates

     phrase(?term, ?list, ?list)
     phrase(?term, ?list)

Description

phrase(Phrase, List, Remainder) succeeds if the list List is in the language defined by the grammar rule body Phrase. Remainder is what remains of the list after a phrase has been found.

phrase(Phrase, List) is equivalent to phrase(Phrase, List, []).

Errors

List is neither a list nor a partial list type error(list, List)

 Remainder is neither a list nor a partial list  type error(list, Remainder)

Portability

GNU Prolog predicates.

7.18 Logic, control and exceptions

7.18.1 abort/0, stop/0, top level/0, break/0, halt/1, halt/0

Templates

     abort
     stop
     top level
     break
     halt(+integer)
     halt

Description

abort aborts the current execution. If this execution was initiated under a top-level the control is given back to the top-level and the message {execution aborted} is displayed. Otherwise, e.g. execution started by a initialization/1 directive (section 6.1.13, page 45), abort/0 is equivalent to halt(1) (see below).

stop stops the current execution. If this execution was initiated under a top-level the control is given back to the top-level. Otherwise, stop/0 is equivalent to halt(0) (see below).

top level starts a new recursive top-level (including the banner display). To end this new top-level simply type the end-of-file key sequence (Ctl-D) or its term representation: end of file.

break invokes a recursive top-level (no banner is displayed). To end this new level simply type the end-of-file key sequence (Ctl-D) or its term representation: end of file.

halt(Status) causes the GNU Prolog process to immediately exit back to the shell with the return code

Status.

halt is equivalent to halt(0).

Errors

 Status is a variable  instantiation error
 Status is neither a variable nor an integer  type error(integer, Status)

Portability

halt/1 and halt/0 are ISO predicates. abort/0, stop/0, top level/0 and break/0 are GNU Prolog

predicates.

7.18.2 once/1, (\+)/1 - not provable, call with args/1-11, call/2

Templates

     once(+callable term)
     \+(+callable term)
     call with args(+atom, +term,..., +term)
     call(+callable term, ?boolean)

Description

once(Goal) succeeds if call(Goal) succeeds. However once/1 is not re-executable on backtracking since all alternatives of Goal are cut. once(Goal) is equivalent to call(Goal), !.

\+ Goal succeeds if call(Goal) fails and fails otherwise. This built-in predicate gives negation by failure.

call with args(Functor, Arg1,..., ArgN) calls the goal whose functor is Functor and whose argu-

ments are Arg1,…, ArgN (0 ≤ N ≤ 10).

call(Goal, Deterministic) succeeds if call(Goal) succeeds and unifies Deterministic with true if

Goal has not created any choice-points, with false otherwise.

\+ is a predefined prefix operator (section 7.14.10, page 100).

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

 Functor is a variable  instantiation error
 Functor is neither a variable nor an atom  type error(atom, Functor)

Deterministic is neither a variable nor a boolean

                                             type error(boolean, Deterministic)

Portability

once/1 and (\+)/1 are ISO predicates, call with args/1-11 and call/2 are GNU Prolog predicates.

7.18.3 repeat/0

Templates

     repeat

Description

repeat generates an infinite sequence of backtracking choices. The purpose is to repeatedly perform some action on elements which are somehow generated, e.g. by reading them from a stream, until some test becomes true. Repeat loops cannot contribute to the logic of the program. They are only meaningful if the action involves side-effects. The only reason for using repeat loops instead of a more natural tailrecursive formulation is efficiency: when the test fails back, the Prolog engine immediately reclaims any working storage consumed since the call to repeat/0.

None. Portability ISO predicate.

7.18.4 for/3 Templates

     for(?integer, +integer, +integer)

Description for(Counter, Lower, Upper) generates an sequence of backtracking choices instantiating Counter to

the values Lower, Lower+1,. . . , Upper.

Errors

 Counter is neither a variable nor an integer  type error(integer, Counter)
 Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

 Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper) Portability GNU Prolog predicate.

7.19 Atomic term processing

These built-in predicates enable atomic terms to be processed as a sequence of characters and character codes. Facilities exist to split and join atoms, to convert a single character to and from the corresponding character code, and to convert a number to and from a list of characters and character codes.

7.19.1 atom length/2

Templates

     atom length(+atom, ?integer)

Description atom length(Atom, Length) succeeds if Length unifies with the number of characters of the name of

Atom.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

 Length is neither a variable nor an integer  type error(integer, Length)
 Length is an integer < 0  domain error(not less than zero, Length)

Portability ISO predicate.

7.19.2 atom concat/3

Templates

     atom concat(+atom, +atom, ?atom)
     atom concat(?atom, ?atom, +atom)

Description

atom concat(Atom1, Atom2, Atom12) succeeds if the name of Atom12 is the concatenation of the name of Atom1 with the name of Atom1. This predicate is re-executable on backtracking (e.g. if Atom12 is instantiated and both Atom1 and Atom2 are variables).

Errors

 Atom1 and Atom12 are variables  instantiation error
 Atom2 and Atom12 are variables  instantiation error

Atom1 is neither a variable nor an atom type error(atom, Atom1) Atom2 is neither a variable nor an atom type error(atom, Atom2)

 Atom12 is neither a variable nor an atom  type error(atom, Atom12)

Portability

ISO predicate.

7.19.3 sub atom/5

Templates

     sub atom(+atom, ?integer, ?integer, ?integer, ?atom)

Description

sub atom(Atom, Before, Length, After, SubAtom) succeeds if atom Atom can be split into three

atoms, AtomL, SubAtom and AtomR such that Before is the number of characters of the name of AtomL, Length is the number of characters of the name of SubAtom and After is the number of characters of the name of AtomR. This predicate is re-executable on backtracking.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

 SubAtom is neither a variable nor an atom  type error(atom, SubAtom)
 Before is neither a variable nor an integer  type error(integer, Before)
 Length is neither a variable nor an integer  type error(integer, Length)

After is neither a variable nor an integer type error(integer, After)

 Before is an integer < 0  domain error(not less than zero, Before)
 Length is an integer < 0  domain error(not less than zero, Length)
 After is an integer < 0  domain error(not less than zero, After)

Portability

ISO predicate.

7.19.4 char code/2

Templates

     char code(+character, ?character code)
     char code(-character, +character code)

Description char code(Char, Code) succeeds if the character code for the one-char atom Char is Code. Errors

 Char and Code are variables  instantiation error

Char is neither a variable nor a one-char atom type error(character, Char) Code is neither a variable nor an integer type error(integer, Code)

 Code is an integer but not a character code  representation error(character code)

Portability ISO predicate.

7.19.5 lower upper/2

Templates

     lower upper(+character, ?character)
     lower upper(-character, +character)

Description lower upper(Char1, Char2) succeeds if Char1 and Char2 are one-char atoms and if Char2 is the upper conversion of Char1. If Char1 (resp. Char2) is a character that is not a lower (resp. upper) letter then

Char2 is equal to Char1.

Errors

 Char1 and Char2 are variables  instantiation error

Char1 is neither a variable nor a one-char atom type error(character, Char1) Char2 is neither a variable nor a one-char atom type error(character, Char2) Portability GNU Prolog predicate.

7.19.6 atom chars/2, atom codes/2

Templates

     atom chars(+atom, ?character list)
     atom chars(-atom, +character list)
     atom codes(+atom, ?character code list)
     atom codes(-atom, +character code list)

Description atom chars(Atom, Chars) succeeds if Chars is the list of one-char atoms whose names are the successive characters of the name of Atom.

atom codes(Atom, Codes) is similar to atom chars/2 but deals with a list of character codes.

Errors

Atom is a variable and Chars (or Codes) is a partial list or a list with an element which is a variable

                                             instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) Chars is neither a list nor a partial list type error(list, Chars) Codes is neither a list nor a partial list type error(list, Codes) Atom is a variable and an element E of the list Chars is neither a variable nor a one-char atom

                                             type error(character, E)

Atom is a variable and an element E of the list Codes is neither a variable nor an integer

                                             type error(integer, E)

Atom is a variable and an element E of the list Codes is an integer but not a character code

                                             representation error(character code)

Portability

ISO predicates. The ISO reference only causes a type error(list, Chars) if Atom is a variable and Chars is neither a list nor a partial list. GNU Prolog always checks if Chars is a list. Similarly for Codes. The type error(integer, E) when an element E of the Codes is not an integer is a GNU Prolog extension. This seems to be an omission in the ISO reference since this error is detected for many other built-in predicates accepting a character code (e.g. char code/2, put code/2).

7.19.7 number atom/2, number chars/2, number codes/2

Templates

     number atom(+number, ?atom)
     number atom(-number, +atom)
     number chars(+number, ?character list)
     number chars(-number, +character list)
     number codes(+number, ?character code list)
     number codes(-number, +character code list)

Description

number atom(Number, Atom) succeeds if Atom is an atom whose name corresponds to the characters of

Number.

number chars(Number, Chars) is similar to number atom/2 but deals with a list of character codes.

number codes(Number, Codes) is similar to number atom/2 but deals with a list of characters.

Number and Atom are variables  instantiation error

Number is a variable and Chars (or Codes) is a partial list or a list with an element which is a variable

                                           instantiation error

Number is neither a variable nor an number  type error(number, Number)

Atom is neither a variable nor an atom type error(atom, Atom) Number is a variable and Chars is neither a list nor a partial list

                                           type error(list, Chars)

Number is a variable and Codes is neither a list nor a partial list

                                           type error(list, Codes)

Number is a variable and an element E of the list Chars is neither a variable nor a one-char atom

                                           type error(character, E)

Number is a variable and an element E of the list Codes is neither a variable nor an integer

                                           type error(integer, E)

Number is a variable and an element E of the list Codes is an integer but not a character code

                                           representation error(character code)
Number is a variable, Atom (or Chars or Codes)

cannot be parsed as a number and the value of

the syntax error Prolog flag is error

(section 7.22.1, page 133)

                                           syntax error(atom explaining the error)

Portability

number atom/2 is a GNU Prolog predicate. number chars/2 and number codes/2 are ISO predicates.

GNU Prolog only raises an error about an element E of the Chars (or Codes) list when Number is a variable while the ISO reference always check this. This seems an error since the list itself is only checked if Number is a variable.

The type error(integer, E) when an element E of the Codes is not an integer is a GNU Prolog extension. This seems to be an omission in the ISO reference since this error is detected for many other built-in predicates accepting a character code (e.g. char code/2, put code/2).

7.19.8 name/2

Templates

    name(+atomic, ?character code list)
    name(-atomic, +character code list)

Description

name(Constant, Codes) succeeds if Codes is a list whose elements are the character codes corresponding to the successive characters of Constant (a number or an atom). However, there atoms are for which name(Constant, Codes) is true, but which will not be constructed if name/2 is called with Constant uninstantiated, e.g. the atom '1024'. For this reason the use of name/2 is discouraged and should be limited to compatibility purposes. It is preferable to use atom codes/2 (section 7.19.6, page 116) or number chars/2 (section 7.19.7, page 117).

Constant is a variable and Codes is a partial list or a list with an element which is a variable

                                             instantiation error

Constant is neither a variable nor an atomic term

                                             type error(atomic, Constant)

Constant is a variable and Codes is neither a list nor a partial list

                                             type error(list, Codes)

Constant is a variable and an element E of the list Codes is neither a variable nor an integer

                                             type error(integer, E)

Constant is a variable and an element E of the list Codes is an integer but not a character code

                                             representation error(character code)

Portability

GNU Prolog predicate.

7.19.9 atom hash/2

Templates

     atom hash(+atom, ?integer)
     atom hash(?atom, +integer)

Description

atom hash(Atom, Hash) succeeds if Hash is the internal key associated to Atom (an existing atom). The internal key of an atom is a unique integer ≥ 0 and < to the max atom Prolog flag (section 7.22.1, page 133).

Errors

 Atom and Hash are both variables  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) Hash is neither a variable nor an integer type error(integer, Hash)

 Hash is an integer < 0  domain error(not less than zero, Hash)

Portability

GNU Prolog predicate.

7.19.10 new atom/3, new atom/2, new atom/1

Templates

     new atom(+atom, +integer, -atom)
     new atom(+atom, -atom)
     new atom(-atom)

Description

new atom(Prefix, Hash, Atom) unifies Atom with a new atom whose name begins with the characters of the name of Prefix and whose internal key is Hash (section 7.19.9, page 119). This predicate is then a symbol generator. It is guaranteed that Atom does not exist before the invocation of new atom/3. The characters appended to Prefix to form Atom are in: A-Z (capital letter), a-z (small letter), 0-9 (digit), #, $, &, , @.

new atom/2 is similar to new atom/3, but the atom generated can have any (free) internal key. new atom/1 is similar to new atom(atom , Atom), i.e. the generated atom begins with atom . Errors

 Prefix is a variable  instantiation error
 Hash is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)

Hash is neither a variable nor an integer type error(integer, Hash)

 Hash is an integer < 0  domain error(not less than zero, Hash)
 Atom is not a variable  type error(variable, Atom)

Portability GNU Prolog predicate.

7.19.11 current atom/1

Templates

     current atom(?atom)

Description current atom(Atom) succeeds if there exists an atom that unifies with Atom. All atoms are found except those beginning with a '$' (system atoms). This predicate is re-executable on backtracking. Errors Atom is neither a variable nor an atom type error(atom, Atom) Portability GNU Prolog predicate.

7.19.12 atom property/2

Templates

     atom property(?atom, ?atom property)

Description

atom property(Atom, Property) succeeds if current atom(Atom) succeeds (section 7.19.11, page 120)

and if Property unifies with one of the properties of the atom. This predicate is re-executable on backtracking. Atom properties:

• length(Length): Length is the length of the name of the atom.

• hash(Hash): Hash is the internal key of the atom (section 7.19.9, page 119).

• prefix op: if there is a prefix operator currently defined with this name.

• infix op: if there is an infix operator currently defined with this name.

• postfix op: if there is a postfix operator currently defined with this name.

• needs quotes: if the atom must be quoted to be read later.

• needs scan: if the atom must be scanned when output to be read later (e.g. contains special characters that must be output with a \ escape sequence).

Errors

Atom is neither a variable nor an atom type error(atom, Atom) Property is neither a variable nor a n atom property term

                                             domain error(atom property, Property)
 Property = length(E) or hash(E) and E is

neither a variable nor an integer

                                             type error(integer, E)

Portability

GNU Prolog predicate.

7.20 List processing

These predicates manipulate lists. They are bootstrapped predicates (i.e. written in Prolog) and no error cases are tested (for the moment). However, since they are written in Prolog using other built-in predicates, some errors can occur due to those built-in predicates.

7.20.1 append/3

Templates

     append(?list, ?list, ?list)

Description

append(List1, List2, List12) succeeds if the concatenation of the list List1 and the list List2 is the list List12. This predicate is re-executable on backtracking (e.g. if List12 is instantiated and both List1 and List2 are variable).

Errors

None.

Portability

GNU Prolog predicate.

7.20.2 member/2, memberchk/2

Templates

     member(?term, ?list)
     memberchk(?term, ?list)

Description

member(Element, List) succeeds if Element belongs to the List. This predicate is re-executable on backtracking and can be thus used to enumerate the elements of List. memberchk/2 is similar to member/2 but only succeeds once. Errors None. Portability GNU Prolog predicate.

7.20.3 reverse/2

Templates

     reverse(?list, ?list)

Description reverse(List1, List2) succeeds if List2 unifies with the list List1 in reverse order. Errors None. Portability GNU Prolog predicate.

7.20.4 delete/3, select/3

Templates

     delete(?list, ?term, ?list)
     select(?term, ?list, ?list)

Description

delete(List1, Element, List2) removes all occurrences of Element in List1 to provide List2. A

strict term equality is required, cf. (==)/2 (section 7.3.2, page 51).

select(Element, List1, List2) removes one occurrence of Element in List1 to provide List2. This

predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

7.20.5 permutation/2

Templates

     permutation(?list, ?list)

Description permutation(List1, List2) succeeds if List2 is a permutation of the elements of List1. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

7.20.6 prefix/2, suffix/2

Templates

     prefix(?list, ?list)
     suffix(?list, ?list)

Description prefix(Prefix, List) succeeds if Prefix is a prefix of List. This predicate is re-executable on backtracking. suffix(Suffix, List) succeeds if Suffix is a suffix of List. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

7.20.7 sublist/2

Templates

     sublist(?list, ?list)

Description sublist(List1, List2) succeeds if List2 is a sub-list of List1. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

7.20.8 last/2

Templates

     last(?list, ?term)

Description

last(List, Element) succeeds if Element is the last element of List.

Errors None. Portability GNU Prolog predicate.

7.20.9 length/2

Templates

     length(?list, ?integer)

Description

length(List, Length) succeeds if Length is the length of List.

Errors None. Portability GNU Prolog predicate.

7.20.10 nth/3 Templates

     nth(?integer, ?list, ?term)

Description

nth(N, List, Element) succeeds if the Nth argument of List is Element.

Errors None. Portability GNU Prolog predicate.

7.20.11 max list/2, min list/2, sum list/2

Templates

     min list(+list, ?number)
     max list(+list, ?number)
     sum list(+list, ?number)

Description min list(List, Min) succeeds if Min is the smallest number in List. max list(List, Max) succeeds if Max is the largest number in List. sum list(List, Sum) succeeds if Sum is the sum of all the elements in List. List must be a list of arithmetic evaluable terms (section 7.6.1, page 57). Errors None. Portability GNU Prolog predicate.

7.20.12 sort/2, sort0/2, keysort/2 sort/1, sort0/1, keysort/1

Templates

     sort(+list, ?list)
     sort0(+list, ?list)
     keysort(+list, ?list)
     sort(+list)
     sort0(+list)
     keysort(+list)

Description sort(List1, List2) succeeds if List2 is the sorted list corresponding to List1 where duplicate elements are merged. sort0/2 is similar to sort/2 except that duplicate elements are not merged. keysort(List1, List2) succeeds if List2 is the sorted list of List1 according to the keys. The list List1 consists of items of the form Key-Value. These items are sorted according to the value of Key yielding the List2. Duplicate keys are not merged. This predicate is stable, i.e. if K-A occurs before K-B in the input, then K-A will occur before K-B in the output.

sort/1, sort0/1 and keysort/1 are similar to sort/2, sort0/2 and keysort/2 but achieve a sort in-

place destructing the original List1 (this in-place assignment is not undone at backtracking). The sorted list occupies the same memory space as the original list (saving thus memory consumption). The time complexity of these sorts is O(N log N), N being the length of the list to sort. These predicates refer to the standard ordering of terms (section 7.3.1, page 51).

Errors

 List1 is a partial list  instantiation error

List1 is neither a partial list nor a list type error(list, List1) List2 is neither a partial list nor a list type error(list, List2)

Portability

GNU Prolog predicates.

7.21 Global variables

7.21.1 Introduction

GNU Prolog provides a simple and powerful way to assign and read global variables. A global variable is associated to each atom, its initial value is the integer 0. A global variable can store 3 kinds of objects:

• a copy of a term (the assignment can be made backtrackable or not).

• a link to a term (the assignment is always backtrackable).

• an array of objects (recursively).

The space necessary for copies and arrays is dynamically allocated and recovered as soon as possible. For instance, when an atom is associated to a global variable whose current value is an array, the space for this array is recovered (unless the assignment is to be undone when backtracking occurs).

When a link to a term is associated to a global variable, the reference to this term is stored and thus the original term is returned when the content of the variable is read.

Global variable naming convention: a global variable is referenced by an atom.

If the variable contains an array, an index (ranging from 0) can be provided using a compound term whose principal functor is the correponding atom and the argument is the index. In case of a multi-dimensional array, each index is given as the arguments of the compound term.

If the variable contains a term (link or copy), it is possible to only reference a sub-term by giving its argument number (also called argument selector). Such a sub-term is specified using a compound term whose principal functor is -/2 and whose first argument is a global variable name and the second argument is the argument number (from 1). This can be applied recursively to specify a sub-term of any depth. In case of a list, a argument number I represents the Ith element of the list. In the rest of this section we use the operator notation since - is a predefined infix operator (section 7.14.10, page 100).

In the following, GVarName represents a reference to a global variable and its syntax is as follows:

GVarName ::= atom  whole content of a variable

                    atom(Integer,. . . ,Integer) element of an array

GVarName-Integer  sub-term selection
Integer ::= integer  immediate value
GVarName  indirect value

When a GVarName is used as an index or an argument number (i.e. indirection), the value of this variable must be an integer.

Here are some examples of the naming convention:

a  the content of variable associated to a (any kind)
t(1) the 2nd element of the array associated to t
t(k) if the value associated to k is I, the Ith element of the array associated to t
a-1-2 if the value associated to a is f(g(a,b,c),2), the sub-term b

Here are the errors associated to global variable names and common to all predicates.

 GVarName is a variable  instantiation error
 GVarName is neither a variable nor a callable term type error(callable, GVarName)

GVarName contains an invalid argument number (or GVarName is an array)

                                             domain error(g argument selector,
                                             GVarName)

GVarName contains an invalid index (or GVarName is not an array)

                                             domain error(g array index, GVarName)

GVarName is used as an indirect index or argument selector and is not an integer

                                             type error(integer, GVarName)

Arrays: the predicates g assign/2, g assignb/2 and g link/2 (section 7.21.2, page 127) can be used to create an array. They recognize some terms as values. For instance, a compound term with principal functor g array is used to define an array of fixed size. There are 3 forms for the term g array:

• g array(Size): if Size is an integer > 0 then defines an array of Size elements which are all initialized with the integer 0.

• g array(Size, Initial): as above but the elements are initialized with the term Initial instead of 0. Initial can contain other array definitions allowing thus for multi-dimensional arrays.

• g array(List): as above if List is a list of length Size except that the elements of the array are initialized according to the elements of List (which can contain other array definitions).

An array can be extended explicitely using a compound term with principal functor g array extend which accept the same 3 forms detailed above. In that case, the existing elements of the array are not initialized. If g array extend is used with an object which is not an array it is similar to g array.

Finally, an array can be automatically expanded when needed. The programmer does not need to explicitely control the expansion of an automatic array. An array is expanded as soon as an index is outside the current size of this array. Such an array is defined using a compound term with principal

functor g array auto:

• g array auto(Size): if Size is an integer > 0 then defines an automatic array whose initial size is Size. All elements are initialized with the integer 0. Elements created during implicit expansions will be initialized with 0.

• g array auto(Size, Initial): as above but the elements are initialized with the term Initial instead of 0. Initial can contain other array definitions allowing thus for multi-dimensional arrays. Elements created during implicit expansions will be initialized with Initial.

• g array auto(List): as above if List is a list of length Size except that the elements of the array are initialized according to the elements of List (which can contain other array definitions). Elements created during implicit expansions will be initialized with 0.

In any case, when an array is read, a term of the form g array([Elem0,..., ElemSize-1]) is returned.

Some examples using global variables are presented later (section 7.21.7, page 130).

7.21.2 g assign/2, g assignb/2, g link/2

     g assign(+callable term, ?term)
     g assignb(+callable term, ?term)
     g link(+callable term, ?term)

Description g assign(GVarName, Value) assigns a copy of the term Value to GVarName. This assignment is not undone when backtracking occurs. g assignb/2 is similar to g assign/2 but the assignment is undone at backtracking. g link(GVarName, Value) makes a link between GVarName to the term Value. This allows the user to give a name to any Prolog term (in particular non-ground terms). Such an assignment is always undone when backtracking occurs (since the term may no longer exist). If Value is an atom or an integer, g link/2 and g assignb/2 have the same behavior. Since g link/2 only handles links to existing terms it does not require extra memory space and is not expensive in terms of execution time. NB: argument selectors can only be used with g assign/2 (i.e. when using an argument selector inside an assignment, this one must not be backtrackable). Errors See common errors detailed in the introduction (section 7.21.1, page 126) GVarName contains an argument selector and the assignment is backtrackable

                                             domain error(g argument selector,
                                             GVarName)

Portability GNU Prolog predicates.

7.21.3 g read/2

Templates

     g read(+callable term, ?term)

Description

g read(GVarName, Value) unifies Value with the term assigned to GVarName.

Errors See common errors detailed in the introduction (section 7.21.1, page 126) Portability GNU Prolog predicate.

7.21.4 g array size/2

Templates

     g array size(+callable term, ?integer)

Description g array size(GVarName, Value) unifies Size with the dimension (an integer > 0) of the array assigned to GVarName. Fails if GVarName is not an array. Errors See common errors detailed in the introduction (section 7.21.1, page 126) Size is neither a variable nor an integer type error(integer, Size) Portability GNU Prolog predicate.

7.21.5 g inc/3, g inc/2, g inco/2, g inc/1, g dec/3, g dec/2, g deco/2, g dec/1

Templates

     g inc(+callable term, ?integer, ?integer)
     g inc(+callable term, ?integer)
     g inco(+callable term, ?integer)
     g inc(+callable term)
     g dec(+callable term, ?integer, ?integer)
     g dec(+callable term, ?integer)
     g deco(+callable term, ?integer)
     g dec(+callable term)

Description

g inc(GVarName, Old, New) unifies Old with the integer assigned to GVarName, increments GVarName

and then unifies New with the incremented value.

g inc(GVarName, New) is equivalent to g inc(GVarName,  , New).
g inco(GVarName, Old) is equivalent to g inc(GVarName, Old,  ).
g inc(GVarName) is equivalent to g inc(GVarName,  ,  ).

Predicates g dec are similar but decrement the content of GVarName instead. Errors See common errors detailed in the introduction (section 7.21.1, page 126) Old is neither a variable nor an integer type error(integer, Old) New is neither a variable nor an integer type error(integer, New)

 GVarName stores an array  type error(integer, g array)

GVarName stores a term T which is not an integer type error(integer, T) Portability GNU Prolog predicates.

7.21.6 g set bit/2, g reset bit/2, g test set bit/2, g test reset bit/2

Templates

     g set bit(+callable term, +integer)
     g reset bit(+callable term, +integer)
     g test set bit(+callable term, +integer)
     g test reset bit(+callable term, +integer)

Description

g set bit(GVarName, Bit) sets to 1 the bit number specified by Bit of the integer assigned to GVarName to 1. Bit numbers range from 0 to the maximum number allowed for integers (this is architecture dependent). If Bit is greater than this limit, the modulo with this limit is taken.

g reset bit(GVarName, Bit) is similar to g set bit/2 but sets the specified bit to 0.

g test set bit/2 succeeds if the specified bit is set to 1.

g test reset bit/2 succeeds if the specified bit is set to 0.

Errors

See common errors detailed in the introduction (section 7.21.1, page 126)

 Bit is a variable  instantiation error

Bit is neither a variable nor an integer type error(integer, Bit)

 Bit is an integer < 0  domain error(not less than zero, Bit)
 GVarName stores an array  type error(integer, g array)

GVarName stores a term T which is not an integer type error(integer, T)

Portability

GNU Prolog predicates.

7.21.7 Examples

Simulating g inc/3: this predicate behaves like: global variable:

     my_g_inc(Var, Old, New) :-
             g_read(Var, Old),
             N is Value + 1,
             g_assign(Var, X),
     New = N.

The query: my g inc(c, X, ) will succeed unifying X with 0, another call to my g inc(a, Y, ) will then unify Y with 1, and so on.

Difference between g assign/2 and g assignb/2: g assign/2 does not undo its assignment when backtracking occurs whereas g assignb/2 undoes it.

     test(Old) :-                       testb(Old) :-
             g assign(x,1),                     g assign(x,1),
             (   g read(x, Old),                (   g read(x, Old),
                 g assign(x, 2)                    g assignb(x, 2)
             ;   g read(x, Old),                ;   g read(x, Old),
                 g assign(x, 3)                     g assign(x, 3)
             ).                                 ).

The query test(Old) will succeed unifying Old with 1 and on backtracking with 2 (i.e. the assignment of the value 2 has not been undone). The query testb(Old) will succeed unifying Old with 1 and on backtracking with 1 (i.e. the assignment of the value 2 has been undone). Difference between g assign/2 and g link/2: g assign/2 (and g assignb/2) creates a copy of the term whereas g link/2 does not. g link/2 can be used to avoid passing big data structures (e.g. dictionaries,…) as arguments to predicates.

test(B) :-                         test(B) :-
        g assign(b, f(X)),                g link(b, f(X)),
        X = 12,                            X = 12,
        g read(b, B).                      g read(b, B).

The query test(B) will succeed unifying B with f( ) (g assign/2 assigns a copy of the value). The query testl(B) will succeed unifying B with f(12) (g link/2 assigns a pointer to the term). Simple array definition: here are some queries to show how arrays can be handled:

| ?- g_assign(w, g_array(3)), g_read(w, X).
X = g_array([0,0,0])
| ?- g_assign(w(0), 16), g_assign(w(1), 32), g_assign(w(2), 64), g_read(w, X).
X = g_array([16,32,64])

this is equivalent to:

| ?- g_assign(k, g_array([16,32,64])), g_read(k, X).
X = g_array([16,32,64])
| ?- g_assign(k, g_array(3,null)), g_read(k, X), g_array_size(k, S).
S = 3
X = g_array([null,null,null])

2-D array definition:

| ?- g_assign(w, g_array(2, g_array(3))), g_read(w, X).
X = g_array([g_array([0,0,0]),g_array([0,0,0])])
| ?- (   for(I,0,1), for(J,0,2), K is I*3+J, g_assign(w(I,J), K),
         fail
     ;   g_read(w, X)
     ).
X = g_array([g_array([0,1,2]),g_array([3,4,5])])
| ?- g_read(w(1),X).
X = g_array([3,4,5])

Hybrid array:

| ?- g_assign(w,g_array([1,2,g_array([a,b,c]), g_array(2,z),5])), g_read(w, X).
X = g_array([1,2,g_array([a,b,c]), g_array([z,z]),5])

| ?- g_read(w(1), X), g_read(w(2,1), Y), g_read(w(3,1), Z).
X = 2
Y = b
Z = z
| ?- g_read(w(1,2),X).
uncaught exception: error(domain_error(g_array_index,w(1,2)),g_read/2)

Array extension:

| ?- g_assign(a, g_array([10,20,30])), g_read(a, X).
X = g_array([10,20,30])
| ?- g_assign(a, g_array_extend(5,null)), g_read(a, X).
X = g_array([10,20,30,null,null])
| ?- g_assign(a, g_array([10,20,30])), g_read(a, X).
X = g_array([10,20,30])
| ?- g_assign(a, g_array_extend([1,2,3,4,5,6])), g_read(a, X).
X = g_array([10,20,30,4,5,6])

Automatic array:

| ?- g_assign(t, g_array_auto(3)), g_assign(t(1), foo), g_read(t,X).
X = g_array([0,foo,0])
| ?- g_assign(t(5), bar), g_read(t,X).
X = g_array([0,foo,0,0,0,bar,0,0])
| ?- g_assign(t, g_array_auto(2, g_array(2))), g_assign(t(1,1), foo),
     g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo])])
| ?- g_assign(t(3,0), bar), g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo]),g_array([0,0]),g_array([bar,0])])
| ?- g_assign(t(3,4), bar), g_read(t,X).
uncaught exception: error(domain_error(g_array_index,t(3,4)),g_assign/2)
| ?- g_assign(t, g_array_auto(2, g_array_auto(2))), g_assign(t(1,1), foo),
     g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo])])
| ?- g_assign(t(3,3), bar), g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo]),g_array([0,0]),
    g_array([0,0,0,bar])])

     | ?- g_assign(t, g_array_auto(2, g_array_auto(2, null))), g_read(t(2,3), U),
          g_read(t, X).
     U = null
     X = g_array([g_array([null,null]),g_array([null,null]),
                  g_array([null,null,null,null]),g_array([null,null])])

7.22 Prolog state

7.22.1 set prolog flag/2

Templates

     set prolog flag(+flag, +term)

Description

set prolog flag(Flag, Value) sets the value of the Prolog flag Flag to Value.

Prolog flags: a Prolog flag is an atom which is associated with a value that is either implementation defined or defined by the user. Each flag has a permitted range of values; any other value is a domain error. The following two tables present available flags, the possible values, a description and if they are ISO or an extension. The first table presents unchangeable flags while the second one the changeable flags. For flags whose default values is machine independent, this value is \underline{underlined}.

Unchangeable flags:

Flag  Values  Description  ISO

 bounded

                            \underline{true} / false

are integers bounded ?

Y

 max integer

an integer greatest integer

Y

 min integer

an integer smallest integer

Y

 integer rounding function

                             toward zero down

rnd(X) = integer part of X

rnd(X) = ⌊X⌋ (section 7.6.1, page 57)

Y

 max arity

an integer maximum arity for compound terms (255)

Y

 max atom

an integer maximum number of atoms

N

 max unget

an integer maximum number of successive ungets

N

 prolog name

an atom name of the Prolog system

N

 prolog version

an atom version number of the Prolog system

N

 prolog date

an atom date of the Prolog system

N

 prolog copyright

an atom copyright message of the Prolog system

N

Changeable flags:

Flag  Values  Description  ISO

char conversion

on / \underline{off}

is character conversion activated ?

Y

debug

on / \underline{off}

is the debugger activated ?

Y

singleton warning

\underline{on} / off

warn about named singleton variables ?

N

strict iso

\underline{on} / off

strict ISO behavior ?

N

double quotes                    atom
chars
\underline{codes}
atom no escape
chars no escape
codes no escape

a double quoted constant is returned as: an atom a list of characters a list of character codes as atom but ignore escape sequences as chars but ignore escape sequences as code but ignore escape sequences

{Y \atop N}

back quotes

atom
chars
codes
atom no escape
chars no escape
codes no escape

a back quoted constant is returned as: an atom a list of characters a list of character codes as atom but ignore escape sequences as chars but ignore escape sequences as code but ignore escape sequences

N

unknown

\underline{error} warning fail

a predicate calls an unknown procedure:

an existence error is raised

a message is displayed then fails quietly fails

Y

syntax error

\underline{error} warning fail

a predicate causes a syntax error:

a syntax error is raised

a message is displayed then fails quietly fails

N

os error

\underline{error} warning fail

a predicate causes an O.S. error:

a system error is raised

a message is displayed then fails quietly fails

N

The strict iso flag is introduced to allow a compatibility with other Prolog systems. When turned off the following relaxations apply:

• a callable term can be given as a predicate indicator.

• built-in predicates are found by current predicate/1 (section 7.8.1, page 64).

Errors

Flag is a variable  instantiation error
Value is a variable  instantiation error

Flag is neither a variable nor an atom type error(atom, Flag)

Flag is an atom but not a valid flag  domain error(prolog flag, Flag)
Value is inappropriate for Flag  domain error(flag value, Flag+Value)

Value is appropriate for Flag but flag Flag is not modifiable

permission error(modify, flag, Flag)

Portability

ISO predicate. All ISO flags are implemented.

7.22.2 current prolog flag/2

     current prolog flag(?flag, ?term)

Description current prolog flag(Flag, Value) succeeds if there exists a Prolog flag that unifies with Flag and whose value unifies with Value. This predicate is re-executable on backtracking. Errors Flag is neither a variable nor an atom type error(atom, Flag)

 Flag is an atom but not a valid flag  domain error(prolog flag, Flag)

Portability ISO predicate.

7.22.3 set bip name/2

Templates

     set bip name(+atom, +arity)

Description set bip name(Functor, Arity) initializes the context of the error (section 5.3.1, page 37) with Functor and Arity (if Arity < 0 only Functor is significant). Errors

 Functor is a variable  instantiation error
 Arity is a variable  instantiation error
 Functor is neither a variable nor an atom  type error(atom, Functor)

Arity is neither a variable nor an integer type error(integer, Arity) Portability GNU Prolog predicate.

7.22.4 current bip name/2

Templates

     current bip name(?atom, ?arity)

Description

current bip name(Functor, Arity) succeeds if Functor and Arity correspond to the context of the

error (section 5.3.1, page 37) (if Arity < 0 only Functor is significant). Errors

 Functor is neither a variable nor an atom  type error(atom, Functor)

Arity is neither a variable nor an integer type error(integer, Arity) Portability GNU Prolog predicate.

7.22.5 write pl state file/1, read pl state file/1

Templates

     write pl state file(+source sink)
     read pl state file(+source sink)

Description

write pl state file(FileName) writes onto FileName all information that influences the parsing of a term (section 7.14, page 91). This allows a sub-process written in Prolog to read this file and then process any Prolog term as done by the parent process. This file can also be passed as argument of the --pl-state option when invoking gplc (section 3.4.3, page 22). More precisely the following elements are saved:

• all operator definitions (section 7.14.10, page 100).

• the character conversion table (section 7.14.12, page 102).

   • the value of char conversion, double quotes, back quotes and singleton warning Prolog flags

(section 7.22.1, page 133).

read pl state file(FileName) reads (restores) from FileName all information previously saved by

write pl state file/1.

Errors

 FileName is a variable  instantiation error
 FileName is neither a variable nor an atom  type error(atom, FileName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.23 Program state

7.23.1 consult/1, '.'/2 - program consult

Templates

     consult(+atom or atom list)
     '.'(+atom, +atom list)

Description

consult(Files) compiles and loads into memory each file of the list Files. Each file is compiled for byte-code using the GNU Prolog compiler (section 3.4, page 20) then loaded using load/1 (section 7.23.2, page 137). It is possible to specify user as a file name to directly enter the program from the terminal. Files can be also a single file name (i.e. an atom). Refer to the section concerning the consult of a Prolog program for more information (section 3.2.3, page 16).

The final file name of a file is computed using the predicates prolog file name/2 (section 7.26.3, page 142) and absolute file name/2 (section 7.26.1, page 141).

[ File | Files ], i.e. '.'(File, Files) is equivalent to consult([ File | Files ]).

Errors

Files is a partial list or a list with an element E which is a variable

                                           instantiation error

Files is neither a partial list nor a list nor an atom

                                           type error(list, Files)

an element E of the Files list is neither a variable nor an atom

                                           type error(atom, E)

an element E of the Files list is an atom but not a valid pathname

                                           domain error(os path, E)

an element E of the Files list is a valid pathname but does not correspond to an existing source

                                           existence error(source sink, E)

an error occurs executing a directive see call/1 errors (section 6.2.3, page 47)

Portability

GNU Prolog predicates.

7.23.2 load/1

Templates

    load(+atom or atom list)

Description

load(Files) loads into memory each file of the list Files. Each file must have been previously compiled for byte-code using the GNU Prolog compiler (section 3.4, page 20). Files can be also a single file name (i.e. an atom).

The final file name of a file is computed using the predicates absolute file name/2 (section 7.26.1, page 141). If no suffix is given '.wbc' is appended to the file name.

Errors

Files is a partial list or a list with an element E which is a variable

                                           instantiation error

Files is neither a partial list nor a list nor an atom

                                           type error(list, Files)

an element E of the Files list is neither a variable nor an atom

                                           type error(atom, E)

an element E of the Files list is an atom but not a valid pathname

                                           domain error(os path, E)

an element E of the Files list is a valid pathname but does not correspond to an existing source

                                           existence error(source sink, E)

an error occurs executing a directive see call/1 errors (section 6.2.3, page 47)

Portability

7.23.3 listing/1, listing/0

Templates

     listing(+predicate indicator)
     listing(+atom)
     listing

Description

listing(Pred) lists the clauses of the consulted predicate whose predicate indicator is Pred. Pred can also be a single atom in which case all predicates whose name is Pred are listed (of any arity). This predicate uses portray clause/2 (section 7.14.8, page 99) to output the clauses.

listing lists all clauses of all consulted predicates.

Errors

 Pred is a variable  instantiation error

Pred is neither a variable nor predicate indicator or an atom

                                             type error(predicate indicator, Pred)

Portability

GNU Prolog predicate.

7.24 System statistics

7.24.1 statistics/0, statistics/2

Templates

     statistics
     statistics(?atom, ?list)

Description

statistics displays statistics about memory usage and run times.

statistics(Key, Value) unifies Value with the current value of the statistics key Key. Value a list of two elements. Times are in milliseconds, sizes of areas in bytes.

Key  Description  Value

 user time  user CPU time  [SinceStart, SinceLast]
 system time system CPU time  [SinceStart, SinceLast]
 cpu time  total CPU time (user + system)  [SinceStart, SinceLast]
 real time  absolute time  [SinceStart, SinceLast]

local stack local stack sizes (control, environments, choices) [UsedSize, FreeSize]
global stack global stack sizes (compound terms)  [UsedSize, FreeSize]
trail stack trail stack sizes (variable bindings to undo)  [UsedSize, FreeSize]
cstr stack  constraint trail sizes (finite domain constraints) [UsedSize, FreeSize]

Note that the key runtime is recognized as user time for compatibility purpose.

 Key is neither a variable nor a valid key  domain error(statistics key, Key)

Value is neither a variable nor a list of two elements

                                             domain error(statistics value, Value)

Value is a list of two elements and an element E is neither a variable nor an integer

                                             type error(integer, E)

Portability GNU Prolog predicates.

7.24.2 user time/1, system time/1, cpu time/1, real time/1

Templates

     user time(?integer)
     system time(?integer)
     cpu time(?integer)
     real time(?integer)

Description user time(Time) unifies Time with the user CPU time elapsed since the start of Prolog. system time(Time) unifies Time with the system CPU time elapsed since the start of Prolog. cpu time(Time) unifies Time with the CPU time (user + system) elapsed since the start of Prolog. real time(Time) unifies Time with the absolute time elapsed since the start of Prolog. Errors Time is neither a variable nor an integer type error(integer, Time) Portability GNU Prolog predicates.

7.25 Random number generator

7.25.1 set seed/1, randomize/0

Templates

     set seed(+integer)
     randomize

Description set seed(Seed) reinitializes the random number generator seed with Seed. randomize reinitializes the random number generator. This predicates calls set seed/1 with a random value depending on the absolute time. Errors

 Seed is a variable  instantiation error

Seed is neither a variable nor an integer type error(integer, Seed)

 Seed is an integer < 0  domain error(not less than zero, Seed)

Portability GNU Prolog predicates.

7.25.2 get seed/1

Templates

     get seed(?integer)

Description get seed(Seed) unifies Seed with the current random number generator seed. Errors Seed is neither a variable nor an integer type error(integer, Seed)

 Seed is an integer < 0  domain error(not less than zero, Seed)

Portability GNU Prolog predicate.

7.25.3 random/1

Templates

     random(-float)

Description random(Number) unifies Number with a random floating point number such that 0.0 ≤ Number < 1.0. Errors

 Number is not a variable  type error(variable, Number)

Portability GNU Prolog predicate.

7.25.4 random/3

Templates

     random(+number, +number, -number)

Description

random(Base, Max, Number) unifies Number with a random number such that Base ≤ Number < Max.

If both Base and Max are integers Number will be an integer, otherwise Number will be a floating point number.

Errors

 Base is a variable  instantiation error

Base is neither a variable nor a number type error(number, Base)

 Max is a variable  instantiation error

Max is neither a variable nor a number type error(number, Max)

 Number is not a variable  type error(variable, Number)

Portability

GNU Prolog predicate.

7.26 File name processing

7.26.1 absolute file name/2

Templates

     absolute file name(+atom, atom)

Description

absolute file name(File1, File2) succeeds if File2 is the absolute pathname associated to the relative file name File1. File1 can contain $VAR NAME sub-strings. When such a sub-string is encountered, it is expanded with the value of the environment variable whose name is VAR NAME if exists (otherwise no expansion is done). File1 can also begin with a sub-string ~USER NAME/, this is expanded as the home directory of the user USER NAME . If USER NAME does not exist File1 is an invalid pathname. If no USER NAME is given (i.e. File1 begins with ~/) the ~ character is expanded as the value of the environment variable HOME. If the HOME variable is not defined File1 is an invalid pathname. Relative references to the current directory (/./ sub-string) and to the parent directory (/../ sub-strings) are removed and no longer appear in File2. File1 is also invalid if it contains too many /../ consecutive sub-strings (i.e. parent directory relative references). Finally if File1 is user then File2 is also unified with user to allow this predicate to be called on Prolog file names (since user in DEC-10 input/output predicates denotes the current input/output stream).

Most predicates using a file name implicitly call this predicate to obtain the desired file, e.g. open/4.

Errors

 File1 is a variable  instantiation error

File1 is neither a variable nor an atom type error(atom, File1) File2 is neither a variable nor an atom type error(atom, File2)

 File1 is an atom but not a valid pathname  domain error(os path, File1)

Portability

GNU Prolog predicate.

7.26.2 decompose file name/4

Templates

     decompose file name(+atom, ?atom, ?atom, ?atom)

Description

decompose file name(File, Directory, Prefix, Suffix) decomposes the pathname File and ex-

tracts the Directory part (characters before the last /), the Prefix part (characters after the last / and before the last . or until the end if there is no suffix) and the Suffix part (characters from the last . to the end of the string). Errors

 File is a variable  instantiation error

File is neither a variable nor an atom type error(atom, File)

 Directory is neither a variable nor an atom  type error(atom, Directory)
 Prefix is neither a variable nor an atom  type error(atom, Prefix)
 Suffix is neither a variable nor an atom  type error(atom, Suffix)

Portability GNU Prolog predicate.

7.26.3 prolog file name/2

Templates

     prolog file name(+atom, ?atom)

Description prolog file name(File1, File2) unifies File2 with the Prolog file name associated to File1. More precisely File2 is computed as follows:

• if File1 has a suffix or if it is user then File2 is unified with File1.

• else if the file whose name is File1 + '.pl' exists then File2 is unified with this name.

• else if the file whose name is File1 + '.pro' exists then File2 is unified with this name.

• else File2 is unified with the name File1 + '.pl'. This predicate uses absolute file name/2 to check the existence of a file (section 7.26.1, page 141). Errors

 File1 is a variable  instantiation error

File1 is neither a variable nor an atom type error(atom, File1) File2 is neither a variable nor an atom type error(atom, File2)

 File1 is an atom but not a valid pathname  domain error(os path, File1)

Portability GNU Prolog predicate.

7.27 Operating system interface

7.27.1 argument counter/1

Templates

     argument counter(?integer)

Description argument counter(Counter) succeeds if Counter is the number of arguments of the command-line. Since the first argument is always the name of the running program, Counter is always ≥ 1. See (section 3.2, page 13) for more information about command-line arguments retrieved under the top level. Errors

 Counter is neither a variable nor an integer  type error(integer, Counter)

Portability GNU Prolog predicate.

7.27.2 argument value/2

Templates

     argument value(+integer, ?atom)

Description argument value(N, Arg) succeeds if the Nth argument on the command-line unifies with Arg. The first argument is always the name of the running program and its number is 0. The number of arguments on the command-line can be obtained using argument counter/1 (section 7.27.1, page 142). Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)

Arg is neither a variable nor an atom type error(atom, Arg) Portability GNU Prolog predicate.

7.27.3 argument list/1

Templates

     argument list(?list)

Description argument list(Args) succeeds if Args unifies with the list of atoms associated to each argument on the command-line other than the first argument (the name of the running program). Errors Args is neither a partial list nor a list type error(list, Args) Portability GNU Prolog predicate.

7.27.4 environ/2

Templates

     environ(?atom, ?atom)

Description environ(Name, Value) succeeds if Name is the name of an environment variable whose value is Value. This predicate is re-executable on backtracking. Errors Name is neither a variable nor an atom type error(atom, Name) Value is neither a variable nor an atom type error(atom, Value) Portability GNU Prolog predicate.

7.27.5 make directory/1, delete directory/1, change directory/1

Templates

     make directory(+atom)
     delete directory(+atom)
     change directory(+atom)

Description make directory(PathName) creates the directory whose pathname is PathName. delete directory(PathName) removes the directory whose pathname is PathName. change directory(PathName) sets the current directory to the directory whose pathname is PathName. See absolute file name/2 for information about the syntax of PathName (section 7.26.1, page 141). Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability GNU Prolog predicates.

7.27.6 working directory/1

Templates

     working directory(?atom)

Description working directory(PathName) succeeds if PathName is the pathname of the current directory. Errors

 PathName is neither a variable nor an atom  type error(atom, PathName)

Portability GNU Prolog predicate.

7.27.7 directory files/2

Templates

     directory files(+atom, ?list)

Description directory files(PathName, Files) succeeds if Files is the list of all entries (files, sub-directories,…) in the directory whose pathname is PathName. See absolute file name/2 for information about the syntax of PathName (section 7.26.1, page 141). Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

Files is neither a partial list nor a list type error(list, Files) an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

7.27.8 rename file/2

Templates

     rename file(+atom, +atom)

Description rename file(PathName1, PathName2) renames the file or directory whose pathname is PathName1 to

PathName2. See absolute file name/2 for information about the syntax of PathName1 and PathName2

(section 7.26.1, page 141). Errors

 PathName1 is a variable  instantiation error
 PathName1 is neither a variable nor an atom  type error(atom, PathName1)
 PathName1 is an atom but not a valid pathname domain error(os path, PathName1)
 PathName2 is a variable  instantiation error
 PathName2 is neither a variable nor an atom  type error(atom, PathName2)
 PathName2 is an atom but not a valid pathname domain error(os path, PathName2)

an operating system error occurs and value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.9 delete file/1, unlink/1

Templates

     delete file(PathName)
     unlink(PathName)

Description

delete file(PathName) removes the existing file whose pathname is PathName.

unlink/1 is similar to delete file/1 except that it never causes a system error (e.g. if PathName does not refer to an existing file).

See absolute file name/2 for information about the syntax of PathName (section 7.26.1, page 141).

Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

7.27.10 file permission/2, file exists/1

Templates

     file permission(+atom, +atom)
     file permission(+atom, +atom list)
     file exists(+atom)

Description

file permission(PathName, Permission) succeeds if PathName is the pathname of an existing file (or

directory) whose permissions include Permission.

File permissions: Permission can be a single permission or a list of permissions. A permission is an atom among:

• read: the file or directory can be read.

• write: the file or directory can be written.

• execute: the file can be executed.

• search: the directory can be searched.

If PathName does not exists or if it its permissions do not include Permission this predicate fails.

file exists(PathName) is equivalent to file permission(PathName, []), i.e. it succeeds if PathName

is the pathname of an existing file (or directory).

See absolute file name/2 for information about the syntax of PathName (section 7.26.1, page 141).

Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

Permission is a partial list or a list with an element which is a variable

                                             instantiation error

Permission is neither an atom nor partial list or a list

                                             type error(list, Permission)

an element E of the Permission list is neither a variable nor an atom

                                             type error(atom, E)

an element E of the Permission is an atom but not a valid permission

                                             domain error(os file permission,
                                             Permission)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

7.27.11 file property/2

Templates

     file property(+atom, ?os file property)

Description

file property(PathName, Property) succeeds if PathName is the pathname of an existing file (or directory) and if Property unifies with one of the properties of the file. This predicate is re-executable on backtracking.

File properties:

• absolute file name(File): File is the absolute file name of PathName (section 7.26.1, page 141).

• real file name(File): File is the real file name of PathName (follows symbolic links).

   • type(Type): Type is the type of PathName. Possible values are: regular, directory, fifo, socket,
     character device, block device or unknown.

• size(Size): Size is the size (in bytes) of PathName.

• permission(Permission): Permission is a permission of PathName (section 7.27.10, page 146).

• last modification(DT): DT is the last modification date and time (section 7.27.14, page 149).

See absolute file name/2 for information about the syntax of PathName (section 7.26.1, page 141).

Errors

PathName is a variable  instantiation error
PathName is neither a variable nor an atom  type error(atom, PathName)
PathName is an atom but not a valid pathname  domain error(os path, PathName)

Property is neither a variable nor a file property term

                                           domain error(os file property, Property)
Property = absolute file name(E),
real file name(E), type(E) or permission(E)

and E is neither a variable nor an atom

                                           type error(atom, E)

Property = last modification(DateTime)

and DateTime is neither a variable nor a compound term

                                           type error(compound, DateTime)

Property = last modification(DateTime)

and DateTime is a compound term but not a structure dt/6

                                           domain error(date time, DateTime)

Property = size(E) or
last modification(DateTime) and DateTime is

a structure dt/6 but an element E is neither a variable nor an integer

                                           type error(integer, E)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.12 temporary name/2

Templates

    temporary name(+atom, ?atom)

Description

temporary name(Template, PathName) creates a unique file name PathName whose pathname begins by Template. Template should contain a pathname with six trailing Xs. PathName is Template with the six Xs replaced with a letter and the process identifier. This predicate is an interface to the C Unix function

mktemp(3).

See absolute file name/2 for information about the syntax of Template (section 7.26.1, page 141).

 Template is a variable  instantiation error
 Template is neither a variable nor an atom  type error(atom, Template)
 Template is an atom but not a valid pathname  domain error(os path, Template)
 PathName is neither a variable nor an atom  type error(atom, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.13 temporary file/3

Templates

     temporary file(+atom, +atom, ?atom)

Description

temporary file(Directory, Prefix, PathName) creates a unique file name PathName whose pathname

begins by Directory/Prefix. If Directory is the empty atom '' a standard temporary directory will be used (e.g. /tmp). Prefix can be the empty atom ''. This predicate is an interface to the C Unix

function tempnam(3).

See absolute file name/2 for information about the syntax of Directory (section 7.26.1, page 141).

Errors

 Directory is a variable  instantiation error
 Directory is neither a variable nor an atom  type error(atom, Directory)
 Directory is an atom but not a valid pathname domain error(os path, Directory)
 Prefix is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)
 PathName is neither a variable nor an atom  type error(atom, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.14 date time/1

Templates

     date time(?compound)

Description

date time(DateTime) unifies DateTime with a compound term containing the current date and time.

DateTime is a structure dt(Year, Month, Day, Hour, Minute, Second). Each sub-argument of the

term dt/6 is an integer.

Errors

DateTime is neither a variable nor a compound term

                                             type error(compound, DateTime)

DateTime is a compound term but not a structure dt/6

                                             domain error(date time, DateTime)

DateTime is a structure dt/6 and an element E is neither a variable nor an integer

                                             type error(integer, E)

Portability

GNU Prolog predicate.

7.27.15 host name/1

Templates

     host name(?atom)

Description

host name(HostName) unifies HostName with the name of the host machine executing the current GNU Prolog process. If the sockets are available (section 7.28.1, page 157), the name returned will be fully qualified. In that case, host name/1 will also succeed if HostName is instantiated to the unqualified name (or an alias) of the machine.

Errors

 Hostname is neither a variable nor an atom  type error(atom, HostName)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.16 os version/1

Templates

     os version(?atom)

Description

os version(OSVersion) unifies OSVersion with the operating system version of the machine executing the current GNU Prolog process.

Errors

 OSVersion is neither a variable nor an atom  type error(atom, OSVersion)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

GNU Prolog predicate.

7.27.17 architecture/1

Templates

     architecture(?atom)

Description architecture(Architecture) unifies Architecture with the name of the machine executing the current GNU Prolog process. Errors

 Architecture is neither a variable nor an atom type error(atom, Architecture)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

7.27.18 shell/2, shell/1, shell/0

Templates

     shell(+atom, ?integer)
     shell(+atom)
     shell

Description shell(Command, Status) invokes a new shell (named by the SHELL environment variable) passing Command for execution and unifies Status with the result of the execution. If Command is the empty atom '' a new interactive shell is executed. The control is returned to Prolog upon termination of the called process.

shell(Command) is equivalent to shell(Command, 0).
shell is equivalent to shell('', 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Status is neither a variable nor an integer  type error(integer, Status)

Portability GNU Prolog predicates.

7.27.19 system/2, system/1

Templates

     system(+atom, ?integer)
     system(+atom)

Description

system(Command, Status) invokes a new default shell passing Command for execution and unifies Status with the result of the execution. The control is returned to Prolog upon termination of the shell process. This predicate is an interface to the C Unix function system(3).

system(Command) is equivalent to system(Command, 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Status is neither a variable nor an integer  type error(integer, Status)

Portability

GNU Prolog predicates.

7.27.20 spawn/3, spawn/2

Templates

     spawn(+atom, +atom list, ?integer)
     spawn(+atom, +atom list)

Description

spawn(Command, Arguments, Status) executes Command passing as arguments of the command-line each element of the list Arguments and unifies Status with the result of the execution. The control is returned to Prolog upon termination of the command.

spawn(Command, Arguments) is equivalent to spawn(Command, Arguments, 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)

Arguments is a partial list or a list with an element which is a variable

                                             instantiation error
 Arguments is neither a partial list nor a list  type error(list, Arguments)

an element E of the Arguments list is neither a variable nor an atom

                                             type error(atom, E)
 Status is neither a variable nor an integer  type error(integer, Status)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

7.27.21 popen/3

Templates

     popen(+atom, +io mode, -stream)

Description

popen(Command, Mode, Stream) invokes a new default shell (by creating a pipe) passing Command for execution and associates a stream either to the standard input or the standard output of the created process. if Mode is read (resp. write) an input (resp. output) stream is created and Stream is unified with the stream-term associated. Writing to the stream writes to the standard input of the command while reading from the stream reads the command’s standard output. The stream must be closed using close/2 (section 7.10.7, page 72). This predicate is an interface to the C Unix function popen(3).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Mode is a variable  instantiation error

Mode is neither a variable nor an atom type error(atom, Mode)

 Mode is an atom but neither read nor write.  domain error(io mode, Mode)
 Stream is not a variable  type error(variable, Stream)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.22 exec/5, exec/4

Templates

     exec(+atom, -stream, -stream, -stream, -integer)
     exec(+atom, -stream, -stream, -stream)

Description

exec(Command, StreamIn, StreamOut, StreamErr, Pid) invokes a new default shell passing Command

for execution and associates streams to standard streams of the created process. StreamIn is unified with the stream-term associated to the standard input stream of Command (it is an output stream). StreamOut is unified with the stream-term associated to the standard output stream of Command (it is an input stream). StreamErr is unified with the stream-term associated to the standard error stream of Command (it is an input stream). Pid is unified with the process identifier of the new process. This information is only useful if it is necessary to obtain the status of the execution using wait/2 (section 7.27.25, page 155). Until a call to wait/2 is done the process remains in the system processes table (as a zombie process if terminated). For this reason, if the status is not needed it is preferable to use exec/4.

exec/4 is similar to exec/5 but the process is removed from system processes as soon as it is terminated.

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 StreamIn is not a variable  type error(variable, StreamIn)
 StreamOut is not a variable  type error(variable, StreamOut)
 StreamErr is not a variable  type error(variable, StreamErr)
 Pid is not a variable  type error(variable, Pid)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

7.27.23 fork prolog/1

Templates

     fork prolog(-integer)

Description

fork prolog(Pid) creates a child process that differs from the parent process only in its PID. In the parent process Pid is unified with the PID of the child while in the child process Pid is unified with 0. In the parent process, the status of the child process can be ontained using wait/2 (section 7.27.25, page 155). Until a call to wait/2 is done the child process remains in the system processes table (as a zombie process if terminated). This predicate is an interface to the C Unix function fork(2).

Errors

 Pid is not a variable  type error(variable, Pid)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.24 create pipe/2

Templates

     create pipe(-stream, -stream)

Description

create pipe(StreamIn, StreamOut) creates a pair of streams pointing to a pipe inode. StreamIn is unified with the stream-term associated to the input side of the pipe and StreamOut is unified with the stream-term associated to output side. This predicate is an interface to the C Unix function pipe(2).

 StreamIn is not a variable  type error(variable, StreamIn)
 StreamOut is not a variable  type error(variable, StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

7.27.25 wait/2 Templates

     wait(+integer, ?integer)

Description wait(Pid, Status) waits for the child process whose identifier is Pid to terminate. Status is then unified with the exit status. This predicate is an interface to the C Unix function waitpid(2). Errors

 Pid is a variable  instantiation error

Pid is neither a variable nor an integer type error(integer, Pid)

 Status is neither a variable nor an integer  type error(integer, Status)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

7.27.26 prolog pid/1

Templates

     prolog pid(?integer)

Description prolog pid(Pid) unifies Pid with the process identifier of the current GNU Prolog process. Errors Pid is neither a variable nor an integer type error(integer, Pid) Portability GNU Prolog predicate.

7.27.27 send signal/2

Templates

     send signal(+integer, +integer)
     send signal(+integer, +atom)

Description

send signal(Pid, Signal) sends Signal to the process whose identifier is Pid. Signal can be specified directly as an integer or symbolically as an atom. Allowed atoms depend on the machine (e.g. 'SIGINT',

'SIGQUIT', 'SIGKILL', 'SIGUSR1', 'SIGUSR2', 'SIGALRM',. . . ). This predicate is an interface to the C

Unix function kill(2).

Errors

 Pid is a variable  instantiation error

Pid is neither a variable nor an integer type error(integer, Pid)

 Signal is a variable  instantiation error

Signal is neither a variable nor an integer or an atom

                                             type error(integer, Signal)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.27.28 sleep/1

Templates

     sleep(+number)

Description

sleep(Seconds) puts the GNU Prolog process to sleep for Seconds seconds. Seconds can be an integer or a floating point number (in which case it can be < 1). This predicate is an interface to the C Unix

function usleep(3).

Errors

 Seconds is a variable  instantiation error
 Seconds is neither a variable nor a number  type error(number, Seconds)
 Seconds is a number < 0  domain error(not less than zero, Seconds)

Portability

GNU Prolog predicate.

7.27.29 select/5

    select(+list, ?list, +list, ?list, +number)

Description

select(Reads, ReadyReads, Writes, ReadyWrites, TimeOut) waits for a number of streams (or file

descriptors) to change status. ReadyReads is unified with the list of elements listed in Reads that have characters available for reading. Similarly ReadyWrites is unified with the list of elements of Writes that are ok for immediate writing. The elements of Reads and Writes are either stream-terms or aliases or integers considered as file descriptors, e.g. for sockets (section 7.28, page 157). Streams that must be tested with select/5 should not be buffered. This can be done at the opening using open/4 (section 7.10.6, page 70) or later using set stream buffering/2 (section 7.10.27, page 82). TimeOut is an upper bound on the amount of time (in milliseconds) elapsed before select/5 returns. If TimeOut ≤ 0 (no timeout) select/5 waits until something is available (either or reading or for writing) and thus can block indefinitely. This predicate is an interface to the C Unix function select(2).

Errors

Reads (or Writes) is a partial list or a list with an element E which is a variable

                                           instantiation error

Reads is neither a partial list nor a list type error(list, Reads) Writes is neither a partial list nor a list type error(list, Writes)

ReadyReads is neither a partial list nor a list  type error(list, ReadyReads)
ReadyWrites is neither a partial list nor a list  type error(list, ReadyWrites)

an element E of the Reads (or Writes) list is neither a stream-term or alias nor an integer

                                           domain error(stream or alias, E)

an element E of the Reads (or Writes) list is not a selectable item

                                           domain error(selectable item, E)

an element E of the Reads (or Writes) list is an integer < 0

                                           domain error(not less than zero, E)

an element E of the Reads (or Writes) list is a stream-tern or alias not associated with an open stream

                                           existence error(stream, E)

an element E of the Reads list is associated to an output stream

                                           permission error(input, stream, E)

an element E of the Writes list is associated to an input stream

                                           permission error(output, stream, E)
TimeOut is a variable  instantiation error
TimeOut is neither a variable nor a number  type error(number, TimeOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28 Sockets input/output

7.28.1 Introduction

This set of predicates provides a way to manipulate sockets. The predicates are straightforward interfaces to the corresponding BSD-type socket functions. This facility is available if the sockets part of GNU Prolog has been installed. A reader familiar with BSD sockets will understand them immediately otherwise a study of sockets is needed.

The domain is either the atom 'AF INET' or 'AF UNIX' corresponding to the same domains in BSD-type sockets.

An address is either of the form 'AF INET'(HostName, Port) or 'AF UNIX'(SocketName). HostName is

an atom denoting a machine name, Port is a port number and SocketName is an atom denoting a socket.

By default, streams associated to sockets are block buffered. The predicate set stream buffering/2 (section 7.10.27, page 82) can be used to change this mode. They are also text streams by default. Use set stream type/2 (section 7.10.25, page 81) to change the type if binary streams are needed.

7.28.2 socket/2

Templates

     socket(+socket domain, -integer)

Description

socket(Domain, Socket) creates a socket whose domain is Domain (section 7.28, page 157) and unifies Socket with the descriptor identifying the socket. This predicate is an interface to the C Unix function

socket(2).

Errors

 Domain is a variable  instantiation error
 Domain is neither a variable nor an atom  type error(atom, Domain)
 Domain is an atom but not a valid socket domain domain error(socket domain, Domain)
 Socket is not a variable  type error(variable, Socket)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28.3 socket close/1

Templates

     socket close(+integer)

Description

socket close(Socket) closes the socket whose descriptor is Socket. This predicate should not be used if Socket has given rise to a stream, e.g. by socket connect/4 (section 7.28.5, page 159). In that case simply use close/2 (section 7.10.7, page 72) on the associated stream.

Errors

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28.4 socket bind/2

Templates

     socket bind(+integer, +socket address)

Description

socket bind(Socket, Address) binds the socket whose descriptor is Socket to the address specified by Address (section 7.28, page 157). If Address if of the form 'AF INET'(HostName, Port) and if HostName is uninstantiated then it is unified with the current machine name. If Port is uninstantiated, it is unified to a port number picked by the operating system. This predicate is an interface to the C Unix function bind(2).

Errors

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)
 Address is a variable  instantiation error
 Address is neither a variable nor a valid address domain error(socket address, Address)
 Address = 'AF UNIX'(E) and E is a variable  instantiation error
 Address = 'AF UNIX'(E) or 'AF INET'(E,  )

and E is neither a variable nor an atom

                                             type error(atom, E)
 Address = 'AF UNIX'(E) and E is an atom but

not a valid pathname

                                             domain error(os path, E)
 Address = 'AF INET'( , E) and E is neither a

variable nor an integer

                                             type error(integer, E)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28.5 socket connect/4

Templates

     socket connect(+integer, +socket address, -stream, -stream)

Description

socket connect(Socket, Address, StreamIn, StreamOut) connects the socket whose descriptor is

Socket to the address specified by Address (section 7.28, page 157). StreamIn is unified with a streamterm associated to the input of the connection (it is an input stream). Reading from this stream gets data from the socket. StreamOut is unified with a stream-term associated to the output of the connection (it is an output stream). Writing to this stream sends data to the socket. The use of select/5 can be useful (section 7.27.29, page 156). This predicate is an interface to the C Unix function connect(2).

Socket is a variable  instantiation error
Socket is neither a variable nor an integer  type error(integer, Socket)
Address is a variable  instantiation error
Address is neither a variable nor a valid address domain error(socket address, Address)
Address = 'AF UNIX'(E) or 'AF INET'(E,  ) or
Address = 'AF INET'( , E) and E is a variable

                                           instantiation error
Address = 'AF UNIX'(E) or 'AF INET'(E,  )

and E is neither a variable nor an atom

                                           type error(atom, E)
Address = 'AF UNIX'(E) and E is an atom but

not a valid pathname

                                           domain error(os path, E)
Address = 'AF INET'( , E) and E is neither a

variable nor an integer

                                           type error(integer, E)
StreamIn is not a variable  type error(variable, StreamIn)
StreamOut is not a variable  type error(variable, StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28.6 socket listen/2

Templates

    socket listen(+integer, +integer)

Description

socket listen(Socket, Length) defines the socket whose descriptor is Socket to have a maximum backlog queue of Length pending connections. This predicate is an interface to the C Unix function

listen(2).

Errors

Socket is a variable  instantiation error
Socket is neither a variable nor an integer  type error(integer, Socket)
Length is a variable  instantiation error
Length is neither a variable nor an integer  type error(integer, Length)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

7.28.7 socket accept/4, socket accept/3

Templates

    socket accept(+integer, -atom, -stream, -stream)
    socket accept(+integer, -stream, -stream)

Description

socket accept(Socket, Client, StreamIn, StreamOut) extracts the first connection to the socket

whose descriptor is Socket. If the domain is 'AF INET', Client is unified with an atom whose name is the Internet host address in numbers-and-dots notation of the connecting machine. StreamIn is unified with a stream-term associated to the input of the connection (it is an input stream). Reading from this stream gets data from the socket. StreamOut is unified with a stream-term associated to the output of the connection (it is an output stream). Writing to this stream sends data to the socket. The use of select/5 can be useful (section 7.27.29, page 156). This predicate is an interface to the C Unix function

accept(2).
socket accept(Socket, StreamIn, StreamOut) is equivalent to socket accept(Socket,  ,
StreamIn, StreamOut).

Errors

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)
 Client is not a variable  type error(variable, Client)
 StreamIn is not a variable  type error(variable, StreamIn)
 StreamOut is not a variable  type error(variable, StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 7.22.1, page 133)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

7.28.8 hostname address/2

Templates

     hostname address(+atom, ?atom)
     hostname address(?atom, +atom)

Description

hostname address(HostName, HostAddress) succeeds if the Internet host address in numbers-and-dots notation of HostName is HostAddress. Hostname can be given as a fully qualified name, or an unqualified name or an alias of the machine. The predicate will fail if the machine name or address cannot be resolved.

Errors

 HostName and HostAddress are variables  instantiation error
 HostName is neither a variable nor an atom  type error(atom, HostName)
 HostAddress is neither a variable nor an atom  type error(atom, HostAddress)
 Address is neither a variable nor a valid address domain error(socket address, Address)

Portability

7.29 Linedit management

The following predicates are only available if the linedit part of GNU Prolog has been installed.

7.29.1 get linedit prompt/1

Templates

     get linedit prompt(?atom)

Description get linedit prompt(Prompt) succeeds if Prompt is the current linedit prompt, e.g. the top-level prompt is '| ?-'. By default all other reads have an empty prompt. Errors Prompt is neither a variable nor an atom type error(atom, Pred) Portability GNU Prolog predicate.

7.29.2 set linedit prompt/1

Templates

     set linedit prompt(+atom)

Description set linedit prompt(Prompt) sets the current linedit prompt to Prompt. This prompt will be displayed for reads from a terminal (except for top-level reads). Errors

 Prompt is a variable  instantiation error

Prompt is neither a variable nor an atom type error(atom, Pred) Portability GNU Prolog predicate.

7.29.3 add linedit completion/1

Templates

     add linedit completion(+atom)

Description add linedit completion(Word) adds Word in the list of completion words maintained by linedit (section 3.2.5, page 18). Only words containing letters, digits and the underscore character are added (if Word does not respect this restriction the predicate fails).

Errors

 Word is a variable  instantiation error

Word is neither a variable nor an atom type error(atom, Word) Portability GNU Prolog predicate.

7.29.4 find linedit completion/2

Templates

     find linedit completion(+atom, ?atom)

Description

find linedit completion(Prefix, Word) succeeds if Word is a word beginning by Prefix and belongs

to the list of completion words maintained by linedit (section 3.2.5, page 18). This predicate is reexecutable on backtracking. Errors

 Prefix is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)

Word is neither a variable nor an atom type error(atom, Word) Portability GNU Prolog predicate.

7.30 Source reader facility

7.30.1 Introduction To be written...

164  7 PROLOG BUILT-IN PREDICATES

7.30.2 sr open/3
7.30.3 sr change options/2
7.30.4 sr close/1
7.30.5 sr read term/4
7.30.6 sr current descriptor/1
7.30.7 sr get stream/2
7.30.8 sr get module/3
7.30.9 sr get file name/2
7.30.10 sr get position/3
7.30.11 sr get include list/2
7.30.12 sr get include stream list/2
7.30.13 sr get size counters/3
7.30.14 sr get error counters/3
7.30.15 sr set error counters/3
7.30.16 sr error from exception/2
7.30.17 sr write message/8, sr write message/6, sr write message/4
7.30.18 sr write error/6, sr write error/4, sr write error/2

8 Finite domain solver and built-in predicates

8.1 Introduction

The finite domain (FD) constraint solver extends Prolog with constraints over FD. This facility is available if the FD part of GNU Prolog has been installed. The solver is an instance of the Constraint Logic Programming scheme introduced by Jaffar and Lassez in 1987 [6]. Constraints on FD are solved using propagation techniques, in particular arc-consistency (AC). The interested reader can refer to “Constraint Satisfaction in Logic Programming” of P. Van Hentenryck (1989) [7]. The solver is based on the clp(FD) solver [4]. The GNU Prolog FD solver offers arithmetic constraints, boolean constraints, reified constraints and symbolic constraints on an new kind of variables: Finite Domain variables.

8.1.1 Finite Domain variables

A new type of data is introduced: FD variables which can only take values in their domains. The initial domain of an FD variable is 0..fd max integer where fd max integer represents the greatest value that any FD variable can take. The predicate fd max integer/1 returns this value which may be different from the max integer Prolog flag (section 7.22.1, page 133). The domain of an FD variable X is reduced step by step by constraints in a monotonic way: when a value has been removed from the domain of X it will never reappear in the domain of X. An FD variable is fully compatible with both Prolog integers and Prolog variables. Namely, when an FD variable is expected by an FD constraint it is possible to pass a Prolog integer (considered as an FD variable whose domain is a singleton) or a Prolog variable (immediately bound to an initial range 0..fd max integer). This avoids the need for specific type declaration. Although it is not necessary to declare the initial domain of an FD variable (since it will be bound 0..fd max integer when appearing for the fist time in a constraint) it is advantageous to do so and thus reduce as soon as possible the size of its domain: particularly because GNU Prolog, for efficiency reasons, does not check for overflows. For instance, without any preliminary domain definitions for X, Y and Z, the non-linear constraint X*Y#=Z will fail due to an overflow when computing the upper bound of the domain of Z: fd max integer × fd max integer. This overflow causes a negative result for the upper bound and the constraint then fails.

There are two internal representations for an FD variable:

• interval representation: only the min and the max of the variable are maintained. In this representation it is possible to store values included in 0..fd max integer.

• sparse representation: an additional bit-vector is used to store the set of possible values for the variable (i.e. the domain). In this representation it is possible to store values included in 0..vector max. By default vector max is set to 127. This value can be redefined via an environment variable VECTORMAX or via the built-in predicate fd set vector max/1 (section 8.2.3, page 167). The predicate fd vector max/1 returns the current value of vector max (section 8.2.1, page 166).

The initial representation for an FD variable X is always an interval representation and is switched to a sparse representation when a “hole” appears in the domain (e.g. due to an inequality constraint). Once a variable uses a sparse representation it will not switch back to an interval representation even if there are no longer holes in its domain. When this switching occurs some values in the domain of X can be lost since vector max is less than fd max integer. We say that “X is extra-constrained” since X is constrained by the solver to the domain 0..vector max (via an imaginary constraint X #=< vector max ). An extra cstr is associated to each FD variable to indicate that values have been lost due to the switch to a sparse representation. This flag is updated on every operations. The domain of an extra-constrained FD variable is output followed by the @ symbol. When a constraint fails on a extra-constrained variable

a message Warning:  Vector too small - maybe lost solutions (FD Var:N) is displayed (N is the

address of the involved variable).

Example 1 (vector max = 127):

Constraint on X Domain of X    extra cstr Lost values

 X #=< 512      0..512            off  none
 X #\= 10       0..9:11..127      on      128..512
 X #=< 100      0..9:11..100      off  none

In this example, when the constraint X #\= 10 is posted some values are lost, the extra cstr is then switched on. However, posting the constraint X #=< 100 will turn off the flag (no values are lost).

Example 2:

Constraint on X Domain of X                     extra cstr Lost values

 X #=< 512      0..512                             off  none
 X #\= 10       0..9:11..127                       on      128..512
 X #>= 256      Warning:  Vector too small...      on      128..512

In this example, the constraint X #>= 256 fails due to the lost of 128..512 so a message is displayed onto the terminal. The solution would consist in increasing the size of the vector either by setting the environment variable VECTORMAX (e.g. to 512) or using fd set vector max(512).

Finally, bit-vectors are not dynamic, i.e. all vectors have the same size (0..vector max). So the use of fd set vector max/1 is limited to the initial definition of vector sizes and must occur before any constraint. As seen before, the solver tries to display a message when a failure occurs due to a too short vector max. Unfortunately, in some cases it cannot detect the lost of values and no message is emitted. So the user should always take care to this parameter to be sure that it is large to encode any vector.

8.2 FD variable parameters

8.2.1 fd max integer/1

Templates

     fd max integer(?integer)

Description

fd max integer(N) succeeds if N is the current value of fd max integer (section 8.1, page 165).

Errors

N is neither a variable nor an integer type error(integer, N)

Portability

GNU Prolog predicate.

8.2.2 fd vector max/1

Templates

     fd vector max(?integer)

fd vector max(N) succeeds if N is the current value of vector max (section 8.1, page 165). Errors N is neither a variable nor an integer type error(integer, N) Portability GNU Prolog predicate.

8.2.3 fd set vector max/1

Templates

     fd set vector max(+integer)

Description fd set vector max(N) initializes vector max based on the value N (section 8.1, page 165). More precisely, on 32 bit machines, vector max is set to the smallest value of (32*k)-1 which is ≥ N. Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)

Portability GNU Prolog predicate.

8.3 Initial value constraints

8.3.1 fd domain/3, fd domain bool/1

Templates

     fd domain(+fd variable list or fd variable, +integer, +integer)
     fd domain(?fd variable, +integer, +integer)
     fd domain bool(+fd variable list)
     fd domain bool(?fd variable)

Description

fd domain(Vars, Lower, Upper) constraints each element X of Vars to take a value in Lower..Upper.

This predicate is generally used to set the initial domain of variables to an interval. Vars can be also a single FD variable (or a single Prolog variable). fd domain bool(Vars) is equivalent to fd domain(Vars, 0, 1) and is used to declare boolean FD variables. Errors Vars is not a variable but is a partial list instantiation error Vars is neither a variable nor an FD variable nor an integer nor a list

                                           type error(list, Vars)

an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, E)
Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper)

Portability

GNU Prolog predicate.

8.3.2 fd domain/2

Templates

    fd domain(+fd variable list, +integer list)
    fd domain(?fd variable, +integer list)

Description

fd domain(Vars, Values) constraints each element X of the list Vars to take a value in the list Values. This predicate is generally used to set the initial domain of variables to a set of values. The domain of each variable of Vars uses a sparse representation. Vars can be also a single FD variable (or a single Prolog variable).

Errors

Vars is not a variable but is a partial list instantiation error Vars is neither a variable nor an FD variable nor an integer nor a list

                                           type error(list, Vars)

an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, E)

Values is a partial list or a list with an element E which is a variable

                                           instantiation error

Values is neither a partial list nor a list type error(list, Values) an element E of the Values list is neither a variable nor an integer

                                           type error(integer, E)

Portability

GNU Prolog predicate.

8.4 Type testing

8.4.1 fd var/1, non fd var/1, generic var/1, non generic var/1

     fd var(?term)
     non fd var(?term)

                                                 generic var(?term)
                                                 non generic var(?term)

Description fd var(Term) succeeds if Term is currently an FD variable. non fd var(Term) succeeds if Term is currently not an FD variable (opposite of fd var/1). generic var(Term) succeeds if Term is either a Prolog variable or an FD variable. non generic var(Term) succeeds if Term is neither a Prolog variable nor an FD variable (opposite of

generic var/1).

Errors None. Portability GNU Prolog predicate.

8.5 FD variable information

These predicate allow the user to get some information about FD variables. They are not constraints, they only return the current state of a variable.

8.5.1 fd min/2, fd max/2, fd size/2, fd dom/2

Templates

     fd min(+fd variable, ?integer)
     fd max(+fd variable, ?integer)
     fd size(+fd variable, ?integer)
     fd dom(+fd variable, ?integer list)

Description fd min(X, N) succeeds if N is the minimal value of the current domain of X. fd max(X, N) succeeds if N is the maximal value of the current domain of X. fd size(X, N) succeeds if N is the number of elements of the current domain of X. fd dom(X, Values) succeeds if Values is the list of values of the current domain of X. Errors

 X is a variable  instantiation error

X is neither an FD variable nor an integer type error(fd variable, X) N is neither a variable nor an integer type error(integer, N) an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, E)

Values is neither a partial list nor a list type error(list, Values)

Portability GNU Prolog predicate.

8.5.2 fd has extra cstr/1, fd has vector/1, fd use vector/1

Templates

     fd has extra cstr(+fd variable)
     fd has vector(+fd variable)
     fd use vector(+fd variable)

Description fd has extra cstr(X) succeeds if the extra cstr of X is currently on (section 8.1, page 165). fd has vector(X) succeeds if the current domain of X uses a sparse representation (section 8.1, page 165). fd use vector(X) enforces a sparse representation for the domain of X (section 8.1, page 165). Errors

 X is a variable  instantiation error

X is neither an FD variable nor an integer type error(fd variable, X) Portability GNU Prolog predicates.

8.6 Arithmetic constraints

8.6.1 FD arithmetic expressions An FD arithmetic expression is a Prolog term built from integers, variables (Prolog or FD variables), and functors (or operators) that represent arithmetic functions. The following table details the components of an FD arithmetic expression:

FD Expression  Result

Prolog variable domain 0..fd max integer

FD variable X        domain of X
integer number N     domain N..N
+ E                same as E
- E                opposite of E

E1 + E2  sum of E1 and E2

E1 - E2 subtraction of E2 from E1 E1 * E2 multiplication of E1 by E2 E1 / E2 integer division of E1 by E2 (only succeeds if the remainder is 0) E1 ** E2 E1 raised to the power of E2 (E1 or E2 must be an integer)

min(E1,E2)  minimum of E1 and E2
max(E1,E2)  maximum of E1 and E2
dist(E1,E2)  distance, i.e. |E1 - E2|

E1 // E2 quotient of the integer division of E1 by E2 E1 rem E2 remainder of the integer division of E1 by E2 quot rem(E1,E2,R) quotient of the integer division of E1 by E2

(R is the remainder of the integer division of E1 by E2)

FD expressions are not restricted to be linear. However non-linear constraints usually yield less constraint propagation than linear constraints.

+, -, *, /, //, rem and ** are predefined infix operators. + and - are predefined prefix operators (section 7.14.10, page 100).

Errors

a sub-expression is of the form ** E and E is a variable

                                           instantiation error

a sub-expression E is neither a variable nor an integer nor an FD arithmetic functor

                                           type error(fd evaluable, E)
an expression is too complex  resource error(too big fd constraint)

8.6.2 Partial AC: (#=)/2 - constraint equal, (#\=)/2 - constraint not equal, (#<)/2 - constraint less than, (#=<)/2 - constraint less than or equal, (#>)/2 - constraint greater than, (#>=)/2 - constraint greater than or equal

Templates

    #=(?fd evaluable, ?fd evaluable)
    #\=(?fd evaluable, ?fd evaluable)
    #<(?fd evaluable, ?fd evaluable)
    #=<(?fd evaluable, ?fd evaluable)
    #>(?fd evaluable, ?fd evaluable)
    #>=(?fd evaluable, ?fd evaluable)

Description

FdExpr1 #= FdExpr2 constrains FdExpr1 to be equal to FdExpr2.
FdExpr1 #\= FdExpr2 constrains FdExpr1 to be different from FdExpr2.
FdExpr1 #< FdExpr2 constrains FdExpr1 to be less than FdExpr2.

FdExpr1 #=< FdExpr2 constrains FdExpr1 to be less than or equal to FdExpr2.

FdExpr1 #> FdExpr2 constrains FdExpr1 to be greater than FdExpr2.

FdExpr1 #>= FdExpr2 constrains FdExpr1 to be greater than or equal to FdExpr2. FdExpr1 and FdExpr2 are arithmetic FD expressions (section 8.6.1, page 170). #=, #\=, #<, #=<, #> and #>= are predefined infix operators (section 7.14.10, page 100). These predicates post arithmetic constraints that are managed by the solver using a partial arc-consistency algorithm to reduce the domain of involved variables. In this scheme only the bounds of the domain of variables are updated. This leads to less propagation than full arc-consistency techniques (section 8.6.3, page 172) but is generally more efficient for arithmetic. These arithmetic constraints can be reified (section 8.7, page 173). Errors Refer to the syntax of arithmetic FD expressions for possible errors (section 8.6.1, page 170). Portability GNU Prolog predicates.

8.6.3 Full AC: (#=#)/2 - constraint equal, (#\=#)/2 - constraint not equal, (#<#)/2 - constraint less than, (#=<#)/2 - constraint less than or equal, (#>#)/2 - constraint greater than, (#>=#)/2 - constraint greater than or equal

Templates

     #=#(?fd evaluable, ?fd evaluable)
     #\=#(?fd evaluable, ?fd evaluable)
     #<#(?fd evaluable, ?fd evaluable)
     #=<#(?fd evaluable, ?fd evaluable)
     #>#(?fd evaluable, ?fd evaluable)
     #>=#(?fd evaluable, ?fd evaluable)

Description

FdExpr1 #=# FdExpr2 constrains FdExpr1 to be equal to FdExpr2.
FdExpr1 #\=# FdExpr2 constrains FdExpr1 to be different from FdExpr2.
FdExpr1 #<# FdExpr2 constrains FdExpr1 to be less than FdExpr2.
FdExpr1 #=<# FdExpr2 constrains FdExpr1 to be less than or equal to FdExpr2.
FdExpr1 #># FdExpr2 constrains FdExpr1 to be greater than FdExpr2.

FdExpr1 #>=# FdExpr2 constrains FdExpr1 to be greater than or equal to FdExpr2. FdExpr1 and FdExpr2 are arithmetic FD expressions (section 8.6.1, page 170). #=#, #\=#, #<#, #=<#, #># and #>=# are predefined infix operators (section 7.14.10, page 100). These predicates post arithmetic constraints that are managed by the solver using a full arc-consistency algorithm to reduce the domain of involved variables. In this scheme the full domain of variables is updated. This leads to more propagation than partial arc-consistency techniques (section 8.6.1, page 170) but is generally less efficient for arithmetic. These arithmetic constraints can be reified (section 8.7.1, page 173).

Errors

Refer to the syntax of arithmetic FD expressions for possible errors (section 8.6.1, page 170).

Portability

GNU Prolog predicates.

8.6.4 fd prime/1, fd not prime/1

Templates

     fd prime(?fd variable)
     fd not prime(?fd variable)

Description

fd prime(X) constraints X to be a prime number between 0..vector max. This constraint enforces a sparse representation for the domain of X (section 8.1, page 165).

fd not prime(X) constraints X to be a non prime number between 0..vector max. This constraint enforces a sparse representation for the domain of X (section 8.1, page 165).

Errors

X is neither an FD variable nor an integer type error(fd variable, X)

Portability

GNU Prolog predicates.

8.7 Boolean and reified constraints

8.7.1 Boolean FD expressions

An boolean FD expression is a Prolog term built from integers (0 for false, 1 for true), variables (Prolog or FD variables), partial AC arithmetic constraints (section 8.6.2, page 171), full AC arithmetic constraints (section 8.6.3, page 172) and functors (or operators) that represent boolean functions. When a subexpression of a boolean expression is an arithmetic constraint c , it is reified. Namely, as soon as the solver detects that c is true (i.e. entailed) the sub-expression has the value 1. Similarly when the solver detects that c is false (i.e. disentailed) the sub-expression evaluates as 0. While neither the entailment nor the disentailment can be detected the sub-expression is evaluated as a domain 0..1. The following table details the components of an FD boolean expression:

FD Expression Result
Prolog variable domain 0..1
FD variable X domain of X, X is constrained to be in 0..1
0 (integer)  0 (false)
1 (integer)  1 (true)

#\ E  not E
E1 #<=> E2     E1 equivalent to E2

E1 #\<=> E2    E1 not equivalent to E2 (i.e. E1 different from E2)
E1 ## E2       E1 exclusive OR E2 (i.e. E1 not equivalent to E2)

E1 #==> E2     E1 implies E2
E1 #\==> E2    E1 does not imply E2
E1 #/\ E2      E1 AND E2
E1 #\/\ E2     E1 NAND E2
E1 #\/ E2      E1 OR E2
E1 #\\/ E2     E1 NOR E2

#<=>, #\<=>, ##, #==>, #\==>, #/\, #\/\, #\/ and #\\/ are predefined infix operators. #\ is a predefined prefix operator (section 7.14.10, page 100).

Errors

a sub-expression E is neither a variable nor an integer (0 or 1) nor an FD boolean functor nor reified constraint

                                           type error(fd bool evaluable, E)

an expression is too complex  resource error(too big fd constraint)

a sub-expression is an invalid reified constraint an arithmetic constraint error (section 8.6.1, page 170)

8.7.2 (#\)/1 - constraint NOT, (#<=>)/2 - constraint equivalent, (#\<=>)/2 - constraint different, (##)/2 - constraint XOR, (#==>)/2 - constraint imply, (#\==>)/2 - constraint not imply, (#/\)/2 - constraint AND, (#\/\)/2 - constraint NAND, (#\/)/2 - constraint OR, (#\\/)/2 - constraint NOR

Templates

    #\(?fd bool evaluable)
    #<=>(?fd bool evaluable, ?fd bool evaluable)
    #\<=>(?fd bool evaluable, ?fd bool evaluable)
    ##(?fd bool evaluable, ?fd bool evaluable)
    #==>(?fd bool evaluable, ?fd bool evaluable)
    #\==>(?fd bool evaluable, ?fd bool evaluable)
    #/\(?fd bool evaluable, ?fd bool evaluable)
    #\/\(?fd bool evaluable, ?fd bool evaluable)
    #\/(?fd bool evaluable, ?fd bool evaluable)
    #\\/(?fd bool evaluable, ?fd bool evaluable)

Description

#\= FdBoolExpr1 constraints FdBoolExpr1 to be false.
FdBoolExpr1 #<=> FdBoolExpr2 constrains FdBoolExpr1 to be equivalent to FdBoolExpr2.
FdBoolExpr1 #\<=> FdBoolExpr2 constrains FdBoolExpr1 to be equivalent to not FdBoolExpr2.
FdBoolExpr1 ## FdBoolExpr2 constrains FdBoolExpr1 XOR FdBoolExpr2 to be true

FdBoolExpr1 #==> FdBoolExpr2 constrains FdBoolExpr1 to imply FdBoolExpr2.
FdBoolExpr1 #\==> FdBoolExpr2 constrains FdBoolExpr1 to not imply FdBoolExpr2.
FdBoolExpr1 #/\ FdBoolExpr2 constrains FdBoolExpr1 AND FdBoolExpr2 to be true.
FdBoolExpr1 #\/\ FdBoolExpr2 constrains FdBoolExpr1 AND FdBoolExpr2 to be false.
FdBoolExpr1 #\/ FdBoolExpr2 constrains FdBoolExpr1 OR FdBoolExpr2 to be true.
FdBoolExpr1 #\\/ FdBoolExpr2 constrains FdBoolExpr1 OR FdBoolExpr2 to be false.

FdBoolExpr1 and FdBoolExpr2 are boolean FD expressions (section 8.7.1, page 173). Note that #\<=> (not equivalent) and ## (exclusive or) are synonymous. These predicates post boolean constraints that are managed by the FD solver using a partial arcconsistency algorithm to reduce the domain of involved variables. The (dis)entailment of reified constraints is detected using either the bounds (for partial AC arithmetic constraints) or the full domain (for full AC arithmetic constraints). #<=>, #\<=>, ##, #==>, #\==>, #/\, #\/\, #\/ and #\\/ are predefined infix operators. #\ is a predefined prefix operator (section 7.14.10, page 100). Errors Refer to the syntax of boolean FD expressions for possible errors (section 8.7.1, page 173). Portability GNU Prolog predicates.

8.7.3 fd cardinality/2, fd cardinality/3, fd at least one/1, fd at most one/1,
       fd only one/1

Templates

     fd cardinality(+fd bool evaluable list, ?fd variable)
     fd cardinality(+integer, ?fd variable, +integer)
     fd at least one(+fd bool evaluable list)
     fd at most one(+fd bool evaluable list)
     fd only one(+fd bool evaluable list)

Description fd cardinality(List, Count) unifies Count with the number of constraints that are true in List. This is equivalent to post the constraint B_{1} + B_{2} + ...+ B_{n} #= Count where each variable Bi is a new variable defined by the constraint B_{i} #<=> C_{i} where C_{i} is the ith constraint of List. Each C_{i} must be a boolean FD expression (section 8.7.1, page 173).

fd cardinality(Lower, List, Upper) is equivalent to fd cardinality(List, Count), Lower #=<
Count, Count #=< Upper
fd at least one(List) is equivalent to fd cardinality(List, Count), Count #>= 1.
fd at most one(List) is equivalent to fd cardinality(List, Count), Count #=< 1.

fd only one(List) is equivalent to fd cardinality(List, 1).

Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) Count is neither an FD variable nor an integer type error(fd variable, Count)

 Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

 Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper) an element E of the List list is an invalid boolean expression an FD boolean constraint (section 8.7.1, page 173)

Portability

GNU Prolog predicates.

8.8 Symbolic constraints

8.8.1 fd all different/1

Templates

     fd all different(+fd variable list)

Description

fd all different(List) constrains all variables in List to take distinct values. This is equivalent to posting an inequality constraint for each pair of variables. This constraint is triggered when a variable becomes ground, removing its value from the domain of the other variables.

Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Portability

GNU Prolog predicate.

8.8.2 fd element/3

Templates

     fd element(?fd variable, +integer list, ?fd variable)

Description

fd element(I, List, X) constraints X to be equal to the Ith integer (from 1) of List.

I is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, I)

X is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, X)

List is a partial list or a list with an element E which is a variable

                                             instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer

                                             type error(integer, E)

Portability

GNU Prolog predicate.

8.8.3 fd element var/3

Templates

     fd element var(?fd variable, +fd variable list, ?fd variable)

Description

fd element var(I, List, X) constraints X to be equal to the Ith variable (from 1) of List. This constraint is similar to fd element/3 (section 8.8.2, page 176) but List can also contain FD variables (rather than just integers).

Errors

I is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, I)

X is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, X)
 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Portability

GNU Prolog predicate.

8.8.4 fd atmost/3, fd atleast/3, fd exactly/3

Templates

     fd atmost(+integer, +fd variable list, +integer)
     fd atleast(+integer, +fd variable list, +integer)
     fd exactly(+integer, +fd variable list, +integer)

Description

fd atmost(N, List, V) posts the constraint that at most N variables of List are equal to the value V.

fd atleast(N, List, V) posts the constraint that at least N variables of List are equal to the value V. fd exactly(N, List, V) posts the constraint that at exactly N variables of List are equal to the value

V.

These constraints are special cases of fd cardinality/2 (section 8.7.3, page 175) but their implementation is more efficient.

Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 V is a variable  instantiation error

V is neither a variable nor an integer type error(integer, V)

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, E)

Portability

GNU Prolog predicates.

8.8.5 fd relation/2, fd relationc/2

Templates

     fd relation(+integer list list, ?fd variable list)
     fd relationc(+integer list list, ?fd variable list)

Description

fd relation(Relation, Vars) constraints the tuple of variables Vars to be equal to one tuple of the list Relation. A tuple is represented by a list.

Example: definition of the boolean AND relation so that X AND Y ⇔ Z:

     and(X,Y,Z):-
             fd_relation([[0,0,0],[0,1,0],[1,0,0],[1,1,1]], [X,Y,Z]).

fd relationc(Columns, Vars) is similar to fd relation/2 except that the relation is not given as the list of tuples but as the list of the columns of the relation. A column is represented by a list.

Example:

     and(X,Y,Z):-
             fd_relationc([[0,0,1,1],[0,1,0,1],[0,0,0,1]], [X,Y,Z]).

Errors

Relation is a partial list or a list with a sub-term E which is a variable

                                             instantiation error
 Relation is neither a partial list nor a list  type error(list, Relation)

an element E of the Relation list is neither a variable nor an integer

                                             type error(integer, E)
 Vars is a partial list  instantiation error

Vars is neither a partial list nor a list type error(list, Vars) an element E of the Vars list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Portability

GNU Prolog predicates.

8.9 Labeling constraints

8.9.1 fd labeling/2, fd labeling/1, fd labelingff/1

Templates

fd labeling(+fd variable list, +fd labeling option list)
fd labeling(+fd variable, +fd labeling option list)
fd labeling(+fd variable list)
fd labeling(+fd variable)
fd labelingff(+fd variable list)
fd labelingff(+fd variable)

Description

fd labeling(Vars, Options) assigns a value to each variable X of the list Vars according to the list of labeling options given by Options. Vars can be also a single FD variable. This predicate is re-executable on backtracking.

FD labeling options: Options is a list of labeling options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• variable method(V): specifies the heuristics to select the variable to enumerate:

– standard: no heuristics, the leftmost variable is selected.

– first fail (or ff): selects the variable with the smallest number of elements in its domain. If several variables have the same number of elements the leftmost variable is selected.

– most constrained: like first fail but when several variables have the same number of elements selects the variable that appears in most constraints.

– smallest: selects the variable that has the smallest value in its domain. If there is more than one such variable selects the variable that appears in most constraints.

– largest: selects the variable that has the greatest value in its domain. If there is more than one such variable selects the variable that appears in most constraints.

– max regret: selects the variable that has the greatest difference between the smallest value and the next value of its domain. If there is more than one such variable selects the variable that appears in most constraints.

– random: selects randomly a variable. Each variable is only chosen once. The default value is standard.

• reorder(true/false): specifies if the variable heuristics should dynamically reorder the list of variable (true) or not (false). Dynamic reordering is generally more efficient but in some cases a static ordering is faster. The default value is true.

• value method(V): specifies the heuristics to select the value to assign to the chosen variable: – min: enumerates the values from the smallest to the greatest (default).

– max: enumerates the values from the greatest to the smallest.

– middle: enumerates the values from the middle to the bounds.

– bounds: enumerates the values from the bounds to the middle.

– random: enumerates the values randomly. Each value is only tried once. The default value is min.

• backtracks(B): unifies B with the number of backtracks during the enumeration.

fd labeling(Vars) is equivalent to fd labeling(Vars, []).
fd labelingff(Vars) is equivalent to fd labeling(Vars, [variable method(ff)]).

Errors

Vars is a partial list or a list with an element E which is a variable

                                             instantiation error

Vars is neither a partial list nor a list type error(list, Vars) an element E of the Vars list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a labeling option

                                             domain error(fd labeling option, E)

Portability

GNU Prolog predicates.

8.10 Optimization constraints

8.10.1 fd minimize/2, fd maximize/2

Templates

     fd minimize(+callable term, ?fd variable)
     fd maximize(+callable term, ?fd variable)

Description

fd minimize(Goal, X) repeatedly calls Goal to find a value that minimizes the variable X. Goal is a Prolog goal that should instantiate X, a common case being the use of fd labeling/2 (section 8.9.1, page 179). This predicate uses a branch-and-bound algorithm with restart: each time call(Goal) succeeds the computation restarts with a new constraint X #< V where V is the value of X at the end of the last call of Goal. When a failure occurs (either because there are no remaining choice-points for Goal or because the added constraint is inconsistent with the rest of the store) the last solution is recomputed since it is optimal.

fd maximize(Goal, X) is similar to fd minimize/2 but X is maximized.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 7.22.1, page 133)

                                             existence error(procedure, Pred)

X is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, X)

182  8 FINITE DOMAIN SOLVER AND BUILT-IN PREDICATES

9 Interfacing Prolog and C

9.1 Calling C from Prolog

9.1.1 Introduction

This interface allows a Prolog predicate to call a C function. Here are some features of this facility:

• implicit Prolog ↔ C data conversions for simple types.

• functions to handle complex types.

• error detection depending on the type of the argument.

• different kinds of arguments: input, output or input/output.

• possibility to write non-deterministic code.

This interface can then be used to write both simple and complex C routines. A simple routine uses either input or output arguments which type is simple. In that case the user does not need any knowledge of Prolog data structures since all Prolog ↔ C data conversions are implicitly achieved. To manipulate complex terms (lists, structures) a set of functions is provided. Finally it is also possible to write nondeterministic C code.

9.1.2 foreign/2 directive

foreign/2 directive (section 6.1.14, page 45) declares a C function interface. The general form is foreign(Template, Options) which defines an interface predicate whose prototype is Template according to the options given by Options. Template is a callable term specifying the type/mode of each argument of the associated Prolog predicate.

Foreign options: Options is a list of foreign options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• fct name(F): F is an atom representing the name of the C function to call. By default the name of the C function is the same as the principal functor of Template. In any case, the atom associated to the name of the function must conforms to the syntax of C identifiers.

• return(boolean/none/jump): specifies the value returned by the C function:

– boolean: the type of the function is Bool (returns TRUE on success, FALSE otherwise).

– none: the type of the function is void (no returned value).

– jump: the type of the function is void(*)() (returns the address of a Prolog code to execute). The default value is boolean.

• bip name(Name, Arity): initializes the error context with Name and Arity. If an error occurs this information is used to indicate from which predicate the error occurred (section 5.3.1, page 37). It is also possible to prevent the initialization of the error context using bip name(none). By default Name and Arity are set to the functor and arity of Template.

• choice size(N): this option specifies that the function implements a non-deterministic code. N is an integer specifying the size needed by the non-deterministic C function. This facility is explained later (section 9.1.7, page 186). By default a foreign function is deterministic.

foreign(Template) is equivalent to foreign(Template, []).

Foreign modes and types: each argument of Template specifies the foreign mode and type of the corresponding argument. This information is used to check the type of effective arguments at run-time and to perform Prolog ↔ C data conversions. Each argument of Template is formed with a mode symbol followed by a type name. Possible foreign modes are:

• +: input argument.

• -: output argument.

• ?: input/output argument.

Possible foreign types are:

Foreign type Prolog type  C type Description of the C type
integer  integer  long  value of the integer
positive  positive integer  long  value of the integer
float  floating point number double value of the floating point number
number  number  double value of the number
atom  atom  int  internal key of the atom
boolean  boolean  int  value of the boolean (0=false, 1=true)
char  character  int  value of (the code of) the character
code  character code  int  value of the character-code
byte  byte  int  value of the byte
in char  in-character  int  value of the character or -1 for end-of-file
in code  in-character code  int  value of the character-code or -1 for end-of-file
in byte  in-byte  int  value of the byte or -1 for the end-of-file
string  atom  char * C string containing the name of the atom
chars  character list  char * C string containing the characters of the list
codes  character-code list  char * C string containing the characters of the list
term  Prolog term  PlTerm generic Prolog term

Simple foreign type: a simple type is any foreign type listed in the above tabled except term. A simple foreign type is an atomic term (character and character-code lists are in fact lists of constants). Each simple foreign type is converted to/from a C type to simplify the writing of the C function.

Complex foreign type: type foreign type term refers to any Prolog term (e.g. lists, structures…). When such an type is specified the argument is passed to the C function as a PlTerm (GNU Prolog C type equivalent to a long). Several functions are provided to manipulate PlTerm variables (section 9.2, page 190). Since the original term is passed to the function it is possible to read its value or to unify it. So the meaning of the mode symbol is less significant. For this reason it is possible to omit the mode symbol. In that case term is equivalent to +term.

9.1.3 The C function

The C code is written in a C file which must first include the GNU Prolog header file called gprolog.h. This file contains all GNU Prolog C definitions (constants, types, prototypes,…).

The type returned by a C function depends on the value of the return foreign option (section 9.1.2, page 183). If it is boolean then the C function is of type Bool and shall return TRUE in case of success and FALSE otherwise. If the return option is none the C function is of type void. Finally if it is jump, the function shall return the address of a Prolog predicate and, at the exit of the function, the control is given to that predicate.

The type of the arguments of the C function depends on the mode and type declaration specified in Template for the corresponding argument as explained in the following sections.

9.1.4 Input arguments

An input argument is tested at run-time to check if its type conforms to the foreign type and then it is passed to the C function. The type of the associated C argument is given by the above table (section 9.1.2, page 183). For instance, the effective argument Arg associated to +positive foreign declaration is submitted to the following process:

• if Arg is a variable an instantiation error is raised.

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise the value of Arg is passed to the C is passed to the C function as an integer (long).

When +string is specified the string passed to the function is the internal string of the corresponding atom and should not be modified.

When +term is specified the term passed to the function is the original Prolog term. It can be read and/or unified. It is also the case when term is specified without any mode symbol.

9.1.5 Output arguments

An output argument is tested at run-time to check if its type conforms to the foreign type and it is unified with the value set by the C function. The type of the associated C argument is a pointer to the type given by the above table (section 9.1.2, page 183). For instance, the effective argument Arg associated to -positive foreign declaration is handled as follows:

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise a pointer to an integer (long *) is passed to the C function. If the function returns TRUE the integer stored at this location is unified with Arg.

When -term is specified, the function must construct a term into the its corresponding argument (which is of type PlTerm *). At the exit of the function this term will be unified with the actual predicate argument.

9.1.6 Input/output arguments

Basically an input/output argument is treated as in input argument if it is not a variable, as an output argument otherwise. The type of the associated C argument is a pointer to a FIOArg (GNU Prolog C type) defined as follows:

     typedef struct
         {
          Bool is_var;
          Bool unify;
          union
             {
              long   l;
              char  *s;
              double d;
             }value;
         }FIOArg;

The field is var is set to TRUE if the argument is a variable and FALSE otherwise. This value can be tested by the C function to determine which treatment to perform. The field unify controls whether the effective argument must be unified at the exit of the C function. Initially unify is set to the same value as is var (i.e. a variable argument will be unified while a non-variable argument will not) but it can be modified by the C function. The field value stores the value of the argument. It is declared as a C union since there are several kinds of value types. The field s is used for C strings, d for C doubles and l otherwise (int, long, PlTerm). if is var is FALSE then value contains the input value of the argument with the same conventions as for input arguments (section 9.1.4, page 185). At the exit of the function, if unify is TRUE value must contain the value to unify with the same conventions as for output arguments (section 9.1.5, page 185).

For instance, the effective argument Arg associated to ?positive foreign declaration is handled as follows:

• if Arg is a variable is var and unify are set to TRUE else to FALSE and its value is copied in

     value.l.

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise a pointer to the FIOArg (FIOArg *) is passed to the C function. If the function returns TRUE and if unify is TRUE the value stored in value.l is unified with Arg.

9.1.7 Writing non-deterministic C code

The interface allows the user to write non-deterministic C code. When a C function is non-deterministic, a choice-point is created for this function. When a failure occurs, if all more recent non-deterministic code are finished, the function is re-invoked. It is then important to inform Prolog when there is no more solution (i.e. no more choice) for a non-deterministic code. So, when no more choices remains the function must remove the choice-point. The interface increments a counter each time the function is re-invoked. At the first call this counter is equal to 0. This information allows the function to detect its first call. When writing non-deterministic code, it is often useful to record data between consecutive re-invocations of the function. The interface maintains a buffer to record such an information. The size of this buffer is given by choice size(N) when using foreign/2 (section 9.1.2, page 183). This size is the number of (consecutive) longs needed by the C function. Inside the function it is possible to call the following functions/macros:

     void Get_Choice_Counter(void)
     TYPE Get_Choice_Buffer (TYPE)
     void No_More_Choice    (void)

The function Get Choice Counter() returns the value of the invocation counter (0 at the first call).

The macro Get Choice Buffer(TYPE) returns a pointer to the buffer (casted to TYPE ).

The function No More Choice() deletes the choice point associated to the function.

9.1.8 Example: input and output arguments

All examples presented here can be found in the ExamplesC sub-directory of the distribution, in the files examp.pl (Prolog part) and examp c.c (C part).

Let us define a predicate first occurrence(A, C, P) which unifies P with the position (from 0) of the first occurrence of the character C in the atom A. The predicate must fail if C does not appear in A.

In the prolog file examp.pl:

:- foreign(first occurrence(+string, +char, -positive)).

In the C file examp c.c:

#include <string.h>
#include "gprolog.h"
Bool
first_occurrence(char *str, long c, long *pos)
{
  char *p;
  p = strchr(str, c);
  if (p == NULL)                /* C does not appear in A */
    return FALSE;               /* fail */
  *pos = p - str;               /* set the output argument */
  return TRUE;                  /* succeed */
}

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Examples of use:

| ?- first_occurrence(prolog, p, X).
X = 0
| ?- first_occurrence(prolog, k, X).
no
| ?- first_occurrence(prolog, A, X).
{exception: error(instantiation_error,first_occurrence/3)}
| ?- first_occurrence(prolog, 1 ,X).
{exception: error(type_error(character,1),first_occurrence/3)}

9.1.9 Example: non-deterministic code

We here define a predicate occurrence(A, C, P) which unifies P with the position (from 0) of one occurrence of the character C in the atom A. The predicate will fail if C does not appear in A. The predicate is re-executable on backtracking. The information that must be recorded between two invocations of the function is the next starting position in A to search for C.

In the prolog file examp.pl:

:- foreign(occurrence(+string, +char, -positive), [choice size(1)]).

In the C file examp c.c:

#include <string.h>
#include "gprolog.h"
Bool
occurrence(char *str, long c, long *pos)
{

  char **info_pos;
  char *p;
  info_pos = Get_Choice_Buffer(char **); /* recover the buffer */
  if (Get_Choice_Counter() == 0)        /* first invocation ? */
    *info_pos = str;
  p = strchr(*info_pos, c);
  if (p == NULL)                /* C does not appear */
    {
      No_More_Choice();         /* remove choice-point */
      return FALSE;             /* fail */
    }
  *pos = p - str;               /* set the output argument */
  *info_pos = p + 1;            /* update next starting pos */
  return TRUE;                  /* succeed */
}

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Examples of use:

| ?- occurrence(prolog, o, X).

X = 2 ?  (here the user presses ; to compute another solution)
X = 4 ?  (here the user presses ; to compute another solution)
no  (no more solution)

| ?- occurrence(prolog, k, X).
no

In the first example when the second (the last) occurrence is found (X=4) the choice-point remains and the failure is detected only when another solution is requested (by pressing ;). It is possible to improve this behavior by deleting the choice-point when there is no more occurrence. To do this it is necessary to do one search ahead. The information stored is the position of the next occurrence. Let us define such a behavior for the predicate occurrence2/3.

In the prolog file examp.pl:

:- foreign(occurrence2(+string, +char, -positive), [choice size(1)]).

In the C file examp c.c:

#include <string.h>
#include "gprolog.h"
Bool
occurrence2(char *str, long c, long *pos)
{
  char **info_pos;
  char *p;
  info_pos = Get_Choice_Buffer(char **); /* recover the buffer */

  if (Get_Choice_Counter() == 0) /* first invocation ? */
    {
      p = strchr(str, c);
      if (p == NULL)            /* C does not appear at all */
        {
          No_More_Choice();     /* remove choice-point */
          return FALSE;         /* fail */
        }
      *info_pos = p;
    }
                                /* info_pos = an occurrence */
  *pos = *info_pos - str;       /* set the output argument */
  p = strchr(*info_pos + 1, c);
  if (p == NULL)                /* no more occurrence */
    No_More_Choice();           /* remove choice-point */
  else
    *info_pos = p;              /* else update next solution */
  return TRUE;                  /* succeed */
}

Examples of use:

| ?- occurrence2(prolog, l, X).

X = 3  (here the user is not prompted since there is no more alternative)

| ?- occurrence2(prolog, o, X).

X = 2 ?  (here the user presses ; to compute another solution)
X = 4  (here the user is not prompted since there is no more alternative)

9.1.10 Example: input/output arguments

We here define a predicate char ascii(Char, Code) which converts in both directions the character Char and its character-code Code. This predicate is then similar to char code/2 (section 7.19.4, page 116).

In the prolog file examp.pl:

:- foreign(char ascii(?char, ?code), [fct name('Char Ascii')]).

In the C file examp c.c:

#include "gprolog.h"
Bool
char_ascii(FIOArg *c, FIOArg *ascii)
{
  if (!c->is_var)               /* Char is not a variable */
    {
      ascii->unify = TRUE;      /* enforce unif. of Code */
      ascii->value.l = c->value.l; /* set Code */
      return TRUE;              /* succeed */

         }
       if (ascii->is_var)            /* Code is also a variable */
         Pl_Err_Instantiation();     /* emit instantiation_error */
       c->value.l = ascii->value.l;  /* set Char */
       return TRUE;                  /* succeed */
     }

If Char is instantiated it is necessary to enforce the unification of Code since it could be instantiated. Recall that by default if an input/output argument is instantiated it will not be unified at the exit of the function (section 9.1.6, page 185). If both Char and Code are variables the function raises an instantiation error. The way to raise Prolog errors is described later (section 9.3, page 196).

The compilation produces an executable called examp:

     % gplc examp.pl examp c.c

Examples of use:

     | ?- char_ascii(a, X).
     X = 97
     | ?- char_ascii(X, 65).
     X = 'A'
     | ?- char_ascii(a, 12).
     no
     | ?- char_ascii(X, X).
     {exception: error(instantiation_error,char_ascii/2)}
     | ?- char_ascii(1, 12).
     {exception: error(type_error(character,1),char_ascii/2)}

9.2 Manipulating Prolog terms

9.2.1 Introduction

In the following we presents a set of functions to manipulate Prolog terms. For simple foreign terms the functions manipulate simple C types (section 9.1.2, page 183).

Functions managing lists handle an array of 2 elements (of type PlTerm) containing the terms corresponding to the head and the tail of the list. For the empty list NULL is passed as the array. These functions require to flatten a list in each sub-list. To simplify the management of proper lists (i.e. lists terminated by []) a set of functions is provided that handle the number of elements of the list (an integer) and an array whose elements (of type PlTerm) are the elements of the list. The caller of these functions must provide the array.

Functions managing compound terms handle a functor (the principal functor of the term), an arity N ≥ 0 and an array of N elements (of type PlTerm) containing the sub-terms of the compound term. Since a list is a special case of compound term (functor = '.' and arity=2) it is possible to use any function managing compound terms to deal with a list but the error detection is not the same. Indeed many functions check if the Prolog argument is correct. The name of a read or unify function checking the Prolog arguments is of the form Name Check(). For each of these functions there is a also check-free version called Name(). We then only present the name of checking functions.

9.2.2 Managing Prolog atoms

Each atom has a unique internal key which corresponds to its index in the GNU Prolog atom table. It is possible to obtain the information about an atom and to create new atoms using:

     char *Atom_Name           (int atom)
     int   Atom_Length         (int atom)
     Bool  Atom_Needs_Quote    (int atom)
     Bool  Atom_Needs_Scan     (int atom)
     Bool  Is_Valid_Atom       (int atom)
     int   Create_Atom         (char *str)
     int   Create_Allocate_Atom(char *str)
     int   Find_Atom           (char *str)
     int   ATOM_CHAR           (char c)
     int   atom_nil
     int   atom_false
     int   atom_true
     int   atom_end_of_file

The macro Atom Name(atom) returns the internal string of atom (this string should not be modified). The function Atom Lengh(atom) returns the length (of the name) of atom.

The function Atom Needs Scan(atom) indicates if the canonical form of atom needs to be quoted as done by writeq/2 (section 7.14.6, page 96). In that case Atom Needs Scan(atom) indicates if this simply comes down to write quotes around the name of atom or if it necessary to scan each character of the name because there are some non-printable characters (or included quote characters). The function Is Valid Atom(atom) is true only if atom is the internal key of an existing atom.

The function Create Atom(str) adds a new atom whose name is the content of str to the system and returns its internal key. If the atom already exists its key is simply returned. The string str passed to the function should not be modified later. The function Create Allocate Atom(str) is provided when this condition cannot be ensured. It simply makes a dynamic copy of str.

The function Find Atom(str) returns the internal key of the atom whose name is str or -1 if does not exist.

All atoms corresponding to a single character already exist and their key can be obtained via the macro ATOM CHAR. For instance ATOM CHAR('.') is the atom associated to '.' (this atom is the functor of lists). The other variables correspond to the internal key of frequently used atoms: [], false, true and

end of file.

9.2.3 Reading Prolog terms

The name of all functions presented here are of the form Rd Name Check(). They all check the validity of the Prolog term to read emitting appropriate errors if necessary. Each function has a check-free version

called Rd Name().

Simple foreign types: for each simple foreign type (section 9.1.2, page 183) there is a read function (used by the interface when an input argument is provided):

     long    Rd_Integer_Check  (PlTerm term)

long    Rd_Positive_Check (PlTerm term)
double  Rd_Float_Check    (PlTerm term)
double  Rd_Number_Check   (PlTerm term)
int     Rd_Atom_Check     (PlTerm term)
int     Rd_Boolean_Check  (PlTerm term)
int     Rd_Char_Check     (PlTerm term)
int     Rd_In_Char_Check  (PlTerm term)
int     Rd_Code_Check     (PlTerm term)
int     Rd_In_Code_Check  (PlTerm term)
int     Rd_Byte_Check     (PlTerm term)
int     Rd_In_Byte_Check  (PlTerm term)
char   *Rd_String_Check   (PlTerm term)
char   *Rd_Chars_Check    (PlTerm term)
char   *Rd_Codes_Check    (PlTerm term)
int     Rd_Chars_Str_Check(PlTerm term, char *str)
int     Rd_Codes_Str_Check(PlTerm term, char *str)

All functions returning a C string (char *) use a same buffer. The function Rd Chars Str Check() is similar to Rd Chars Check() but accepts as argument a string to store the result and returns the length of that string (which is also the length of the Prolog list). Similarly for Rd Codes Str Check().

Complex terms: the following functions return the sub-arguments (terms) of complex terms as an array of PlTerm except Rd Proper List Check() which returns the size of the list read (and initializes the array element). Refer to the introduction of this section for more information about the arguments of complex functions (section 9.2.1, page 190).

int     Rd_Proper_List_Check(PlTerm term, PlTerm *arg)
PlTerm *Rd_List_Check       (PlTerm term)
PlTerm *Rd_Compound_Check   (PlTerm term, int *functor, int *arity)
PlTerm *Rd_Callable_Check   (PlTerm term, int *functor, int *arity)

9.2.4 Unifying Prolog terms

The name of all functions presented here are of the form Un Name Check(). They all check the validity of the Prolog term to unify emitting appropriate errors if necessary. Each function has a check-free version

called Un Name().

Simple foreign types: for each simple foreign type (section 9.1.2, page 183) there is an unify function (used by the interface when an output argument is provided):

Bool Un_Integer_Check (long n,    PlTerm term)
Bool Un_Positive_Check(long n,    PlTerm term)
Bool Un_Float_Check   (double n,  PlTerm term)
Bool Un_Number_Check  (double n,  PlTerm term)
Bool Un_Atom_Check    (int atom,  PlTerm term)
Bool Un_Boolean_Check (int b,     PlTerm term)
Bool Un_Char_Check    (int c,     PlTerm term)
Bool Un_In_Char_Check (int c,     PlTerm term)
Bool Un_Code_Check    (int c,     PlTerm term)
Bool Un_In_Code_Check (int c,     PlTerm term)
Bool Un_Byte_Check    (int b,     PlTerm term)
Bool Un_In_Byte_Check (int b,     PlTerm term)
Bool Un_String_Check  (char *str, PlTerm term)
Bool Un_Chars_Check   (char *str, PlTerm term)
Bool Un_Codes_Check   (char *str, PlTerm term)

The function Un Number Check(n, term) unifies term with an integer if n is an integer, with a floating point number otherwise. The function Un String Check(str, term) creates the atom corresponding to

str and then unifies term with it (same as Un Atom Check(Create Allocate Atom(str), term)).

Complex terms: the following functions accept the sub-arguments (terms) of complex terms as an array of PlTerm. Refer to the introduction of this section for more information about the arguments of complex functions (section 9.2.1, page 190).

Bool Un_Proper_List_Check(int size, PlTerm *arg, PlTerm term)
Bool Un_List_Check       (PlTerm *arg, PlTerm term)
Bool Un_Compound_Check   (int functor, int arity, PlTerm *arg,
                          PlTerm term)
Bool Un_Callable_Check   (int functor, int arity, PlTerm *arg,
                          PlTerm term)

All these functions check the type of the term to unify and return the result of the unification. Generally if an unification fails the C function returns FALSE to enforce a failure. However if there are several arguments to unify and if an unification fails then the C function returns FALSE and the type of other arguments has not been checked. Normally all error cases are tested before doing any work to be sure that the predicate fails/succeeds only if no error condition is satisfied. So a good method is to check if the validity of all arguments to unify and later to do the unification (using check-free functions). Obviously if there is only one to unify it is more efficient to use a unify function checking the argument. For the other cases the interface provides a set of functions to check the type of a term.

Simple foreign types: for each simple foreign type (section 9.1.2, page 183) there is check-for-unification function (used by the interface when an output argument is provided):

void Check_For_Un_Integer (PlTerm term)
void Check_For_Un_Positive(PlTerm term)
void Check_For_Un_Float   (PlTerm term)
void Check_For_Un_Number  (PlTerm term)
void Check_For_Un_Atom    (PlTerm term)
void Check_For_Un_Boolean (PlTerm term)
void Check_For_Un_Char    (PlTerm term)
void Check_For_Un_In_Char (PlTerm term)
void Check_For_Un_Code    (PlTerm term)
void Check_For_Un_In_Code (PlTerm term)
void Check_For_Un_Byte    (PlTerm term)
void Check_For_Un_In_Byte (PlTerm term)
void Check_For_Un_String  (PlTerm term)
void Check_For_Un_Chars   (PlTerm term)
void Check_For_Un_Codes   (PlTerm term)

Complex terms: the following functions check the validity of complex terms:

void Check_For_Un_List    (PlTerm term)
void Check_For_Un_Compound(PlTerm term)
void Check_For_Un_Callable(PlTerm term)
void Check_For_Un_Variable(PlTerm term)

The function Check For Un List(term) checks if term can be unified with a list. This test is done for the entire list (not only for the functor/arity of term but also recursively on the tail of the list). The function Check For Un Variable(term) ensures that term is not currently instantiated. These functions can be defined using functions to test the type of a Prolog term (section 9.2.6, page 194) and functions to raise Prolog errors (section 9.3, page 196). For instance Check For Un List(term) is defined as follows:

void Check_For_Un_List(PlTerm term)
{
 if (!Blt_List_Or_Partial_List(term))
     Pl_Err_Type(type_list, term);
}

9.2.5 Creating Prolog terms

These functions are provided to creates Prolog terms. Each function returns a PlTerm containing the created term.

Simple foreign types: for each simple foreign type (section 9.1.2, page 183) there is a creation function:

PlTerm Mk_Integer (long n)
PlTerm Mk_Positive(long n)
PlTerm Mk_Float   (double n)
PlTerm Mk_Number  (double n)
PlTerm Mk_Atom    (int atom)
PlTerm Mk_Boolean (int b)
PlTerm Mk_Char    (int c)
PlTerm Mk_In_Char (int c)
PlTerm Mk_Code    (int c)
PlTerm Mk_In_Code (int c)
PlTerm Mk_Byte    (int b)
PlTerm Mk_In_Byte (int b)
PlTerm Mk_String  (char *str)
PlTerm Mk_Chars   (char *str)
PlTerm Mk_Codes   (char *str)

The function Mk Number(n, term) initializes term with an integer if n is an integer, with a floating point number otherwise. The function Mk String(str) first creates an atom corresponding to str and then returns that Prolog atom (i.e. equivalent to Mk Atom(Create Allocate Atom(str))).

Complex terms: the following functions accept the sub-arguments (terms) of complex terms as an array of PlTerm. Refer to the introduction of this section for more information about the arguments of complex functions (section 9.2.1, page 190).

PlTerm Mk_Proper_List(int size, PlTerm *arg)
PlTerm Mk_List       (PlTerm *arg)
PlTerm Mk_Compound   (int functor, int arity, PlTerm *arg)
PlTerm Mk_Callable   (int functor, int arity, PlTerm *arg)

9.2.6 Testing the type of Prolog terms

The following functions test the type of a Prolog term. Each function corresponds to a type testing built-in predicate (section 7.1.1, page 49).

Bool Blt_Var                 (PlTerm term)
Bool Blt_Non_Var             (PlTerm term)
Bool Blt_Atom                (PlTerm term)
Bool Blt_Integer             (PlTerm term)
Bool Blt_Float               (PlTerm term)
Bool Blt_Number              (PlTerm term)
Bool Blt_Atomic              (PlTerm term)
Bool Blt_Compound            (PlTerm term)
Bool Blt_Callable            (PlTerm term)
Bool Blt_List                (PlTerm term)
Bool Blt_Partial_List        (PlTerm term)
Bool Blt_List_Or_Partial_List(PlTerm term)
Bool Blt_Fd_Var              (PlTerm term)
Bool Blt_Non_Fd_Var          (PlTerm term)
Bool Blt_Generic_Var         (PlTerm term)

     Bool Blt_Non_Generic_Var     (PlTerm term)
     int  Type_Of_Term            (PlTerm term)
     int  List_Length             (PlTerm list)

The function Type Of Term(term) returns the type of term, the following constants can be used to test this type (e.g. in a switch instruction):

• PLV: Prolog variable.

• FDV: finite domain variable.

• INT: integer.

• FLT: floating point number.

• ATM: atom.

• LST: list.

• STC: structure

The tag LST means a term whose principal functor is '.' and whose arity is 2 (recall that the empty list is the atom []). The tag STC means any other compound term.

The function List Length(list) returns the number of elements of the list (0 for the empty list). If list is not a list this function returns -1.

9.2.7 Comparing Prolog terms

The following functions compares Prolog terms. Each function corresponds to a comparison built-in predicate (section 7.3.2, page 51).

     Bool Blt_Term_Eq (PlTerm term1, PlTerm term2)
     Bool Blt_Term_Neq(PlTerm term1, PlTerm term2)
     Bool Blt_Term_Lt (PlTerm term1, PlTerm term2)
     Bool Blt_Term_Lte(PlTerm term1, PlTerm term2)
     Bool Blt_Term_Gt (PlTerm term1, PlTerm term2)
     Bool Blt_Term_Gte(PlTerm term1, PlTerm term2)

All these functions are based on a general comparison function returning a negative integer if term1 is less than term2, 0 if they are equal and a positive integer otherwise:

     int Term_Compare(PlTerm term1, PlTerm term2)

9.2.8 Copying Prolog terms

The following functions make a copy of a Prolog term:

     void Copy_Term           (PlTerm *dst_adr, PlTerm *src_adr)
     void Copy_Contiguous_Term(PlTerm *dst_adr, PlTerm *src_adr)
     int  Term_Size           (PlTerm term)

The function Copy Term(dst adr, src adr) makes a copy of the term located at src adr and stores it from the address given by dst adr. The result is a contiguous term. If it can be ensured that the source term is a contiguous term (i.e. result of a previous copy) the function Copy Contiguous Term() can be used instead (it is faster). In any case, sufficient space should be available for the copy (i.e. from dst adr). The function Term Size(term) returns the number of PlTerm needed by term.

9.2.9 Comparing and evaluating arithmetic expressions

The following functions compare arithmetic expressions. Each function corresponds to a comparison built-in predicate (section 7.6.3, page 60).

     Bool Blt_Eq (PlTerm expr1, PlTerm expr2)
     Bool Blt_Neq(PlTerm expr1, PlTerm expr2)
     Bool Blt_Lt (PlTerm expr1, PlTerm expr2)
     Bool Blt_Lte(PlTerm expr1, PlTerm expr2)
     Bool Blt_Gt (PlTerm expr1, PlTerm expr2)
     Bool Blt_Gte(PlTerm expr1, PlTerm expr2)

The following function evaluates the expression expr and stores its result as a Prolog number (integer or floating point number) in result:

     void Math_Load_Value(PlTerm expr, PlTerm *result)

This function can be followed by a read function (section 9.2.3, page 191) to obtain the result.

9.3 Raising Prolog errors

The following functions allows a C function to raise a Prolog error. Refer to the section concerning Prolog errors for more information about the effect of raising an error (section 5.3, page 37).

9.3.1 Managing the error context

When one of the following error function is invoked it refers to the implicit error context (section 5.3.1, page 37). This context indicates the name and the arity of the concerned predicate. When using a foreign/2 declaration this context is set by default to the name and arity of the associated Prolog predicate. This can be controlled using the bip name option (section 9.1.2, page 183). In any case, the following functions can also be used to modify this context:

     void Set_C_Bip_Name  (char *functor, int arity)
     void Unset_C_Bip_Name(void)

The function Set C Bip Name(functor, arity) initializes the context of the error with functor and arity (if arity<0 only functor is significant). The function Unset C Bip Name() removes such an initialization (the context is then reset to the last Functor/Arity set by a call to set bip name/2 (section 7.22.3, page 135). This is useful when writing a C routine to define a context for errors occurring in this routine and, before exiting to restore the previous context.

9.3.2 Instantiation error

The following function raises an instantiation error (section 5.3.2, page 37):

     void Pl Err Instantiation(void)

9.3.3 Type error

The following function raises a type error (section 5.3.3, page 38):

     void Pl Err Type(int atom type, PlTerm culprit)

atom type is (the internal key of) the atom associated to the expected type. For each type name T there is a corresponding predefined atom stored in a global variable whose name is of the form type T . culprit is the argument which caused the error.

Example: x is an atom while an integer was expected: Pl Err Type(type integer, x).

9.3.4 Domain error

The following function raises a domain error (section 5.3.4, page 38):

     void Pl Err Domain(int atom domain, PlTerm culprit)

atom domain is (the internal key of) the atom associated to the expected domain. For each domain name D there is a corresponding predefined atom stored in a global variable whose name is of the form domain D . culprit is the argument which caused the error.

Example: x is < 0 but should be ≥ 0: Pl Err Domain(domain not less than zero, x).

9.3.5 Existence error

The following function raises an existence error (section 5.3.5, page 39):

     void Pl Err Existence(int atom object, PlTerm culprit)

atom object is (the internal key of) the atom associated to the type of the object. For each object name O there is a corresponding predefined atom stored in a global variable whose name is of the form existence O . culprit is the argument which caused the error.

Example: x does not refer to an existing source: Pl Err Existence(existence source sink, x).

9.3.6 Permission error

The following function raises a permission error (section 5.3.6, page 39):

     void Pl Err Permission(int atom operation, int atom permission, PlTerm culprit)

atom operation is (the internal key of) the atom associated to the operation which caused the error. For each operation name O there is a corresponding predefined atom stored in a global variable whose name is of the form permission operation O . atom permission is (the internal key of) the atom associated to the tried permission. For each permission name P there is a corresponding predefined atom stored in a global variable whose name is of the form permission type P . culprit is the argument which caused the error.

Example: reading from an output stream x: Pl Err Permission(permission operation input,
permission type stream, x).

9.3.7 Representation error

The following function raises a representation error (section 5.3.7, page 39):

     void Pl Err Representation(int atom limit)

atom limit is (the internal key of) the atom associated to the reached limit. For each limit name L there is a corresponding predefined atom stored in a global variable whose name is of the form representation L .

Example: an arity too big occurs: Pl Err Representation(representation max arity).

9.3.8 Evaluation error

The following function raises an evaluation error (section 5.3.8, page 40):

     void Pl Err Evaluation(int atom error)

atom error is (the internal key of) the atom associated to the error. For each evaluation error name E there is a corresponding predefined atom stored in a global variable whose name is of the form

evaluation E .
Example: a division by zero occurs: Pl Err Evaluation(evluation zero divisor).

9.3.9 Resource error

The following function raises a resource error (section 5.3.9, page 40):

     void Pl Err Resource(int atom resource)

atom resource is (the internal key of) the atom associated to the resource. For each resource error name R there is a corresponding predefined atom stored in a global variable whose name is of the form

resource R .
Example: too many open streams: Pl Err Resource(resource too many open streams).

9.3.10 Syntax error

The following function raises a syntax error (section 5.3.10, page 40):

     void Pl Err Syntax(int atom error)

atom error is (the internal key of) the atom associated to the error. There is no predefined syntax error atoms.

Example: a / is expected: Pl Err Syntax(Create Atom("/ expected")).

The following function emits a syntax error according to the value of the syntax error Prolog flag (section 7.22.1, page 133). This function can then return (if the value of the flag is either warning or fail). In that case the calling function should fail (e.g. returning FALSE). This function accepts a file name (the empty string C "" can be passed), a line and column number and an error message string. Using this function makes it possible to further call the built-in predicate syntax error info/4 (section 7.14.4, page 95):

     void Emit Syntax Error(char *file name, int line, int column, char *message)
Example: a / is expected: Emit Syntax Error("data", 10, 30, "/ expected").

9.3.11 System error

The following function raises a system error (4.3.11, page *):

     void Pl Err System(int atom error)

atom error is (the internal key of) the atom associated to the error. There is no predefined system error atoms.

Example: an invalid pathname is given: Pl Err System(Create Atom("invalid path name")).

The following function emits a system error associated to an operating system error according to the value of the os error Prolog flag (section 7.22.1, page 133). This function can then return (if the value of the flag is either warning or fail). In that case the calling function should fail (e.g. returning FALSE). This function uses the value of the errno C library variable:

     void Os Error(void)

Example: a call to the C Unix function chdir(3) returns -1: Os Error().

9.4 Calling Prolog from C

9.4.1 Introduction

The following functions allows a C function to call a Prolog predicate:

     void   Pl_Query_Begin        (Bool recoverable)
     int    Pl_Query_Call         (int functor, int arity, PlTerm *arg)
     int    Pl_Query_Next_Solution(void)
     void   Pl_Query_End          (int op)
     PlTerm Pl_Get_Exception      (void)
     void   Pl_Exec_Continuation  (int functor, int arity, PlTerm *arg)

The invocation of a Prolog predicate should be done as follows:

• open a query using Pl Query Begin()

• compute the first solution using Pl Query Call()

• eventually compute next solutions using Pl Query Next Solution()

• close the query using Pl Query End()

The function Pl Query Begin(recoverable) is used to initialize a query. The argument recoverable shall be set to TRUE if the user wants to recover, at the end of the query, the memory space consumed by the query (in that case an additional choice-point is created). All terms created in the heap, e.g. using Mk ... family functions (section 9.2.5, page 194), after the invocation of Pl Query Begin() can be recovered when calling Pl Query End(TRUE) (see below).

The function Pl Query Call(functor, arity, arg) calls a predicate passing arguments. It is then used to compute the first solution. The arguments functor, arity and arg are similar to those of the functions handling complex terms (section 9.2.1, page 190). This function returns:

• PL FAILURE (a constant equal to FALSE, i.e. 0) if the query fails.

• PL SUCCESS (a constant equal to TRUE, i.e. 1) in case of success. In that case the argument array arg can be used to obtain the unification performed by the query.

• PL EXCEPTION (a constant equal to 2). In that case function Pl Get Exception() can be used to obtained the exceptional term raised by throw/1 (section 6.2.4, page 47).

The function Pl Query Next Solution() is used to compute a new solution. It must be only used if the result of the previous solution was PL SUCCESS. This functions returns the same kind of values as

Pl Query Call() (see above).

The function Pl Query End(op) is used to finish a query. This function mainly manages the remaining alternatives of the query. However, even if the query has no alternatives this function must be used to correctly finish the query. The value of op is:

• PL RECOVER: to recover the memory space consumed by the query. After that the state of Prolog stacks is exactly the same as before opening the query. To use this option the query must have been initialized specifying TRUE for recoverable (see above).

• PL CUT: to cut remaining alternatives. The effect of this option is similar to a cut after the query.

• PL KEEP FOR PROLOG: to keep the alternatives for Prolog. This is useful when the query was invoked in a foreign C function. In that case, when the predicate corresponding to the C foreign function is invoked a query is executed and the remaining alternatives are then available as alternatives of that predicate.

Note that several queries can be nested since a stack of queries is maintained. For instance, it is possible to call a query and before terminating it to call another query. In that case the first execution of Pl Query End() will finish the second query (i.e. the inner) and the next execution of Pl Query End() will finish the first query.

Finally, the function Pl Exec Continuation(functor, arity, arg) replaces the current calculus by the execution of the specified predicate. The arguments functor, arity and arg are similar to those of the functions handling complex terms (section 9.2.1, page 190).

9.4.2 Example: my call/1 - a call/1 clone

We here define a predicate my call(Goal) which acts like call(Goal) except that we do not handle exceptions (if an exception occurs the goal simply fails):

In the prolog file examp.pl:

     :- foreign(my call(term)).

In the C file examp c.c:

     #include <string.h>
     #include "gprolog.h"
     Bool
     my_call(PlTerm goal)
     {
       PlTerm *arg;
       int functor, arity;
       int result;
       arg = Rd_Callable_Check(goal, &functor, &arity);
       Pl_Query_Begin(FALSE);
       result = Pl_Query_Call(functor, arity, arg);
       Pl_Query_End(PL_KEEP_FOR_PROLOG);
       return (result == PL_SUCCESS);
     }

The compilation produces an executable called examp:

     % gplc examp.pl examp c.c

Examples of use:

| ?- my call(write(hello)).
hello
| ?- my call(for(X,1,3)).

X = 1 ?  (here the user presses ; to compute another solution)
X = 2 ?  (here the user presses ; to compute another solution)
X = 3  (here the user is not prompted since there is no more alternative)

| ?- my call(1).
{exception: error(type error(callable,1),my call/1)}
| ?- my call(call(1)).
no

When my call(1) is called an error is raised due to the use of Rd Callable Check(). However the error raised by my call(call(1)) is ignored and FALSE (i.e. a failure) is returned by the foreign function.

To really simulate the behavior of call/1 when an exception is recovered it should be re-raised to be captured by an earlier handler. The idea is then to execute a throw/1 as the continuation. This is what it is done by the following code:

#include <string.h>
#include "gprolog.h"
Bool
my_call(PlTerm goal)
{
  PlTerm *args;
  int functor, arity;
  int result;
  args = Rd_Callable_Check(goal, &functor, &arity);
  Pl_Query_Begin(FALSE);
  result = Pl_Query_Call(functor, arity, args);
  Pl_Query_End(PL_KEEP_FOR_PROLOG);
  if (result == PL_EXCEPTION)
    {
      PlTerm except = Pl_Get_Exception();
      Pl_Exec_Continuation(Find_Atom("throw"), 1, &except);
    }
  return result;
}

The following code propagates the error raised by call/1.

| ?- my call(call(1)).
{exception: error(type error(callable,1),my call/1)}

Finally note that a simpler way to define my call/1 is to use Pl Exec Continuation() as follows:

#include <string.h>
#include "gprolog.h"
Bool
my_call(PlTerm goal)

{
PlTerm *args;
int functor, arity;
args = Rd_Callable_Check(goal, &functor, &arity);
Pl_Exec_Continuation(functor, arity, args);
return TRUE;
}

9.4.3 Example: recovering the list of all operators

We here define a predicate all op(List) which unifies List with the list of all currently defined operators

as would be done by: findall(X,current op( , ,X),List).

In the prolog file examp.pl:

:- foreign(all op(term)).

In the C file examp c.c:

#include <string.h>
#include "gprolog.h"
Bool
all_op(PlTerm list)
{
PlTerm op[1024];
PlTerm args[3];
int n = 0;
int result;
Pl_Query_Begin(TRUE);
args[0] = Mk_Variable();
args[1] = Mk_Variable();
args[2] = Mk_Variable();
result = Pl_Query_Call(Find_Atom("current_op"), 3, args);
while (result)
  {
    op[n++] = Mk_Atom(Rd_Atom(args[2])); /* arg #2 is the name of the op */
    result = Pl_Query_Next_Solution();
  }
Pl_Query_End(PL_RECOVER);
return Un_Proper_List_Check(n, op, list);
}

Note that we know here that there is no source for exception. In that case the result of Pl Query Call and Pl Query Next Solution can be considered as a boolean.

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Example of use:

| ?- all_op(L).
L = [:-,:-,\=,=:=,#>=,#<#,@>=,-->,mod,#>=#,**,*,+,+,',',...]

     | ?- findall(X,current_op(_,_,X),L).
     L = [:-,:-,\=,=:=,#>=,#<#,@>=,-->,mod,#>=#,**,*,+,+,',',...]

9.5 Defining a new C main() function

GNU Prolog allows the user to define his own main() function. This can be useful to perform several tasks before starting the Prolog engine. To do this simply define a classical main(argc, argv) function. The following functions can then be used:

     int  Start_Prolog         (int argc, char *argv[])
     void Stop_Prolog          (void)
     void Reset_Prolog         (void)
     Bool Try_Execute_Top_Level(void)

The function Start Prolog(argc, argv) initializes the Prolog engine (argc and argv are the commandline variables). This function collects all linked objects (issued from the compilation of Prolog files) and initializes them. The initialization of a Prolog object file consists in adding to appropriate tables new atoms, new predicates and executing its system directives. A system directive is generated by the Prolog to WAM compiler to reflect a (user) directive executed at compile-time such as op/3 (section 6.1.10, page 44). Indeed, when the compiler encounters such a directive it immediately executes it and also generates a system directive to execute it at the start of the executable. When all system directives have been executed the Prolog engine executes all initialization directives defined with initialization/1 (section 6.1.13, page 45). The function returns the number of user directives (i.e. initialization/1) executed. This function must be called only once.

The function Stop Prolog() stops the Prolog engine. This function must be called only once after all Prolog treatment have been done.

The function Reset Prolog() reinitializes the Prolog engine (i.e. reset all Prolog stacks).

The function Try Execute Top Level() executes the top-level if linked (section 3.4.3, page 22) and returns TRUE. If the top-level is not present the functions returns FALSE.

Here is the definition of the default GNU Prolog main() function:

     int
     Main_Wrapper(int argc, char *argv[])
     {
       int nb_user_directive;
       Bool top_level;
       nb_user_directive = Start_Prolog(argc, argv);
       top_level = Try_Execute_Top_Level();
       Stop_Prolog();
       if (top_level || nb_user_directive)
         return 0;
       fprintf(stderr,
               "Warning: no initial goal executed\n"
               "   use a directive :- initialization(Goal)\n"
               "   or remove the link option --no-top-level"

          " (or --min-bips or --min-size)\n");
  return 1;
}
int
main(int argc, char *argv[])
{
  return Main_Wrapper(argc, argv);
}

Note that under some circumstances it is necessary to encapsulate the code of main() inside an intermediate function called by main(). Indeed, some C compilers (e.g. gcc) treats main() particularly, producing an uncompatible code w.r.t GNU Prolog. So it is a good idea to always use a wrapper function as shown above.

9.5.1 Example: asking for ancestors

In this example we use the following Prolog code (in a file called new main.pl):

parent(bob,   mary).
parent(jane,  mary).
parent(mary,  peter).
parent(paul,  peter).
parent(peter, john).
anc(X, Y):-
        parent(X, Y).
anc(X, Z) :-
        parent(X, Y),
        anc(Y, Z).

The following file (called new main c.c) defines a main() function readinf the name of a person and displaying all successors of that person. This is equivalent to the Prolog query: anc(Result, Name).

static int
Main_Wrapper(int argc, char *argv[])
{
  int func;
  WamWord arg[10];
  char str[100];
  char *sol[100];
  int i, nb_sol = 0;
  Bool res;
  Start_Prolog(argc, argv);
  func = Find_Atom("anc");
  for (;;)
    {
      printf("\nEnter a name (or 'end' to finish): ");
      scanf("%s", str);
      if (strcmp(str, "end") == 0)
        break;

      Pl_Query_Begin(TRUE);
      arg[0] = Mk_Variable();
      arg[1] = Mk_String(str);
      nb_sol = 0;
      res = Pl_Query_Call(func, 2, arg);
      while (res)
        {
          sol[nb_sol++] = Rd_String(arg[0]);
          res = Pl_Query_Next_Solution();
        }
      Pl_Query_End(PL_RECOVER);
      for (i = 0; i < nb_sol; i++)
        printf("  solution: %s\n", sol[i]);
      printf("%d solution(s)\n", nb_sol);
    }
  Stop_Prolog();
  return 0;
}
int
main(int argc, char *argv[])
{
  return Main_Wrapper(argc, argv);
}

The compilation produces an executable called new main:

% gplc new main.pl new main c.c

Examples of use:

Enter a name (or 'end' to finish): john
  solution: peter
  solution: bob
  solution: jane
  solution: mary
  solution: paul
5 solution(s)
Enter a name (or 'end' to finish): mary
  solution: bob
  solution: jane
2 solution(s)
Enter a name (or 'end' to finish): end

206  9 INTERFACING PROLOG AND C

References

[1] H. Aït-Kaci. Warren’s Abstract Machine, A Tutorial Reconstruction.

Logic Programming Series, MIT Press, 1991.

http://www.isg.sfu.ca/~hak/documents/wam.html

[2] W.F. Clocksin and C.S. Mellish. Programming in Prolog, Springer-Verlag, 1981.

[3] P. Codognet and D. Diaz. wamcc: Compiling Prolog to C.

In 12th International Conference on Logic Programming, Tokyo, Japan, MIT Press, 1995.

ftp://ftp.inria.fr/INRIA/Projects/loco/publications/wamcc/wamcc.ps

[4] P. Codognet and D. Diaz. Compiling Constraint in clp(FD).

Journal of Logic Programming, Vol. 27, No. 3, June 1996.

ftp://ftp.inria.fr/INRIA/Projects/loco/publications/clp fd/long clp fd.ps

[5] Information technology - Programming languages - Prolog - Part 1: General Core.

ISO/IEC 13211-1, 1995. http://www.logic-programming.org/prolog std.html

[6] J. Jaffar and J-L. Lassez. Constraint Logic Programming.

In Principles Of Programming Languages, Munich, Germany, January 1987.

[7] P. Van Hentenryck. Constraint Satisfaction in Logic Programming.

Logic Programming Series, The MIT Press, 1989.

[8] D. H. D. Warren. An Abstract Prolog Instruction Set.

Technical Report 309, SRI International, Oct. 1983.

208  REFERENCES

\underline{INDEX 209}

Index

!/0, 46, 47

'.'/2, 136
(',')/2, 46
(-->)/2, 109
(->)/2, 46
(;)/2, 46
(=)/2, 50
(=..)/2, 53
(=:=)/2, 60

(==)/2, 51, 122

(=<)/2, 60
(=\=)/2, 60
(@=<)/2, 51
(@<)/2, 51
(@>)/2, 51
(@>=)/2, 51

(#/\)/2 (FD), 174 (#=)/2 (FD), 171

(#==>)/2 (FD), 174

(#=#)/2 (FD), 172 (#=<)/2 (FD), 171

(#=<#)/2 (FD), 172

(##)/2 (FD), 174 (#<)/2 (FD), 171

(#<=>)/2 (FD), 174

(#<#)/2 (FD), 172 (#>)/2 (FD), 171 (#>=)/2 (FD), 171

(#>=#)/2 (FD), 172

(#>#)/2 (FD), 172 (#\)/1 (FD), 174 (#\/)/2 (FD), 174

(#\/\)/2 (FD), 174

(#\=)/2 (FD), 171

(#\==>)/2 (FD), 174
(#\=#)/2 (FD), 172
(#\<=>)/2 (FD), 174
(#\\/)/2 (FD), 174
(is)/2, 59
(<)/2, 60
(>)/2, 60
(>=)/2, 60
(\+)/1, 113
(\=)/2, 50
(\==)/2, 51

--, 13

--assembly, 22
--aux-father, 27
--aux-father2, 27
--c-compiler, 23
--cmd-line, 27
--comment, 23, 23
--compile-msg, 22
--cstr-size, 23

                                               --encode, 27
                                               --entry-goal, 13
                                               --fast-math, 22, 59
                                               --fd-to-c, 22
                                               --fixed-sizes, 19, 23
                                               --foreign-only, 22
                                               --global-size, 23

--help, 13, 22, 27

                                               --init-goal, 13
                                               --keep-void-inst, 22
                                               --local-size, 19, 23
                                               --min-bips, 23
                                               --min-fd-bips, 23
                                               --min-pl-bips, 23
                                               --min-reg-opt, 22
                                               --min-size, 23
                                               --mini-assembly, 22
                                               --no-call-c, 22
                                               --no-debugger, 23, 23
                                               --no-decode-hexa, 22
                                               --no-del-temp, 22
                                               --no-fd-lib, 23
                                               --no-inline, 22
                                               --no-opt-last-subterm, 22
                                               --no-redef-error, 22
                                               --no-reg-opt, 22
                                               --no-reorder, 22
                                               --no-singl-warn, 22
                                               --no-susp-warn, 22
                                               --no-top-level, 23
                                               --object, 22
                                               --output, 22
                                               --pl-state, 22, 136
                                               --printf, 27
                                               --query-goal, 13
                                               --relax, 27
                                               --statistics, 22
                                               --strip, 23
                                               --temp-dir, 22
                                               --trail-size, 23
                                               --verbose, 22
                                               --version, 13, 22, 27
                                               --wam-for-byte-code, 22
                                               --wam-for-native, 22

\underline{210 INDEX}

-A, 23 -C, 23 -F, 22 -H, 27 -L, 23 -M, 22 -P, 27 -S, 22 -W, 22 -c, 22 -h, 22 -o, 22 -s, 23 -v, 22 -w, 22

abolish/1, 63

abort/0, 17, 32, 112

absolute file name (property), 147

absolute file name/2, 44, 70, 136, 137, 141,

142, 144–149

add linedit completion/1, 162
add stream alias/2, 67, 79
add stream mirror/2, 68, 79

alias (option), 71 alias (property), 73 append (mode), 70

append/1, 107
append/3, 121
architecture/1, 151
arg/3, 53

argument selector, 126

argument counter/1, 142
argument list/1, 14, 143
argument value/2, 14, 143
asserta/1, 61
assertz/1, 61
at end of stream/0, 74
at end of stream/1, 74
atom/1, 49
atom chars/2, 116
atom codes/2, 116
atom concat/3, 115
atom hash/2, 119
atom length/2, 114
atom property/2, 120
atomic/1, 49

back quotes (flag), 14, 91, 134, 136

back quotes (token), 94

backtracks (FD option), 180

bagof/3, 66

binary (option), 70, 81

bind variables/2, 56

bip name (option), 183, 196 block (option), 71, 82 block device (permission), 147 bof (whence), 76 boolean (option), 183, 184 bounded (flag), 133 bounds (FD option), 179 break/0, 17, 32, 112 buffering (option), 71 buffering (property), 74 built in (property), 43, 65 built in/0 (directive), 43 built in/1 (directive), 43 built in fd (property), 43, 65 built in fd/0 (directive), 43 built in fd/1 (directive), 43

                                               call/1, 47
                                               call/2, 113
                                               call with args/1-11, 113
                                               callable/1, 49

catch/3, 29, 37, 47

                                               change directory/1, 144
                                               char code/2, 116, 189

char conversion (flag), 91, 103, 134, 136

                                               char conversion/2 (directive), 44
                                               char conversion/2, 45, 102
                                               character count/2, 76

character device (permission), 147 choice size (option), 183, 186

                                               clause/2, 63
                                               close/1, 72

close/2, 72, 153, 158

                                               close input atom stream/1, 83
                                               close input chars stream/1, 83
                                               close input codes stream/1, 83
                                               close output atom stream/2, 84
                                               close output chars stream/2, 84
                                               close output codes stream/2, 84
                                               compare/3, 52

completion, 18, 162, 163

                                               compound/1, 49

consult/1, 16, 17, 20, 21, 136

                                               copy term/2, 54
                                               cpu time/1, 139
                                               create pipe/2, 154

current (whence), 76

                                               current alias/2, 79
                                               current atom/1, 120
                                               current bip name/2, 37, 135
                                               current char conversion/2, 103
                                               current input/1, 69
                                               current mirror/2, 80
                                               current op/3, 102
                                               current output/1, 69
                                               current predicate/1, 62, 64
                                               current prolog flag/2, 134
                                               current stream/1, 73
                                               date time/1, 149

debug (flag), 134

debug/0 (debug), 17, 29 debugging/0 (debug), 29, 32

                                               decompose file name/4, 141

Definite clause grammars, see DCG

                                               delete/3, 122
                                               delete directory/1, 144
                                               delete file/1, 146

directory (permission), 147

                                               directory files/2, 145

discontiguous/1 (directive), 42
display/1, 96

display/2, 96, 105, 106

display to atom/2, 105
display to chars/2, 106
display to codes/2, 106

double quotes (flag), 91, 134, 136 dynamic (property), 65 dynamic/1 (directive), 41, 60

end of stream (property), 74 end of term (option), 92

ensure linked/1 (directive), 43
ensure loaded/1 (directive), 44
environ/2, 144

eof (whence), 76 eof action (option), 71 eof action (property), 74 eof code (option), 71, 81 error (option), 71, 81, 92 escape sequence, 14, 121, 134 exclude (option), 56

exec/4, 153
exec/5, 153

execute (permission), 147

expand term/2, 111

extended (token), 94 extra-constrained, see extra cstr extra cstr (FD), 165, 170

fail (option), 92

fail/0, 46

fct name (option), 183

fd all different/1 (FD), 176
fd at least one/1 (FD), 175
fd at most one/1 (FD), 175
fd atleast/3 (FD), 177
fd atmost/3 (FD), 177
fd cardinality/2 (FD), 175, 178
fd cardinality/3 (FD), 175

fd dom/2 (FD), 169

fd domain/2 (FD), 168
fd domain/3 (FD), 167
fd domain bool/1 (FD), 167
fd element/3 (FD), 176
fd element var/3 (FD), 177
fd exactly/3 (FD), 177
fd has extra cstr/1 (FD), 170
fd has vector/1 (FD), 170
fd labeling/1 (FD), 179
fd labeling/2 (FD), 179, 180
fd labelingff/1 (FD), 179

fd max/2 (FD), 169

fd max integer (FD), 165, 166
fd max integer/1 (FD), 166
fd maximize/2 (FD), 180

fd min/2 (FD), 169

                                               fd minimize/2 (FD), 180
                                               fd not prime/1 (FD), 173
                                               fd only one/1 (FD), 175
                                               fd prime/1 (FD), 173
                                               fd relation/2 (FD), 178
                                               fd relationc/2 (FD), 178
                                               fd set vector max/1 (FD), 165, 167
                                               fd size/2 (FD), 169
                                               fd use vector/1 (FD), 170

fd var/1 (FD), 168

                                               fd vector max/1 (FD), 165, 166

fifo (permission), 147

                                               file exists/1, 146

file name (property), 73

                                               file permission/2, 146
                                               file property/2, 147
                                               find linedit completion/2, 163
                                               findall/3, 66

first fail (FD option), 179

flag, see Prolog flag

                                               float/1, 49
                                               flush output/0, 72
                                               flush output/1, 68, 72
                                               for/3, 114

force (option), 72

foreign/1 (directive), 45, 183 foreign/2 (directive), 45, 183

                                               fork prolog/1, 154
                                               format/2, 98

format/3, 98, 105, 106

                                               format to atom/3, 105
                                               format to chars/3, 106
                                               format to codes/3, 106

from (option), 56 full (debug), 30

                                               functor/3, 52

g array (global var.), 127

g array auto (global var.), 127 g array extend (global var.), 127

                                               g array size/2, 128
                                               g assign/2, 127
                                               g assignb/2, 127
                                               g dec/1, 129
                                               g dec/2, 129
                                               g dec/3, 129
                                               g deco/2, 129
                                               g inc/1, 129
                                               g inc/2, 129
                                               g inc/3, 129
                                               g inco/2, 129
                                               g link/2, 127
                                               g read/2, 128
                                               g reset bit/2, 130
                                               g set bit/2, 130
                                               g test reset bit/2, 130
                                               g test set bit/2, 130

generic var/1 (FD), 168
get/1, 108
get0/1, 108
get byte/1, 89
get byte/2, 68, 89
get char/1, 85
get char/2, 85
get code/1, 85
get code/2, 85, 86
get key/1, 86
get key/2, 86
get key no echo/1, 86
get key no echo/2, 86
get linedit prompt/1, 162
get print stream/1, 100
get seed/1, 140

gplc, 22, 24, 26, 136

half (debug), 30 halt/0, 13, 17, 112

halt/1, 112

hash (property), 120

hexgplc, 26
host name/1, 150
hostname address/2, 161

ignore ops (option), 96 include/1 (directive), 44 infix op (property), 120 initialization/1 (directive), 24, 25, 45, 203 input (property), 73

integer/1, 49
integer rounding function (flag), 59, 133

interpreter, see top-level

jump (option), 183, 184

keysort/1, 125
keysort/2, 125

largest (FD option), 179

last/2, 124
last modification (property), 148
last read start line column/2, 95

leash/1 (debug), 30, 31 length (property), 120

length/2, 124

line (option), 71, 82

line count/2, 77, 78
line position/2, 77

linedit, 18, 86, 162, 163

list/1, 49
list or partial list/1, 49
listing/0, 138

listing/1, 32, 99, 138 load/1, 17, 21, 23, 137 loose (debug), 30

lower upper/2, 116

MA, 20

make directory/1, 144

max (FD option), 179 max arity (flag), 133 max atom (flag), 119, 133 max depth (option), 97 max integer (flag), 133, 165

max list/2, 125

max regret (FD option), 179 max unget (flag), 88, 90, 133

member/2, 121
memberchk/2, 121

middle (FD option), 179 min (FD option), 179

min integer (flag), 133
min list/2, 125

mini-assembly, 11, 20, 26 mirror (option), 71 mirror (property), 73 mode (property), 73

most constrained (FD option), 179

multifile/1 (directive), 42

name/2, 118
name query vars/2, 55
name singleton vars/1, 55, 99

namevars (option), 16, 56, 96 native code (property), 65 needs quotes (property), 121 needs scan (property), 121

new atom/1, 119
new atom/2, 119
new atom/3, 119

next (option), 56

nl/0, 88
nl/1, 88

nodebug/0 (debug), 29, 32

non fd var/1 (FD), 168
non generic var/1 (FD), 168

none (debug), 30 none (option), 71, 82, 183, 184

nonvar/1, 49

nospy/1 (debug), 30, 32

nospyall/0 (debug), 30

notrace/0 (debug), 29

nth/3, 124
number/1, 49
number atom/2, 117
number chars/2, 117
number codes/2, 117

numbervars (option), 16, 56, 96

numbervars/1, 56, 99
numbervars/3, 56
once/1, 113

op/3 (directive), 44 op/3, 44, 100

open/3, 70

open/4, 67, 70, 81, 82, 157

open input atom stream/2, 83
open input chars stream/2, 83
open input codes stream/2, 83
open output atom stream/1, 84
open output chars stream/1, 84
open output codes stream/1, 84

os error (flag), 134, 199

os version/1, 150

output (property), 73

partial list/1, 49
peek byte/1, 90
peek byte/2, 90
peek char/1, 87
peek char/2, 87
peek code/1, 87
peek code/2, 87

permission (property), 148

permutation/2, 122
phrase/2, 111
phrase/3, 111
popen/3, 67, 153
portray/1, 96, 100
portray clause/1, 99
portray clause/2, 99, 138

portrayed (option), 96 position (property), 74 postfix op (property), 120

predicate property/2, 65
prefix/2, 123

prefix op (property), 120

print/1, 96, 98

print/2, 96, 100, 105, 106

print to atom/2, 105
print to chars/2, 106
print to codes/2, 106

priority (option), 97 private (property), 65 Prolog flag, 14, 36, 45, 59, 64, 88, 90–92, 94, 103,

119, 133, 135, 136, 165, 198, 199

prolog copyright (flag), 133
prolog date (flag), 133

prolog file (property), 65

prolog file name/2, 136, 142

prolog line (property), 65

prolog name (flag), 133
prolog pid/1, 155
prolog version (flag), 133

public (property), 65 public/1 (directive), 41, 61 punct (token), 94

put/1, 109
put byte/1, 91
put byte/2, 91
put char/1, 88

                                               put char/2, 88
                                               put code/1, 88
                                               put code/2, 88

quoted (option), 16, 96

random (FD option), 179

                                               random/1, 140
                                               random/3, 140
                                               randomize/0, 139

read (mode), 70

read (permission), 147

                                               read/1, 92, 95

read/2, 92, 95, 104, 105

                                               read atom/1, 93, 95
                                               read atom/2, 93, 95, 103
                                               read from atom/2, 104
                                               read from chars/2, 104
                                               read from codes/2, 105
                                               read integer/1, 93, 95
                                               read integer/2, 93, 95, 103
                                               read number/1, 93, 95
                                               read number/2, 93, 95, 103
                                               read pl state file/1, 136
                                               read term/2, 92, 95

read term/3, 92, 95, 103–105

                                               read term from atom/3, 14, 92, 104
                                               read term from chars/3, 104
                                               read term from codes/3, 105
                                               read token/1, 94, 95

read token/2, 94, 95, 103–105

                                               read token from atom/2, 104
                                               read token from chars/2, 104
                                               read token from codes/2, 105

real file name (property), 147

                                               real time/1, 139

regular (permission), 147

                                               remove stream mirror/2, 68, 80, 80
                                               rename file/2, 145

reorder (FD option), 179

                                               repeat/0, 113

reposition (option), 71 reposition (property), 73 reset (option), 71, 81

                                               retract/1, 62
                                               retractall/1, 62

return (option), 183, 184

                                               reverse/2, 122

search (permission), 147

                                               see/1, 107
                                               seeing/1, 108
                                               seek/4, 75
                                               seen/0, 108
                                               select/3, 122

select/5, 68, 156, 159, 161

                                               send signal/2, 156

set bip name/2, 37, 135, 196
set input/1, 68, 69
set linedit prompt/1, 162
set output/1, 68, 70
set prolog flag/2 (directive), 45
set prolog flag/2, 45, 133
set seed/1, 139
set stream buffering/2, 68, 82, 157, 158
set stream eof action/2, 81
set stream line column/3, 78
set stream position/2, 68, 75
set stream type/2, 81, 158
setarg/3, 54
setarg/4, 54
setof/3, 66
shell/0, 151
shell/1, 151
shell/2, 151
singleton warning (flag), 134, 136

singletons (option), 55, 56, 92 size (property), 148

skip/1, 108
sleep/1, 156

smallest (FD option), 179 socket (permission), 147

socket/2, 158
socket accept/3, 160
socket accept/4, 160
socket bind/2, 159
socket close/1, 158
socket connect/4, 67, 158, 159
socket listen/2, 160
sort/1, 125
sort/2, 125
sort0/1, 125
sort0/2, 125

space args (option), 96

spawn/2, 152
spawn/3, 152

spy/1 (debug), 30, 32

spypoint condition/3 (debug), 30, 32
sr change options/2, 164
sr close/1, 164
sr current descriptor/1, 164
sr error from exception/2, 164
sr get error counters/3, 164
sr get file name/2, 164
sr get include list/2, 164
sr get include stream list/2, 164
sr get module/3, 164
sr get position/3, 164
sr get size counters/3, 164
sr get stream/2, 164
sr open/3, 164
sr read term/4, 164
sr set error counters/3, 164
sr write error/2, 164

                                               sr write error/4, 164
                                               sr write error/6, 164
                                               sr write message/4, 164
                                               sr write message/6, 164
                                               sr write message/8, 164

standard (FD option), 179 static (property), 65

                                               statistics/0, 138
                                               statistics/2, 138
                                               stop/0, 112
                                               stream line column/3, 77
                                               stream position/2, 74, 75
                                               stream property/2, 73, 74, 75, 79, 81

strict iso (flag), 36, 64, 134

string (token), 94

                                               sub atom/5, 115
                                               sublist/2, 123
                                               suffix/2, 123
                                               sum list/2, 125

syntax error (flag), 92, 134, 198

                                               syntax error (option), 92
                                               syntax error info/4, 95, 198
                                               system/1, 152
                                               system/2, 152
                                               system time/1, 139
                                               tab/1, 109
                                               tell/1, 107
                                               telling/1, 108
                                               temporary file/3, 149
                                               temporary name/2, 148
                                               term ref/2, 57

text (option), 70, 81 throw/1, 29, 37, 47, 199 tight (debug), 30

                                               told/0, 108

top-level, 13, 18, 23, 25, 112, 162, 203

                                               top level/0, 13, 112

trace/0 (debug), 17, 29

                                               true/0, 46

type (option), 70

type (property), 73, 147

                                               unget byte/1, 90
                                               unget byte/2, 90
                                               unget char/1, 88
                                               unget char/2, 88
                                               unget code/1, 88
                                               unget code/2, 88
                                               unify with occurs check/2, 50

unknown (flag), 134

unknown (permission), 147

                                               unlink/1, 146

user (property), 65

user, 107, 108, 136, 141, 142 user input, 68, 72, 107, 108 user output, 68, 72, 107, 108

user time/1, 139

value method (FD option), 179 var (token), 94

var/1, 49
variable method (FD option), 179

variable names (option), 55, 56, 92 variables (option), 92 vector max (FD), 165, 167, 173

wait/2, 155

WAM, 11, 20, 21, 32 wam debug/0 (debug), 29, 32 warning (option), 92 Warren Abstract Machine, see WAM

working directory/1, 144

write (mode), 70 write (permission), 147

write/1, 96, 98

write/2, 96, 105, 106

write canonical/1, 96, 98
write canonical/2, 96, 105, 106
write canonical to atom/2, 105
write canonical to chars/2, 106
write canonical to codes/2, 106
write pl state file/1, 24, 136
write term/2, 96

write term/3, 16, 31, 96, 105, 106

write term to atom/3, 105
write term to chars/3, 106
write term to codes/3, 106
write to atom/2, 105
write to chars/2, 106
write to codes/2, 106
writeq/1, 96, 98

writeq/2, 96, 105, 106, 191

writeq to atom/2, 105
writeq to chars/2, 106
writeq to codes/2, 106